### v1

# Linying Lv

Abstract
I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports. Using embeddings from state-of-the-art large language models, I show that textual information in analyst reports explains 10.19% of contemporaneous stock returns out-of-sample, a value that is economically more significant than quantitative forecasts. I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns. The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power. Expressing these findings in economic terms, I estimate that early acquisition of analysts’ reports can yield significant profits. Analysts’ information value peeks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data. JEL Classification: G11, G14, G24
Keywords: Sell-side Analysts, Value of Information, Large Language Model, Explainable AI
“No one ever made a decision because of a number. They need a story.”
― Daniel Kahneman
1 Introduction
Sell-side analysts play a crucial role in financial markets by processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. While prior studies have demonstrated the incremental information value of report text, these analyses have predominantly centered on the sentiment dimension (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information perceived as at least as important as the quantifiable outputs. For example, according to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is currently limited, and it remains unclear which specific component of analyst reports holds the most value for investors.
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning, capturing more granular and rich information than before. This advancement facilitates a systematic analysis of the comprehensive outputs from sell-side analysts. For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the ”earnings analysis” topic, the second provides additional contextual information that LLMs can interpret and analyze.
I quantify the value of information in analyst reports in three general steps: analyst outputs representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices. For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs. Each report is mapped into a structured embedding space using the LLMs’ transformer architecture, allowing these to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2023).
I evaluate the information content of analyst reports using an R-squared measure to quantify how well the information set explains stock returns. A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings. To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients. This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns. Additionally, I assess the explanatory power of both quantitative and qualitative measures for abnormal stock returns using out-of-sample R-squared values with a recursively expanding window.
I compile a comprehensive corpus of analyst reports from the Mergent Investext database, merging these written reports with detailed analyst forecast summaries from the I/B/E/S database. The final sample comprises 122,252 analyst reports on S&P 100 constituent firms spanning from 2000 to 2023. Using this sample, the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually. Economically, a one standard deviation increase in report text favorableness raises the realized CAR by 120 basis points, compared to 100 basis points for forecast revisions. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content winthin analyst report text.
The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations. My analysis reveals that both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions. Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%. In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
Since a notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al. (2014), Kim and Song (2015), Lobo et al. (2017), and Barron et al. (2017), I examine the information content of analyst reports both within and beyond the earnings announcement window. This investigation reveals that the information content of analyst reports peaks in the first week following earnings announcements, supporting Huang et al. (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients. Using separately trained models on samples within and beyond earnings announcement windows, I find that the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods. This significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
To address potential concerns that analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models. By doing so, I can isolate the unique information of analyst reports. The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights, highlighting the different information sets provided by the two sources.
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question. To quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers. This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
To translate the statistic measure of R-squared into economic terms, I quantify the dollar value of analyst information to strategic investors using Kadan and Manela (forthcoming)’s framework, which builds on Back et al. (2000) imperfect competition insider trading model. The information value can be interpreted as the total expected profits for strategic investors either because some investors have early access to analyst views or because they conduct independent research and derive insights comparable to those of analysts. This multi-client framework provides a realistic foundation for analyzing the complex dynamics of information dissemination in analyst reports. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
The findings reveal significant value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports are \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million from both combined. The annualized aggregate information value is estimated at \$6.89 million, based on an average of 15 days with reports released per year. According to Kadan and Manela (forthcoming), this estimate represents a conservative lower bound. Notably, these values increase for large stocks, bold analysts, and reports released promptly following earnings announcements.
This paper relates to three strands of literature. First, This paper relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights. In contrast, my work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations. A direct comparison with the Naive Bayes tone classification used in Huang et al. (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone. To quantify these gains, I employ a BERT-based tone measure, which increases the out-of-sample $R^{2}$ significantly by 2%. Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure to above 10%. The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information, offering a more nuanced and comprehensive understanding of analyst narratives.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2023), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in financial contexts. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI. This approach enables a precise assessment of the relative importance of the topics discussed in analyst reports.
Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010). Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends prior work by quantifying the economic dollar value of such information, accounting for trading costs.
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation. Section 3 summarizes data construction and empirical results. Section 4 concludes.
2 Methodology
In this section, I introduce the large language models used for topic classification and text embeddings, the information content estimation of out-of-sample $R^{2}$, the strategic information value estimation of analyst reports, and the decomposition metric used to evaluate the information content of specific topics. [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
2.1 Large Language Models
Large Language Models such as Bidirectional Encoder Representations from Transformers (BERT), Generative Pre-trained Transformer (GPT), and Large Language Model Meta AI (LLaMA) are trending in literature for understanding textual data. These models represent a paradigm shift in NLP, offering capabilities that go beyond traditional methods like word-to-vec (W2V) or Latent Dirichlet Allocation (LDA) topic models. In this study, I harness the power of these advanced generative AI models to conduct a comprehensive analysis of analyst reports. The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors. This approach not only enriches information extraction but also provides a nuanced understanding of subtle linguistic cues and thematic elements that may influence market reactions.
2.1.1 ChatGPT Prompting
To effectively categorize the topics discussed in analyst reports, I employ ChatGPT prompting to extract exclusive and consistent topics and avoid ad-hoc classification. By leveraging AI, my goal is to identify the primary topics covered in these reports and assign each sentence to a single, specific topic category. This approach ensures a systematic and coherent classification of the content within analyst reports.
In the first step, to derive manageable and meaningful topics, I input a random sample of 100 analyst reports into ChatGPT, requesting an analysis of the content to identify high-level categories.
Prompt 1:Please read the provided text file of sell-side analyst reports carefully. What are high-level mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
It ends up with 13 categories defined as follows.
  • Executive Summary: Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
  • Company Overview: Offers a comprehensive description of the company, including its history, business model, and key products or services; discusses the company’s organizational structure, management team, and corporate governance; analyzes the company’s mission, vision, and strategic objectives.
  • Industry Analysis: Provides an in-depth analysis of the industry in which the company operates; includes information on market size, growth trends, and key drivers; discusses the regulatory environment, technological advancements, and other external factors affecting the industry; analyzes the industry’s competitive dynamics and the company’s position within the industry.
  • Competitive Landscape: Identifies the company’s main competitors and their market share; compares the company’s products, services, and pricing strategies with those of its competitors; analyzes the strengths and weaknesses of the company and its competitors; discusses potential new entrants and substitutes that could disrupt the competitive landscape.
  • Financial Analysis: Analyzes the company’s revenue, expenses, profitability, assets, liabilities, shareholders’ equity, liquidity, solvency, capital structure, cash flows, and key financial ratios; compares the company’s financial performance with that of its competitors and industry benchmarks; discusses trends in the company’s financial performance over time.
  • Business Segments: Provides a detailed analysis of the company’s various business segments or divisions; discusses the financial performance, growth prospects, and challenges of each segment; analyzes the contribution of each segment to the company’s overall revenue and profitability.
  • Growth Strategies: Discusses the company’s strategies for driving future growth, such as organic growth initiatives, product innovations, and geographic expansions; analyzes the company’s mergers and acquisitions (M&A) strategy and potential targets; examines the company’s investments in research and development (R&D) and marketing.
  • Risk Factors: Identifies and analyzes the key risks facing the company, such as market risks, operational risks, financial risks, and legal/regulatory risks; discusses the potential impact of these risks on the company’s financial performance and growth prospects; examines the company’s risk management strategies and mitigation measures.
  • Management and Governance: Provides an overview of the company’s management team, including their experience, expertise, and track record; analyzes the company’s corporate governance practices, such as board composition, executive compensation, and shareholder rights; discusses the company’s succession planning and key person risks.
  • Environmental, Social, and Governance (ESG) Factors: Analyzes the company’s performance and initiatives related to environmental sustainability, social responsibility, and corporate governance; discusses the potential impact of ESG factors on the company’s reputation, risk profile, and financial performance; examines the company’s compliance with ESG regulations and industry standards.
  • Valuation: Estimates the intrinsic value of the company’s shares using various valuation methodologies, such as discounted cash flow (DCF) analysis, relative valuation multiples, and sum-of-the-parts analysis; compares the company’s valuation with that of its peers and historical benchmarks; discusses the key assumptions and sensitivities underlying the valuation analyses.
  • Investment Thesis: Summarizes the key reasons for investing (or not investing) in the company’s shares; discusses the potential catalysts and risks that could impact the company’s valuation and stock price performance; provides a target price or price range for the company’s shares based on the valuation analyses and investment thesis.
  • Appendices and Disclosures: Includes additional supporting materials, such as financial statements, ratio calculations, and detailed segment data; provides important disclosures, such as the analyst’s rating system, potential conflicts of interest, and disclaimers; discusses the limitations and uncertainties of the analysis and the need for further due diligence by investors.
The Financial Analyses category is further divided into four subcategories: Income Statement Analyses, Balance Sheet Analyses, Cash Flow Analyses, and Financial Ratios. Additionally, a category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the predefined topics. In total, this process generates 17 distinct topics for classification. The distribution of topics is relatively stable across years, as illustrated in Figure 1. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, a random sample of 17,028 sentences from 100 analyst reports is selected from analyst reports. I employ the GPT-4-Turbo model through the ChatGPT API to categorize each sentence in this sample into a single, relevant topic.
Prompt 2: Please read the following sentence from the sell-side analyst report of the company {firm} ({ticker}) carefully. Determine which category of information it belongs to in the following 17 categories: {categories}. Pay attention to the sentence in the context of the report.
Output your response in JSON format.
Here is the sentence from the analyst report: {sentence}.
This training sample serves as the foundation for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Conversely, Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level. Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in focus on these issues.
2.1.2 Text Embeddings
Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary. This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand. [^2] BERT uses 30,000 tokens, OpenAI GPT models use around 50,257 tokens, and LLaMA models use approximately 32,000 tokens.
To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”. This granular approach allows for nuanced interpretation of complex terms. For instance, ‘next-gen workloads’ becomes five separate tokens: ‘next’, ‘-’, ‘gen’, ‘work’, and ‘loads’.
At the heart of LLMs lies the transformer architecture. Introduced by Vaswani et al. (2017), the transformer architecture revolutionized the field of NLP with its novel approach to sequence-to-sequence tasks. Unlike recurrent neural networks (RNNs) and long short-term memory networks (LSTMs), transformers use a self-attention mechanism that processes input sequences in parallel. This approach allows the model to weigh the importance of different words in relation to each other, effectively capturing context regardless of word distance. The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance in various NLP tasks.
Text embedding, a core output of the transformer architecture, transforms language into a format suitable for computational analysis. These embeddings are dense, continuous vectors that capture semantic and syntactic information. In models like BERT, GPT, and LLaMA, self-attention mechanisms generate these embeddings, encoding contextual information for each token. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. The embedding dimensions vary: BERT uses 768 or 1,024, GPT-2 Small uses 768, and LLaMA-2-13B model uses 5,120. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The tokenizer processes the text data, converting it into tokens that the model can understand. By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports. These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and text classification. The reason I employ LLaMA2 as a baseline Language Model is due to its 4,096-token input limit, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens.
2.2 Information Content
I use out-of-sample $R^{2}$ as a measure for the information content of analyst reports. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
$CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (1)
[…]
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ is the structured representation of information from analyst report $j$ on stock $i$ released at day $t$. I use the $60\%$ of data as the initial training sample. Therefore, the out-of-sampel test starts in 2015 and ends in 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ as
$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum CAR_{[-1,+1],it}^{2}},$ (3)
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation of the model. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
To evaluate the statistical significance of the $R_{\mathrm{oos}}^{2}$, I employ a modified version of the Diebold and Mariano (2002) test for comparing predictive accuracy between two models. Given the strong cross-sectional dependence in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional average of prediction errors from each model rather than individual returns.
2.3 Information Value
I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal based on the theoretical framework of Back et al. (2000):
$\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, multiplied by the closing price of the day prior to the [t-1, t+1] window. [^3] Due to risk neutrality, the equilibrium of the continuous-time Kyle model holds if $\tilde{v}$ is an unbiased signal of the asset price. The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
Specifically, $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ denotes the number of analyst reports about stock $i$ on day $t$. To estimate the price impact $\lambda_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals. The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^4] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression model used is
$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
where $\widehat{\lambda}_{it}$ represents the estimated price impact (or Kyle’s lambda). In this context, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$. Thus, each regression has $1170(=390*3)$ observations.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. This setting is particularly apt for analyzing the value of analyst reports, as it accounts for the correlated nature of information and the strategic interactions among market participants. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To tackle this issue, I apply the delta method in estimating the mean and variance of information value over sample $s$. The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
2.4 Shapley Value Decomposation
To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology, as detailed by Lundberg and Lee (2017) and applied to financial contexts by Chen et al. (2022). SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. In my context, this allows me to attribute the overall explanatory power of an analyst report to its constituent topics. Leveraging the additive nature of text embeddings, I design a novel decomposition method based on SHAP values. This approach enables me to systematically assess the relative importance of various topics in analyst reports.
First, LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ is the total number of words in the report. If the tokens can be classified into different topics, the embedding can be fully decomposed into different topic-specific embeddings:
$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
One concern in the architecture of a transformer model is that each token in the input sequence can pay attention to other words through a mechanism called “self-attention” or “scaled dot-product attention.” This mechanism allows the model to weigh the importance of each word relative to every other word in the sequence, thereby capturing contextual information. Consequently, the embedding of a word contains information from other words, which may pertain to different topics. The embedding of the word might predominantly reflect the context information from other topics. To isolate the self-attention across topics, I process the embedding for each sentence in analyst reports and take the token-weighted mean as the report embedding.
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context in analyst reports.
Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Empirically, the Shapley value can be written as:
$\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly. [^5] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fairly distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions to analyst reports.
3 Data and Empirical Results
This section presents the empirical analysis of analyst report information. I begin by examining the information content of these reports and decomposing it across different topics. I then quantify the strategic dollar value of analyst information. Cross-sectional analyses and robustness checks are conducted to validate the main findings.
3.1 Data
Sell-side analyst reports are obtained from the Mergent Investext Database. I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023. The initial reports are in PDF format, and I use Adobe Acrobat to extract structured text from each report. The segmented text is chunked into sentences. To get clean content of analyst discussions, I train a BERT model to remove the boilerplate segments from each report. [^6] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2023), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports. The BERT model trained for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
The Mergent Investext Database provides information on analyst names and the release date for each report. To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al. (2023). I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database. Following Huang et al. (2014), I define the matching window for each report as 2 days before the announcement date to 2 days after the revision date, accounting for multiple reports released between these dates. [^7] In I/B/E/S, each analyst forecast for a specific firm has two time stamps. The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one. The revision date (REVDATS) is the date the forecast becomes invalid or is revised. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and this time window. The numerical measures include earnings forecasts, stock recommendations, and price target information from analyst reports. I also compile firm characteristics from the CRSP and Compustat databases. [^8] See Table A2 and Table A3 for detailed definitions and summary statistics. After excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 firms from 1,305 analysts across 140 unique brokerage firms.
Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts for each year, or across the Fama-French 12 industries. It also reports the average number of pages and tokens per report, which are important indicators of report lengths. Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al. (2023), and the average number of pages per report is consistent with that in Huang et al. (2014). [^9] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline. This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
The daily stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023. To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
In Section 3.2.2, I explore textual information from earnings conference calls and subsequent analyst reports. The earnings call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005101010Available at https://seekingalpha.com/earnings/earnings-call-transcripts. These transcripts provide comprehensive accounts of the calls, including executive presentations and Q&A sessions with analysts, offering valuable insights into manager-analyst interactions. I sampled earnings call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks. This process yielded a final sample of 7,909 firm-quarter-level conference calls.
Given my goal to estimate the information value of analyst outputs, I need to define the information set available to investors. The numerical information and variables used as inputs to machine learning models are as follows:
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ indicates target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $Boldness$ is an indicator variable that takes the value of 1 if revisions are either above both the consensus and previous forecasts or below both. $SR$ represents stock recommendation, following the I/B/E/S rating system, where a rating of 5 denotes a strong sell and a rating of 1 indicates a strong buy. $ERet$ represents the 12-month return forecast, calculated by scaling the 12-month price target by the stock price 50 days before the report release. $Prior\_CAR$ represents the cumulative abnormal return over time period [-10, -2].
Firm-Level Measures: $Size$ is measured by market equity. The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value. Following Huang et al. (2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value. The $Miss$ variable is a binary indicator set to one if $SUE$ is positive. $TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding. The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures the firm’s product market fluidity.
Industry-Level Measures: $IndustryRecession$ is an indicator variable equal to one if the firm’s industry (defined by Fama-French 48 classification) experiences a negative monthly return in the bottom quintile.
Macroeconomic Measures: The $TimeTrend$ variable captures the temporal dimension by recording the number of years elapsed since the beginning of the sample period.
3.2 Information Content of Analyst Reports
In this section, I provide yearly $R^{2}_{\mathrm{OOS}}$ of analyst reports for explaining contemporaneous cumulative abnormal returns to examine the information content of analysts’ quantitative and qualitative outputs. Furthermore, I attribute the information content of analyst report text across topics using Shapley value decomposition.
3.2.1 Information Content of Quantitative and Qualitative Information
Table 2 compares the information content embedded in numerical and textual measures, quantified through their $R^{2}_{\mathrm{OOS}}$. The estimation model used in this table is the Ridge regression, where the objective variable is $CAR_{[-1,+1]}$ centered at the analyst report release date. For reports released on non-trading days, I adjust t=0 to the next trading day as recorded in the CRSP database. The estimation methodology employs an expanding window approach, constructing training sets from 2000 to the year preceding each target year. I estimate separate Ridge regression models for each input configuration, conducting yearly out-of-sample estimations of $\widehat{CAR}_{[-1,+1]}$, allowing for dynamic assessment of model performance over time.
Panel A of the table delineates the performance of four distinct input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) comprehensive numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, the 5,120-dimensional full-context representations of each report generated by LLaMA-2-13B model; and (4) a hybrid input combining revision measures with textual embeddings.
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for the revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions. The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features. I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4). The Diebold-Mariano test follows a standard normal distribution under the null hypothesis of equal predictive accuracy between two competing forecasting models, allowing similar interpretations of t-statistics as in regression analysis (Gu et al., 2020). The models incorporating revision and numerical inputs significantly outperform a zero-estimation benchmark.
Columns (5) and (6) elucidate the performance of text embedding input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20. This underscores the significant information content of the report text in explaining contemporaneous stock returns. The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession. Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs. Language model outputs may incorporate future information that should not be available. Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period. The Diebold-Mariano test comparing revision-only and text-only inputs yields a t-statistic of 1.66, highlighting the substantial information content in the report text.
In columns (7) and (8), I further investigate the performance of combining both revision and text input in the Ridge regression. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model. I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5). The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$. The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by market investors.
To address the concern that the explanatory power might be derived from quantitative information in analyst reports, I conduct robustness checks using content with numbers removed. In Table A4, I exclude numerical data from the written reports before generating text embeddings. The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, still significantly higher than that of quantitative revision measures (t-statistic = 2.40). This suggests that the superior performance of the text-based model is not driven by numerical information within the reports.
LLMs are known to exhibit improved language skills as the size of their parameters increases. I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI GPT-3 embedding model, and LLaMA-3 model. The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between each model and the LLaMA-2 model. The input in Panel B is report text, comparable to column (3) in Panel A. A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark. Overall, the BERT and OpenAI models underperform the LLaMA-2 model at the 1% level for each year. The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model. The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs. The pairwise comparisons between LLMs also shed light on how the scale of the model influences its performance capabilities.
To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al. (2014) and Asquith et al. (2005), respectively. Additionally, I examine the abnormal stock return on the days that reports are released. In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023. In Figure A3, I plot the information value of analysts’ information by industry, revealing a pattern consistent with the information content results.
In Table A6, I present the information content of analyst report text across the Fama-French 12 industries. The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables. Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors. The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value.
To assess economic magnitude, I perform OLS regressions as specified in Equation (10):
$CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (10)
This model incorporates the out-of-sample expanding estimation of $CAR_{[-1,+1]}$ using report text embeddings as input. For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014. The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension. For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
Table 3 summarizes the summary statistics of the sample used and the regression results. The out-of-sample period starts in 2015. I constrain the samples to include only those with valid previous and current stock recommendations, earnings forecasts, and target prices. The final sample for estimating Equation (10) contains 28,837 observations.
Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al. (2014). Column (2) reveals that $\widehat{CAR}{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2. In column (3), the inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% to 14.9%, underscoring the incremental value of text information.
The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by quantitative analyst revisions. Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the three-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics. The $\widehat{CAR}{rev}$ model alone achieves an $R^{2}$ of 8.9%, while the combined input model improves this to 14.7%. Economically, a one standard deviation increase in $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises. The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged. This stability in coefficients suggests that the market-valued analyst information is distinct from, and not merely chasing back on, past returns or recent news events.
3.2.2 Information Content around Earnings Announcements
Some studies have emphasized the role of analysts surrounding corporate earnings announcements (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information events reinforce each other in sequence. I, therefore, further study the information content of sell-side analyst reports around this most significant corporate event.
Following Chen et al. (2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week. Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week. A finding consistent with Huang et al. (2018) is that an overwhelming number of analyst reports are released immediately following the earnings announcements. Consistent with Huang et al. (2018), a large number of analyst reports are released immediately after earnings announcements. Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17. This performance drops sharply in the second week, turns negative in the third and fourth weeks, and does not recover thereafter.
The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports during earnings announcement weeks. To account for potential temporal variation, I train Ridge models using separate samples for periods surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports relative to earnings announcements. Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained outside these periods. Additionally, the $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen. This trend suggests that analysts react swiftly to earnings announcements, with the information content of their reports declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week.
One potential concern in assessing the value of analyst reports is that analysts might merely be repeating the content of earning announcement calls, with the earnings announcement itself driving market reactions. If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls. To address this issue, I use earnings conference call transcripts data to differentiate the information content between the two text sources. The analysis focuses on reports released within a one-day window following earnings announcements, as this is when analysts are most likely to piggyback on the announcements.
In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcripts as input. [^11] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C. This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study. I truncate the text at the length limits to maintain maximum information. The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16. with a Diebold-Mariano (DM) t-statistic of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods of both texts. The input comprised the mean of all report text embeddings on the day following the earnings announcement. This firm-time level Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
Incorporating both analyst report embeddings and earnings conference call embeddings in the Ridge regression resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3). The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports. These findings collectively refute the idea that analysts merely restate information from earnings announcements. Instead, they emphasize the unique and timely value analysts bring to the market.
3.2.3 What Content in Analyst Reports is Valued by Market Investors?
Given the comprehensive discussion in equity reports, what specific content in analyst reports is the most valuable to the market? I employ the Shapley value approach in Section 2.4 to attribute the $R^{2}_{\mathrm{OOS}}$ of the full text to 17 topics. In this section, I use the token-weighted average of each sentence embedding as the report embedding. This sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another. The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings. This reduction roughly quantifies the magnitude of contextual information content in analyst reports.
Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations. These reductions are then averaged into a single importance measure for each topic. By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring a comprehensive attribution of explanatory power. Figure 2 ranks the Shapley values of 17 topics. Topics are ordered from left to right, with the most important topics on the left and the least important on the right. Each topic is represented by a bar, with a color gradient (from lightest to darkest) indicating the magnitude of its importance in terms of Shapley values.
The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports. The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each comprising 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$. Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively. The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
The robustness of these findings is further reinforced by additional analyses. As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses, underlining the consistency of the results.
Figure A5 in the Online Appendix reveals how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time. The annual variations align with the full-context embedding results in Table 2, exhibiting a consistent story across different text embeddings. A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy. Despite this volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable, indicating that analyst discussions of income statements consistently provide valuable information, even during market disruptions.
Figure A6 further supports this by showing a consistent pattern of key topics across industries. This suggests that the market values Income Statement Analyses and Financial Ratio Analyses universally, beyond specific sectors.
To investigate whether this topic ranking is driven by analysts’ expertise in certain topics (supply effect) or investors’ differential responses to different topics (demand effect), Figure A7 compares Shapley values across analysts with varying experience levels and from different-sized brokerage firms. The remarkable consistency in topic importance across all four subsamples (experienced vs. inexperienced analysts, large vs. small brokerages) eliminates the concern that the topic ranking merely reflects analysts’ varying skill levels in different topics.
Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements. Drawing inspiration from the literature, particularly Huang et al. (2018) and Blankespoor et al. (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
To facilitate this analysis, I further categorize sentences within the Income Statement Analyses topic into two sets of sub-topics: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a sophisticated prompt-based classification system, as detailed in Prompt 3:
Prompt 3: Please read the following sentence from a sell-side analyst report for the company {firm} ({ticker}) and classify it based on two criteria:
Information Type:
Information Acquisition (0): Sentences that directly report quantitative financial data, such as earnings, revenue, expenses, or other metrics. Example: ‘Reported EPS of \$1.40 beat consensus of \$0.94 and our estimate of \$1.17.’
Information Interpretation (1): Sentences that provide analysis, context, or interpretation of the financial data, such as discussing trends, comparing performance to forecasts, assessing market impacts, or considering strategic implications. Example: ‘Biogen reported 1Q12 MS franchise sales below our estimate and the Street, as Avonex sales were negatively impacted by unfavorable distribution channel dynamics.’
Time Reference:
Realized Income (0): Sentences referring to concrete, historical results from completed periods. Example: ‘The company’s revenue for the fiscal year 2022 was \$500 million.’
Expected Income (1): Sentences containing subjective predictions, or expectations for future periods. Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
Output your classification as two comma-separated numbers: the first for Information Type (0 for Acquisition, 1 for Integration) and the second for Time Reference (0 for Actual, 1 for Forecast).
Output format: InfoType, TimeRef
Sentence to classify: {report_sent}.
Figure 3 offers a detailed breakdown of the $R^{2}_{\mathrm{OOS}}$ across 18 topics, with the Income Statement category further divided into sub-topics. Specifically, the income interpretation subtopic contributes approximately three times more to the $R^{2}_{\mathrm{OOS}}$ than income acquisition, underscoring the market’s preference for context and explanation over mere data reporting. Additionally, the income realization subtopic also contributes about three times more to the $R^{2}_{\mathrm{OOS}}$ than income expectation, indicating a stronger market preference for analysis of actual financial results over forward-looking projections. These findings suggest that the market values analysts’ interpretation of realized income to a greater extent than their ability to simply report public information or make expectations about future performance.
This insight is reinforced by the findings in section 3.2.2, which show that analyst reports issued promptly after earnings announcements are the most informative. Taken together, these results suggest that the market places the highest value on analysts who can quickly digest financial information from earnings announcements and provide insightful interpretations.
3.2.4 Comparison of Revision Reports and Reiteration Reports
Researchers have documented a robust relationship between market prices and analysts’ forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021). Analysts’ forecast revisions provide a significant and timely source of information to the financial market and may constitute a more valuable signal than reports that merely reiterate previous forecasts.
Therefore, I compare the information content of analyst report text associated with forecast revisions and forecast reiterations. Following the criteria of Huang et al. (2014), I consider a report as a revision report if its release date is within two days of the I/B/E/S forecast announcement date. For each forecast target in recommendations, target prices, and earnings forecasts, I categorize the report independently. Consequently, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 5 summarizes the out-of-sample performance of numerical and textual inputs in both reiteration and revision samples. Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions. When recommendations are revised, the explanatory power of various information sources becomes particularly pronounced. The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%. The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%. The combination of both quantitative and qualitative information sets further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%. This finding aligns with the literature that attributes large stock price changes to analyst recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions, considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), also reveal a clear contrast between revisions and reiterations. While target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ of 20.88%, t-stat = 11.18).
Earnings forecast revisions are the most frequent, comprising 56.57% of the reports. In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively. As Kothari et al. (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output. Interestingly, the $R^{2}_{\mathrm{OOS}}$ for report text in earnings forecast reiteration reports is negative, indicating no explanatory power for contemporaneous stock returns. However, when focusing on analyst earnings forecast revisions,, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant. The combination of these inputs yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
Collectively, Table 5 shows that report content is more informative when it accompanies forecast revisions rather than reiterations. Additionally, analyst reports associated with recommendation revisions demonstrate the highest value.
3.2.5 Comparison with Analyst Report Sentiment
Huang et al. (2014) emphasize the importance of analyst report text by extracting tone using a Naive Bayes machine learning approach. My study extends this research by incorporating more advanced techniques, including BERT-based tone measures and text embeddings.
I construct report-level tone measures using both the Naive Bayes approach and BERT classification. Following Huang et al. (2014), I manually label 10,000 sentences as a training sample for both models. Each sentence is categorized into positive, neutral, or negative using these pre-trained models, then average the tone at the report level to derive $Tone\_NB$ and $Tone\_BERT$ measures. I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
Table 6 presents a comparison of performance between text tone measures and text embeddings. Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures. The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al. (2014). A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, which is comparable with the 41 basis points increase in two-day abnormal return in Huang et al. (2014). Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics. Both coefficients of tone are significantly positive. Column (3) reports regression results with firm-fixed and year-fixed effects, consistent with the findings in column (2).
Columns (4)-(6) present regression results with BERT-based tone measures as explanatory variables. The coefficient estimates’ magnitude increases dramatically, along with significance levels. Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points. When both tone measures are included in an OLS regression in column (7), Naive Bayes-based tone measure coefficients become insignificant, suggesting BERT-based measures capture sentiment more accurately and effectively.
Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of extracted tone measures and embedding representations. The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows. The $R^{2}_{\mathrm{OOS}}$ increases from 9.49% to 12.27% after incorporating analyst revision and text embeddings, confirming the superior performance of state-of-the-art LLMs across various NLP tasks. The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The $R^{2}_{\mathrm{OOS}}$ difference between these models is not statistically distinguishable from zero for both Naive Bayes-based and BERT-based tone measures.
In general, the empirical results in Table 6 support two key findings regarding the superiority of LLMs compared to conventional NLPs. First, overall report text tone measures extracted using small-size LLMs are more explanatory for contemporaneous stock returns than those extracted using the Naive Bayes algorithm. Second, tone measures (extracted with both Naive Bayes and BERT) add no incremental value to text embeddings, suggesting that LLMs provide more comprehensive contextualized representations.
3.2.6 Information Content with Alternative Machine Learning Models
In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions. Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$. To ensure out-of-sample evaluation, I train the models yearly and estimate for the following year’s sample. The $R^{2}_{\mathrm{OOS}}$ is calculated by averaging the estimated and realized $CAR$ values. Additionally, I perform pairwise Diebold-Mariano tests using a zero benchmark to assess predictive accuracy.
I begin with PLS regression, summarized in Table 7. PLS is notable for extracting the most relevant common components from sentiment proxies that align with the target variable (Huang et al., 2015). If there is a strong common component in the text embeddings that is informative of stock returns, the PLS regression architecture should effectively capture this relevant information. Examining the model performance, I observe that PLS regressions produce an overall $R^{2}_{\mathrm{OOS}}$ of 8.26%. This measure of information content ranges from 2.14% in the 2021 sample to 12.23% in the 2015 sample, showing a similar pattern across years as seen in Ridge models. Diebold-Mariano tests indicate the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all periods, demonstrating strong evidence of information content in the analyst report text.
Next, I replicate the analyses with the XGBoost algorithm. Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees. Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information. Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns. Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable. As summarized in Table 7, the XGBoost algorithm underperforms with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
Neural Networks, forming the core of LLMs and the transformer architecture, have shown significant advancements in various NLP tasks (Cao et al., 2024; Li et al., 2023). In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain. Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text. I construct the Neural Network estimations as an ensemble of five models for each NN configuration. As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms. Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al. (2022).
The experiments with machine learning models yield noteworthy findings. First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns. Second, the superior performance of deep learning architectures emphasizes the importance of capturing contextual information and word interdependencies for textual analyses.
3.3 Information Value of Analyst Reports
How can the statistical measure $R^{2}_{\mathrm{OOS}}$ be translated into economic terms? In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3. This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal while considering price impact. In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
I first summarize the value of information from analyst reports in Table 8, covering the period from 2015 to 2023. Both the mean and standard deviation are derived using the delta method. For S&P 100 constituent stocks, the average three-day window information value from sell-side analysts is \$0.47 million, with a standard error of \$0.05 million. The mean information value of text alone is \$0.38 million, and the mean value for revisions is \$0.34 million. Considering that a stock typically has reports released on an average of 15 days each year, the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm is \$6.89 million. The narrow standard errors and confidence intervals across different information values indicate consistent estimates. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, indicating that \$1 million dollar order pushes the stock price up by 0.034%. These statistics illustrate that information from analysts is economically meaningful, with both textual information and revisions contributing significantly.
Figure 4 describes the evolution of information value from analyst reports over time. The solid line represents the strategic value of combined numerical and textual information, while the dashed line shows the value of textual information alone. Both lines exhibit significant fluctuations over the period. Despite these fluctuations, there is a general upward trend in information value. For most of the period, the combined value is higher than the textual value alone, suggesting that numerical information often complements the textual content in analyst reports.
The shaded area represents the 95% confidence interval. After 2020, there appears to be greater variability in the information value, as indicated by the wider confidence intervals and more pronounced peaks and troughs. This period corresponds with the global COVID-19 pandemic, which likely introduced additional uncertainty and volatility in financial markets, impacting the value of information from analyst reports.
These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions and external events. The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value. Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million. Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time. These alternative estimates support my main findings, reinforcing the substantial value of analyst information and the robustness of the approach.
3.3.1 The Value of Analyst Information in the Cross-section of Stocks
It is likely that the information value of sell-side analysts will vary in the cross-section of stocks. Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984). For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995). I plot the scatter of information value and market capitalization by stock in Figure A9. There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values. However, the relationship is not strictly linear, as there are notable exceptions. I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects. I also control for the book-to-market ratio in the regressions.
Table 9 shows the results. Stocks with larger market equity have significantly higher information value from analysts. Economically, column (4) means a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value. I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact for large stocks. This decomposition can be expressed logarithmically as:
$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
Panel B of Table 9 demonstrates that the main reason behind the positive size elasticity of information value is the higher liquidity of large stocks. Overall, analyst information does not appear to explain the higher contemporaneous return volatility of large stocks, yet it remains more valuable due to the higher liquidity of these stocks.
3.3.2 The Value of Analyst Information in the Cross-section of Analysts
I next examine how the information value of analysts differs with herding behavior. Prior research has stated that analysts making bold forecasts incorporate their private information more completely and provide more relevant information to investors than those making herding forecasts (Clement and Tse, 2005).
To investigate the cross-section of analyst characteristics, I regress the stock-analyst-day information value on an indicator of forecast boldness, controlling for firm-fixed effects, analyst-fixed effects, and year-fixed effects. I also control for broker size, as Jacob et al. (1999) suggests that larger brokerage firms provide better research. The analyst-specific information value measure is defined as
$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
Table 10 presents the results. In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts. On average, information from bold reports is 29% more valuable than that from herding reports. The magnitude remains largely unchanged when testing with information value derived from report text. However, the insignificant coefficient estimates for $Brokersize$ suggest that the value of analyst information is not necessarily tied to the size of the brokerage firm.
The subsequent research question is whether the association between boldness and information value is driven by higher information content or lower price impact. In Panel B, I regress the two decompositions of information value on the boldness indicator. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information. In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant. This finding suggests that while bold forecasts contain more valuable information, they do not necessarily lead to a higher price impact compared to herding forecasts.
3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. To expand on this, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 illustrates the information value of analyst reports between earnings announcements on a weekly basis, from the first week after the earnings announcement up to the $13^{th}$ week. The information value of analyst reports is highest immediately following earnings announcements, particularly in the first week, peaking at a sample mean of \$0.84 million. This value sharply declines by the second week and stabilizes thereafter, with minor fluctuations. These patterns suggest that the market values analyst reports most highly immediately after earnings announcements, with diminishing importance in subsequent weeks.
To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$). This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010). The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al. (2024). The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect information value. The model includes stock and year-fixed effects to control for time-invariant stock characteristics and market-wide yearly trends.
Table 11 presents regression results showing the influence of earnings announcements and trading volume on information value. Panel A indicates that the information value is significantly higher for reports released within one week of earnings announcements ($Week$), with positive and significant coefficients across all models. The interaction between $Week$ and $TradingVolume$ also shows a significant positive effect, suggesting that the information value increases with higher trading volumes around earnings announcements. The rationale is that investors may rely more on analyst opinions when there is greater disagreement in the interpretation of earnings information.
Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable. The interaction term ($Week\times TradingVolume$) is significantly positive in columns (1)-(4), indicating that explained return volatility rises with higher trading volumes during post-earnings periods. Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact. These findings suggest that analyst reports provide more valuable insights in the immediate post-earnings announcement period, particularly when investor uncertainty is high. This increased value likely reflects analysts’ ability to interpret complex earnings information, helping investors make sense of the latest financial data during these critical periods.
4 Conclusion
In this paper, I examine both quantitative and qualitative content of analyst reports using $R^{2}_{\mathrm{OOS}}$. The results show that report text contains more market-valued information content than quantifiable forecast revisions economically, especially right following earnings announcements. A Shapley value decomposition method reveals that analysts’ interpretation of realized income statements is the most valuable content within written reports. The text embedding representation extracts more comprehensive information from analyst reports beyond text sentiment.
I further translate the information content into dollar terms to provide a sense of economic magnitude. Using a measure of explained return volatility relative to price impact, I find that the annualized expected profit from early information acquisition of the analyst reports for an average S&P 100 constituent firm is larger than \$6.89 million. While these results are economically meaningful, readers should exercise caution in generalizing beyond the S&P 100 sample. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The significant potential profit from early access to analyst reports raises concerns about selective distribution. Such practices would violate FINRA Rule 2241, which aims to ensure the objectivity of research and prevent the selective distribution of research reports to preferred clients ahead of others. It is crucial to emphasize that this study presents a hypothetical scenario without assessing or endorsing actual market practices. Rather, by quantifying the potential profit of privileged access to analyst information, this research emphasizes the importance of regulations in ensuring fair disclosure.
A. Additional Figures
(a) (1) Executive Summary
(b) (2) Company Overview
(c) (3) Industry Analysis
(d) (4) Competitive Landscape
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (5) Income Statement Analysis
(b) (6) Balance Sheet Analysis
(c) (7) Cash Flow Analysis
(d) (8) Financial Ratios
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (9) Business Segments
(b) (10) Growth Strategies
(c) (11) Risk Factors
(d) (12) Management and Governance
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (13) ESG Factors
(b) (14) Valuation
(c) (15) Investment Thesis
(d) (16) Appendices and Disclosures
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
Figure A2: Large Language Model Architecture
This figure shows an example of how large language models process a sentence from an analyst report on Nvidia. The model breaks this sentence into tokens, and then generates two types of embeddings: token embeddings and position embeddings. These embeddings then pass through the core of the model - a stack of transformer layers. The output of each transformer layer is a series of vectors representing the tokens in their full context.
Figure A3: Analyst Information Value by Industry
This figure shows the estimated information value of analyst reports for Fama-French 12 industries from 2015Q1 to 2023Q4, quantified in millions of dollars. The solid green line marks the sample mean value of information, with green bars showing the 95% confidence intervals. The horizontal dashed red line marks the sample mean of stocks from all industries. Information value estimates are derived using the delta method and adjusted to 2020 levels using the Consumer Price Index (CPI).
Figure A4: Shapley Values Scaled by Length
This figure presents the SHAP values of topics scaled by the total number of sentences (Panel A) and the total number of tokens (Panel B) for topics with more than 100,000 sentences. The scaling process involves two steps: First, normalize the Shapley value of each topic by dividing it by the number of sentences in that topic. Then, adjust the normalized values to ensure the total sum of SHAP values remains consistent. This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
Figure A5: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
This figure shows the $R^{2}_{\mathrm{OOS}}$ of sentence-segmented report embeddings from 2015 to 2023. The color gradient indicates the magnitude of $R^{2}_{\mathrm{OOS}}$. The red line represents the SHAP value of the Income Statements Analyses topic across the years. The shaded area marks the pandemic recession in 2020.
Figure A6: Shapley Values for Topic Importance across Industries
This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value. For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings. The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
Figure A7: Shapley Values for Topic Importance by Analyst Characteristics
This figure illustrates how 17 different topics contribute to the predictive power of analyst reports across four subsamples, measured using SHAP values. The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), and large (Panel C) versus small brokerage firms (Panel D), with subsamples determined by yearly median. For each subsample, the bars show the SHAP values of individual topics, whose sum equals the out-of-sample R-squared ($R^{2}_{\mathrm{OOS}}$) from the sentence-segmented report embeddings. The red line indicates each topic’s relative importance, calculated as the topic’s SHAP value divided by the total $R^{2}_{\mathrm{OOS}}$.
Figure A8: Alternative Analyst Information Value Estimations Over Time
This figure presents the alternative estimations of analyst information value from 2015Q1 to 2024Q4, reported in millions of dollars and adjusted to 2020 dollars. Quarterly mean approximations are calculated using the delta method. TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Figure A9: Analyst Information Value by Stock
This figure shows the estimated information value of analyst reports for individual stocks from 2015Q1 to 2023Q4, quantified in millions of dollars. Estimates are derived using the delta method. Market equity, averaged daily at the time of each report release, is reported in billions of dollars. All values are adjusted to 2020 levels using the Consumer Price Index (CPI).
B. Additional Tables
Table A1: Topic-Specific Sentence Examples from Analyst Reports
Topics Examples
Executive Summary
Our key takeaways from CEO Jeff Immelt’s presentation at EPG were: Outlook for substantial EPS growth over 2010/12 driven by abatement of credit losses (\$8-9bn in 2010E tapering to \$4bn run-rate) and CRE impair. Looking ahead, guidance was tightened, essentially framing the Street, and commentary suggests a strong outlook for the balance of 2007. Key topics: 1) Update on ABTÕs portfolio of COVID-19 tests; 2) Libre trends, Libre 2 launch and expectations for Libre 3; 3) Elective surgery trends exiting Q1, expectations for 2021 and an update on recent and upcoming new product approvals; and 4) Global trends and impact on EPD and Nutrition.
Company Overview
Oracle Corporation, founded in 1977 and headquartered in Redwood Shores, California, is one of the largest and most prominent companies in the software space – and a technology bellwether. As it has grown, Microsoft has expanded into enterprise software with Windows Server, SQL Server, Dynamics CRM, SharePoint, Azure and Lync; hardware with the Xbox gaming/media platform and the Surface tablet; and online services through MSN and Bing. Altria Group, Inc., is the world’s largest producer and marketer of consumer products, and had revenues of \$80 billion in 2002.
Industry Analysis
According to our global Immunology market model, US Psoriasis (PsO) represented a \$7.7B market in 2016 and is expected to grow at a low-teens CAGR to \$11.8B in 2019E and  \$13B by 2021E driven by more highly effective therapies. Despite record prices, oil demand continues to grow, while supply growth lags and spare production and refining capacity is almost nonexistent. Add to that that some new advertising expectations from PricewaterhouseCoopers of a decline in ad spending of 12% worldwide and 15% in the U.S. for 2009 and continue to decline in 2010.
Competitive Landscape
According to Mercury Research, NVIDIA is now the 3rd largest chipset supplier (consisting of desktop and mobile chipsets, and integrated and non-integrated chipsets), shipping 5.4 million units in calendar Q3 for an 8.2% market share, versus Intel’s shipments of 51 million units (62% share), VIA’s shipments of 9.6 million units (14.4% share), SiS’s shipments of 5.3 million units (8% share), and ATI’s shipments of 4.4 million units (6.6% share). Further, competition in the CDK-4/6 space is rising with Verzenio (abemaciclib) & Kisqali launches placing downward pressure on Ibrance trajectory. We believe that Accenture has recognized that web services could compete directly with client/server as the new systems architecture.
Financial Statement Analysis: Income Statement Analysis
We also note that Alico cannot dividend any of its 2010 earnings to AIG, implying that the income recognized this quarter and throughout 2010 will benefit MetLife results. At the respective midpoints, sales of \$8.5 billion would be down 22% annually and 8% sequentially; and non-GAAP EPS would be down 39%. The growth in operating income from Q1 2009 is largely due to recoveries in partnership income.
Financial Statement Analysis: Balance Sheet Analysis
Long-term borrowings of \$6.59 billion at September 30, 2022 were modest compared to shareholders’ equity of \$37.3 billion. The company finished 2Q16 with \$21.4 billion in cash and short-term investments, up from \$15.6 Growth & Valuation Analysis GROWTH ANALYSIS RISK ANALYSIS billion at the end of 4Q15. Inventory turns improved to 4.5x in 4Q from 4.4x in the same period last year.
Financial Statement Analysis: Cash Flow Analysis
UPS generated \$2.3 billion in operating cash for the quarter. Assuming that the company completes a large portion of its current \$1 billion stock buyback plan in 2008, we estimate that cash per share will be about \$6.20 by the end of the year. We view eBays FCF generation as relatively defensible even in the case of a revenue shortfall.
Financial Statement Analysis: Financial Ratios
The company achieves average scores on our three main measures of financial strength: leverage based on debt-to-cap, profitability and interest coverage. P/S is at 0.6x and EV/EBITDA is 7.7x. The index members currently trade at an average of 16.3-times trailing earnings, which is below the five-year average of 19-times.
Business Segments
The company is organized into three businesses, software, representing the majority of the company’s total revenues, hardware systems and services. Results in 2008 also benefited from the absence of the significant level of mark-to-market losses in the company’s Gas Marketing segment in 2007. The company operates five distinct segments: Americas (71% of FY15 profit); Europe, Middle East, and Africa (4% of FY15 profit); China and Asia Pacific (12% of FY15 profit); Channel Development (13% of FY15 profit).
Growth Strategies
Visa has been especially active on the acquisition front over the last several months. Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle. Specific areas where more investment may be needed: Over the past decade, ROK management often claim that Process automation represents the growth opportunity for the company, but its sales in this market have barely grown in recent years, despite its much smaller sales base compared with established incumbents.
Risk Factors
Risks to our BUY thesis have to do with global competition, changing user behavior, global macro uncertainty, and anything else that can affect FB’s relationship with members, its advertisers or its publishing partners. Risks to achieving our price target include: 1) Apple crushing PayPal; 2) increasing competition in the payments space; 3) heavy investment spending on marketing, point of sale, or technology; and 4) legislative action. In addition to the expenses incurred by patent challenges, product liability and other legal suits could occur and lead to additional liabilities and revenue loss, which could substantially change our financial assumptions.
Management and Governance
Top management changes can be unsettling, and the resulting uncertainty has caused 3M shares to decline. Chairman, President, and CEO Charles Ergen beneficially owns about 53.6% of DISH’s equity securities and has 90.5% voting power. Bill Johnson, the present CEO of Progress Energy, will become president and chief executive officer of the new company.
Environmental, Social, and Governance (ESG) Factors
The assessment of ESG (Environmental, Social & Governance) risks by Baptista Research includes a wide range of considerations that pertain to the long-term sustainability of a company. Sustainalytics assesses the degree to which a company’s enterprise business is affected by ESG issues. Failure to adequately address social risks like labor disputes and community relations could jeopardize the company’s social license to operate in certain regions.
Valuation
Our DCF derives an intrinsic value of ¿\$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC. We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment. Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
Investment Thesis
The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road. As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high. Clearly, our Ford Investment Thesis, which was based in large part on our belief that Ford would be able to offset headwinds (slowing cyclical tailwinds in North America, weakness in South America, weak growth in Europe, slowing growth in China, and regulatory cost headwinds), has been thrown into question.
Appendices and Disclosures
Although the information contained in the subject report has been obtained from sources, we believe to be reliable, its accuracy and completeness cannot be guaranteed. For a complete discussion of the risk factors that could affect the market price of a company’s shares, refer to the most recent Form 10-Q or 10-K that a company has filed with the Securities and Exchange Commission. The Benchmark Company, LLC makes every effort to use reliable, comprehensive information, but we make no representation that it is accurate or complete.
Table A2: Numerical Variables Description
This table shows the definitions of numerical measures.
Numerical Measures Definition and/or sources
Panel A: Firm Level Measures
Size The market value equity of the firm ($CSHOQ*PRCCQ$) at the end of the quarter prior to report release.
BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of quarter prior the report release.
Prior_CAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement. Consensus EPS is the median value of 1-year EPS forecast within 90-days window of all analysts following the firm. The unexpected earnings is scaled by price per share at the fiscal quarter end.
AbsSUE Absolue value of SUE, representing the distance between realized EPS and EPS expectation.
Miss Dummy variable that equals one if the actual EPS is less than the last consensus forecast, and 0 otherwise.
Trading volume Trading volume at earnings announcement day (or the first trading day post earnings announcement), calculated as $VOL/SHROUT$.
Distance to default The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
Fluidity A measure of firms’ product market competition introduced by Hoberg et al. (2014). The data is compiled from the Hoberg and Phillips database.
Panel B: Industry Level Measures
Industry recession An indicator variable that equals one if the FF-48 industry return is negative and in the bottom quintile of FF-48 industry returns and zero otherwise.
Panel C: Macroeconomic Measures
Time trend (ttr) The number of years elapsed from the beginning of the sample.
Panel D: Report Level Measures
$REC\_REV$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
$EF\_REV$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
$TP\_REV$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
Boldness An indicator variable that takes the value of 1 if the EPS forecast revision is above both the analyst’s own prior forecast and the consensus forecast, or else below both, and zero otherwise.
SR Stock recommendation from I/B/E/S rating, with 1 being the most bullish (Strong Buy) and 5 being the most bearish (Sell), based on the ratings provided by the Institutional Brokers’ Estimate System (I/B/E/S).
ERet 12-month return forecast by scaling the 12-month price target by the stock price 1-day before release.
Table A3: Sumamry Statistics for Numerical Measures
This table reports summary statistics of numerical measures.
Mean Std p25 P50 P75 N
Miss 0.35 0.48 0.00 0.00 1.00 102776
Trading volume 18.19 21.98 7.24 11.46 20.29 99908
Prior_CAR 0.00 0.05 -0.02 0.00 0.02 122251
Size 2.39 0.11 2.33 2.39 2.46 117534
BtoM 0.46 0.43 0.18 0.32 0.61 117534
Distance to default 0.00 0.01 0.00 0.00 0.00 86693
Fluidity 7.17 3.66 4.40 6.57 9.44 114942
Time trend 13.02 5.71 9.00 13.00 17.00 122252
Industry recession 0.14 0.35 0.00 0.00 0.00 122252
BrokerSize 71.92 56.71 26.00 53.00 113.00 119233
Firm Experience 7.68 6.72 3.00 6.00 11.00 111192
Number of firms 20.08 17.08 14.00 18.00 23.00 119233
REC_REV 0.00 0.16 0.00 0.00 0.00 120673
EF_REV 0.00 0.00 0.00 0.00 0.00 90625
TP_REV 0.01 0.07 0.00 0.00 0.00 84108
ERet 0.19 0.23 0.07 0.18 0.31 84108
Boldness 0.75 0.43 1.00 1.00 1.00 81182
SR 2.77 0.84 2.00 3.00 3.00 122252
Table A4: Information Content of Analyst Reports with Numbers Removed
This table presents the $R^{2}_{\mathrm{OOS}}$ using analyst reports with numbers removed. The input sets include revision numerical measures, text embeddings from number-free report text, and a combination of both. The estimation targets are the three-day $CAR$ centered around the analyst report release date. The machine learning model employed is Ridge regression, with training samples expanding yearly. The $t$-statistic for the $R^{2}_{\mathrm{OOS}}$ is calculated following Gu et al. (2020). In columns (2), (4), and (6), the benchmark estimation is set to zero. Columns (3)-(1) and (5)-(3) present pairwise tests comparing the performance of models using revision-only input versus text-only input, and text-only input versus combined text and revision input, respectively.
Year Revision only t-stat Text only t-stat Revision plus text t-stat t-stat t-stat
(1) (2) (3) (4) (5) (6) (3)- (1) (5)- (3)
2015 10.30% 3.59 13.19% 5.87 11.10% 3.24 3.97 -1.45
2016 15.47% 11.80 13.73% 5.00 17.91% 9.02 -0.81 3.13
2017 9.16% 5.28 11.66% 5.98 12.33% 6.12 10.58 1.56
2018 9.79% 2.98 11.46% 9.12 13.62% 7.97 1.22 4.09
2019 9.76% 15.69 12.61% 14.73 13.90% 15.17 4.26 2.79
2020 5.20% 3.43 4.42% 4.38 6.98% 5.81 -0.70 8.37
2021 5.31% 4.80 8.77% 5.06 10.66% 29.67 1.54 1.56
2022 10.24% 6.16 16.71% 10.47 17.90% 8.94 9.93 2.38
2023 5.09% 3.86 9.27% 3.60 10.39% 5.09 2.60 1.81
Overall 8.93% 8.51 10.95% 7.87 12.51% 8.83 2.40 3.00
Table A5: Cumulative Abnormal Returns with Alternative Windows
This table reports the $R^{2}_{\mathrm{OOS}}$ of using analyst report text embeddings to estimate $CAR$ with alternative windows. The $R^{2}_{\mathrm{OOS}}$ is calculated annually using a training sample from 2000 to the preceding year. $CAR$ measures are centered on the release day of analyst reports ($T_{0}$). Panel A shows results for the full sample, while Panels B and C focus on periods within and beyond 1-day earnings announcement windows, respectively. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
AR[0] CAR[0,+1] CAR[-2,+2]
year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
Panel A: Full sample period
2015 8.03% 8.53 8.24% 8.36 10.87% 9.40
2016 7.88% 8.27 6.09% 4.68 6.60% 3.11
2017 8.75% 14.35 8.56% 9.29 9.00% 8.78
2018 9.09% 14.95 9.39% 11.26 9.75% 10.54
2019 9.89% 99.70 7.68% 9.45 8.66% 18.93
2020 2.33% 2.33 1.76% 3.15 4.13% 2.14
2021 5.75% 7.32 5.80% 6.62 4.05% 3.57
2022 10.54% 12.32 11.23% 18.20 10.91% 13.65
2023 5.38% 15.44 3.39% 7.00 0.53% 1.48
Overall 7.55% 8.07 6.92% 6.77 7.57% 6.79
Panel B: Earnings announcement period
Overall 9.78% 5.58 7.87% 5.10 13.14% 8.02
Panel C: Non-earnings announcement period
Overall 5.77% 5.78 5.62% 5.64 4.96% 4.91
Table A6: Information Content of Analyst Reports across Industries
This table reports the $R^{2}_{\mathrm{OOS}}$ of Fama-French 12 industries, using the definition from Kenneth French’s website. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
year Shops Other Manuf Chems
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 13.27% 6.03 657 17.69% 5.03 585 10.84% 3.19 808 15.05% 1.74 147
2016 10.43% 2.46 556 3.34% 0.48 593 10.37% 5.38 811 23.91% 3.25 111
2017 16.65% 3.49 506 20.45% 2.67 544 25.03% 13.64 722 11.04% 0.59 82
2018 10.86% 4.32 406 21.65% 3.40 405 3.32% 1.55 553 25.80% 1.87 77
2019 21.04% 9.24 394 7.49% 2.37 610 16.25% 12.07 568 5.00% 3.22 106
2020 4.39% 2.55 499 -0.30% -0.53 499 4.06% 1.41 542 2.65% 1.60 115
2021 10.27% 2.13 312 9.70% 3.66 362 8.66% 4.32 367 1.46% 0.30 102
2022 22.75% 3.83 320 15.98% 2.66 437 14.26% 11.46 420 7.00% 2.10 110
2023 17.78% 1.76 353 3.49% 2.00 339 7.15% 2.71 360 19.31% 6.54 76
Overall 13.88% 6.03 4003 11.27% 3.69 4374 11.39% 5.57 5151 11.35% 3.55 926
year Durbl BusEq Hlth NoDur
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 16.83% 5.77 216 15.66% 6.41 1378 5.75% 2.72 1242 0.68% 0.31 241
2016 4.95% 0.79 209 10.24% 6.09 1303 11.07% 2.90 1185 17.14% 5.18 242
2017 -2.17% -0.30 282 10.36% 3.12 1154 3.37% 2.39 1060 -2.14% -0.68 278
2018 18.29% 5.57 208 0.32% 0.31 1002 12.47% 6.58 918 1.29% 0.13 324
2019 15.31% 2.26 222 10.17% 6.13 1036 10.02% 7.68 884 18.51% 2.30 277
2020 5.73% 0.79 168 8.11% 4.14 1100 2.86% 1.84 950 -0.37% -0.28 265
2021 10.01% 5.31 150 11.43% 1.49 731 7.66% 4.58 592 -6.38% -0.60 195
2022 7.99% 2.58 162 12.08% 4.77 822 12.05% 2.71 850 5.24% 1.11 187
2023 12.05% 2.37 140 10.45% 11.80 962 3.52% 1.32 844 -5.69% -0.71 158
Overall 11.02% 3.24 1757 9.68% 6.77 9488 7.53% 5.71 8525 7.50% 1.38 2167
year Money Telcm Enrgy Utils
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 18.13% 4.41 1153 1.68% 0.00 403 -4.21% -1.45 303 18.47% 3.69 401
2016 7.64% 3.53 1108 -10.18% -1.22 329 15.80% 2.83 252 -1.84% -0.23 310
2017 1.77% 0.46 1071 8.60% 5.16 307 -18.18% -8.75 307 -39.46% -6.99 315
2018 13.55% 8.48 879 10.65% 2.46 219 8.38% 1.29 275 -23.76% -5.53 215
2019 2.10% 2.25 977 1.53% 0.52 351 -10.28% -9.07 340 -2.36% -0.27 193
2020 1.42% 0.76 952 -8.84% -2.83 360 5.72% 2.62 434 -7.29% -2.36 165
2021 2.62% 0.47 614 6.15% 3.27 188 -15.04% -6.83 251 -4.01% -1.16 103
2022 4.34% 2.13 667 12.84% 3.50 143 16.14% 4.44 286 8.31% 2.71 132
2023 -0.63% -0.18 569 8.99% 6.89 182 -27.33% -5.03 290 -6.43% -1.12 120
Overall 5.18% 3.31 7990 3.32% 1.39 2482 1.90% 0.31 2738 -1.97% -0.88 1954
Table A7: Summary Statistics for Alternative Analyst Information Value Estimations
The analyst information value is measured as the explained return variance divided by the price impact, with the results reported in millions of dollars. All dollar values are adjusted for inflation to reflect 2020 values using the Consumer Price Index (CPI). The sample comprises analyst reports targeting common constituent firms of the S&P 100 index from 2015Q1 to 2023Q4. Both mean and standard deviation are computed using the delta method.TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. Intraday denotes that $\sigma_{v}^{2}$ does not include overnight return volatility. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Mean SE 95%CI 99%CI N
Information value (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17669
Information value of text (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17672
Information value of revisions (TAQ volitility), \$M 0.42 0.04 [0.34, 0.50] [0.31, 0.53] 17672
Information value (Intraday), \$M 0.42 0.04 [0.34, 0.51] [0.32, 0.53] 17672
Information value of text (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17669
Information value of revisions (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17672
Information value (EMO), \$M 0.34 0.04 [0.27, 0.41] [0.25, 0.43] 17672
Information value of text (EMO), \$M 0.35 0.04 [0.27, 0.42] [0.25, 0.44] 17672
Information value of revisions (EMO), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17669
Information value (CLNV), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17672
Information value of text (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.41] 17672
Information value of revisions (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.42] 17672
C. Machine Learning Models
Ridge Regression
Ridge regression addresses multicollinearity by adding a regularization term to the least squares objective function. The ridge regression estimator is given by:
[…]
where $\alpha$ is the regularization parameter that controls the trade-off between fitting the data and shrinking the coefficients.
To find the optimal value of $\alpha$, cross-validation is used over a grid of values ranging from $10^{-10}$ to $10^{10}$. The cross-validation process ensures that the chosen model generalizes well to unseen data, preventing overfitting while capturing the predictive power of the text embeddings.
Partial Least Square Regression
To mitigate the risk of overfitting inherent in high-dimensional text embeddings, I employ Partial Least Squares (PLS) for dimensionality reduction.
The optimization problem can be expressed as follows:
[…]
where $\Omega$ is a $K\times P$ transformation matrix that reduces the $K$ predictors in $X_{i}$ to $P$ lower-dimensional components.
The extraction of the $j$-th PLS component is guided by the following objective function:
[…]
In essence, this approach sequentially extracts components that maximize the covariance with the outcome variable, while ensuring orthogonality to previously extracted components.
Extreme Gradient Boosting
Tree-based approaches are commonly applied in stock return forecasting literature (see, e.g., Gu et al., 2020; Cao et al., 2024; Bonini et al., 2023). XGBoost is an advanced implementation of tree-based machine learning models that builds an ensemble of decision trees. In XGBoost, each tree is added sequentially to correct the errors of previous trees. The main idea is to combine the outputs of multiple weak learners (decision trees) to create a strong learner. XGBoost incorporates regularization techniques, such as L1 (Lasso) and L2 (Ridge), to prevent overfitting and enhance model generalization.
In comparison to random forests, which build a multitude of independent trees and aggregate their predictions, XGBoost constructs trees sequentially, with each tree designed to correct the errors of the preceding ones. While random forests rely on bagging, a method that combines the predictions of various trees to reduce variance, XGBoost uses boosting, an approach that aims to reduce both bias and variance by focusing on difficult-to-predict instances in subsequent iterations. This boosting approach allows XGBoost to effectively capture and leverage the nuanced information embedded in textual data, leading to a more accurate estimation of stock returns.
Neural Networks
To extend beyond the linear modeling approach, I explore the use of Neural Networks to estimate $CAR$ using text embeddings derived from analyst reports. A Neural Network can capture complex non-linear relationships between the text embeddings and the $CAR$, potentially improving the model’s fitting capabilities. Consider a three-layer Neural Network as the prediction model. The prediction problem can be formulated as follows:
$f(X_{i};\theta)=W_{3}\sigma\left(W_{2}\sigma\left(W_{1}X_{i}+b_{1}\right)+b_{2}\right)+b_{3},$ (16)
where $\sigma(\cdot)$ is the ReLU activation function, $W_{i}$ and $b_{i}$ represent the weights and biases for layer $i$, respectively. The Neural Network architecture consists of an input layer representing the text embeddings, followed by multiple hidden layers. The specific architecture employed includes 32 neurons in the first hidden layer, followed by 16, 8, 4, and 2 neurons in subsequent layers. This structure is flexible and can be adjusted by adding or removing layers as necessary to optimize performance.
The training process involves optimizing the weights and biases to minimize the loss function, which, in this case, is the mean-square loss. To regularize the model and prevent overfitting, early stopping is implemented, halting training once the validation loss ceases to decrease. Additionally, following Gu et al. (2020), the models are retrained five times, and the final estimation is obtained by averaging the outputs of these five models, forming an ensemble estimation.
D. Theory
I provide an intuition for the measure of strategic value by discussing a simple extension of Kyle (1985) model. In the Kyle model, there is one risky asset with a payoff $\tilde{v}\sim N\left(p_{0},\Sigma_{0}\right)$. Three types of traders exist: a strategic trader with insider information, a market maker who sets prices in a perfectly competitive market, and an uninformed trader who trades $\tilde{u}\sim N\left(0,\sigma_{u}^{2}\right)$, where $\tilde{u}$ is independent of $\tilde{v}$. Illiquidity is measured by Kyle’s lambda ($\lambda$). Kyle’s lambda depends on private information and liquidity trading. I extend the model by considering a case where $\phi$ percentage of variance in $\tilde{v}$ is explained by the informed trader’s information.
A Single Auction Model
There are two periods, $t_{0}$ and $t_{1}$. The asset is traded with asymmetric information at period $t_{0}$, and the value $\tilde{v}$ is realized at period $t_{1}$. I assume without loss of generality that
$\tilde{v}=P_{0}+\tilde{s}+\tilde{\epsilon},$
where $\tilde{s}$ is a mean-zero signal observed by the informed trader at $t_{0}$, and $\tilde{\epsilon}$ is the combination of residual information and noise. I assume that the signal $\tilde{s}$ and residual $\tilde{\epsilon}$ are uncorrelated, that is, $\sigma_{v}^{2}=\sigma_{s}^{2}+\sigma_{\epsilon}^{2}$.
Let
$\phi=\frac{\operatorname{var}(\tilde{s})}{\operatorname{var}\left(\tilde{v}\right)}=\frac{\sigma_{s}^{2}}{\sigma_{v}^{2}}.$
This measure $\phi$ is the ”R-square” of projecting $\tilde{v}$ on $\tilde{s}$, or the percentage of explainable variance in $\tilde{v}$ using signal $\tilde{s}$.
After observing $\tilde{s}$, the informed trader submits a market order $\tilde{x}=X(\tilde{s})$, and the uninformed trader trades a zero-mean random variable $\tilde{z}$ that is normally distributed and independent of $\tilde{v}$. The profit of the informed trader is given by: $\tilde{\pi}=(\tilde{v}-\tilde{p})\tilde{v}$. The insider has a rational guess of $P(\tilde{x}+\tilde{u})$ and understands that his order $\tilde{x}$ will move the price against him.
The market maker observes the order flow $\tilde{y}\stackrel{{\scriptstyle\text{def}}}{{=}}\tilde{x}+\tilde{z}$ and determines the equilibrium price $\tilde{p}=P(\tilde{x}+\tilde{u})$ to break even. The assumptions for the market maker are risk neutrality and perfect competition, which drives the profits for market makers to zero.
Equilibrium
An equilibrium is a set of $X$ and $P$ satisfying
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s],$
$\tilde{p}(X,P)=E[\tilde{v}\mid\tilde{x}+\tilde{u}].$
The first condition is profit maximization, stating that given the market maker’s pricing rule, the insider chooses a strategy $X$ maximizing her conditional expected profit, taking into account the pricing rule. The second condition is market efficiency. Given the insiders’ trading strategy, the market maker sets the price to be the expected value of the security.
Conjecture:
$P(\tilde{y})=\mu+\lambda\tilde{y},$
$X(\tilde{s})=\alpha+\beta\tilde{s}.$
The profit of the insider can be written as:
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s]$ $=E[(\tilde{v}-\mu-\lambda(\tilde{u}+x))x\mid\tilde{s}=s]$
$=(P_{0}+s+-\mu-\lambda x)x.$
Traders take into account that her order flow will move the price against her, which serves to restrain her position size.
Solving for optimal profit, I get:
$x^{*}=\frac{P_{0}+s-\mu}{2\lambda}=\alpha+\beta v.$
Hence, I can express $\alpha$ and $\beta$ :
$\begin{gathered}\beta=\frac{1}{2\lambda},\\ \alpha=\frac{P_{0}-\mu}{2\lambda}=(P_{0}-\mu)\beta.\end{gathered}$
When the market maker puts a higher weight on the order flow in setting the price, the trader puts a lower weight on her information.
I now look at the price-setting rule:
$\mu+\lambda y=E\{\tilde{v}\mid\alpha+\beta\tilde{s}+\tilde{u}=y\}.$
Essentially, the market maker observes a normally distributed signal about
$\mu+\lambda y$ $=E[\tilde{v}\mid y]$
$=\bar{v}+\frac{\operatorname{cov}(\tilde{v},\alpha+\beta\tilde{s}+\tilde{u})}{\operatorname{var}(\alpha+\beta\tilde{s}+\tilde{u})}(\beta\tilde{s}-\beta\bar{s}+\tilde{u})$
$=p_{0}+\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}}(y-\alpha).$
Hence, I can express $\lambda$ and $\mu$ :
$\lambda=\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}},$
$\mu=p_{0}-\alpha\lambda.$
There is a unique linear equilibrium given by
$\mu=P_{0},$
$\lambda=\frac{\sigma_{s}}{2\sigma_{u}},$
$\alpha=0,$
$\beta=\frac{\sigma_{u}}{\sigma_{s}}.$
Discussion
Kyle’s Lambda
The parameter $\lambda$ is universally known as Kyle’s lambda. Formally, it is the impact on the equilibrium price of a unit order. Its reciprocal ($1/\lambda$) measures the liquidity (or depth) of the market. If $1/\lambda=\frac{2\sigma_{u}}{\sigma_{s}}$ is larger, then the market is more liquid, either because there is less private information in $\sigma_{s}$ or there is more liquid trade in the sense of $\sigma_{u}$.
Value of Private Information
Notice that the equilibrium strategy of the informed trader is $\tilde{x}=\beta s$. The unconditional expected gain of the informed trader is
$\mathrm{E}[\tilde{x}[\tilde{v}-p(\tilde{x}+\tilde{z})]]$ $=\beta\mathrm{E}[s(\tilde{v}-\mu-\lambda\beta s-\lambda\tilde{u})]$
$=\beta(1-\lambda\beta)\sigma_{s}^{2}$
$=\frac{\sigma_{s}\sigma_{u}}{2}$
$=\frac{\phi\sigma_{v}^{2}}{4\lambda}.$
The expected gain for the informed trader is maximized when she has more private information or when there is more liquidity trading. On the other hand, liquidity traders incur losses equivalent to the informed trader’s gains, but they accept these losses due to other motives for trading. The equilibrium price ensures that market makers do not profit or lose in expectation.
E. Delta Method Approximation
I use the delta method to approximate the expected value and variance of a ratio of two random variables. This method relies on a first-order Taylor expansion. Specifically, consider two random variables $X$ and $Y$ with means $\mu_{X}$ and $\mu_{Y}$, and variances $\sigma_{X}^{2}$ and $\sigma_{Y}^{2}$, respectively. We are interested in the ratio $Z=\frac{X}{Y}$ and want to approximate the mean $E[Z]$ and the variance $\text{Var}(Z)$.
Delta Method Approximation
The key idea is to approximate the function $g(X,Y)=\frac{X}{Y}$ using a Taylor series expansion around the means $\mu_{X}$ and $\mu_{Y}$. For the function $g(X,Y)$, I use a first-order Taylor expansion around $(\mu_{X},\mu_{Y})$:
$g(X,Y)\approx g(\mu_{X},\mu_{Y})+\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}(X-\mu_{X})+\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}(Y-\mu_{Y}).$
The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are:
$\frac{\partial g}{\partial X}=\frac{1}{Y},\quad\frac{\partial g}{\partial Y}=-\frac{X}{Y^{2}}.$
Evaluating these at $(\mu_{X},\mu_{Y})$, I get:
$\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}=\frac{1}{\mu_{Y}},\quad\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}=-\frac{\mu_{X}}{\mu_{Y}^{2}}.$
Substituting the partial derivatives into the Taylor expansion, I obtain:
$\frac{X}{Y}\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}(X-\mu_{X})-\frac{\mu_{X}}{\mu_{Y}^{2}}(Y-\mu_{Y}).$
Taking the expectation on both sides:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}E[X-\mu_{X}]-\frac{\mu_{X}}{\mu_{Y}^{2}}E[Y-\mu_{Y}].$
Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation simplifies to:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}.$
Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by:
$\operatorname{Var}(Z)\approx\left(\frac{\partial g}{\partial X}\right)^{2}\operatorname{Var}(X)+\left(\frac{\partial g}{\partial Y}\right)^{2}\operatorname{Var}(Y)+2\left(\frac{\partial g}{\partial X}\right)\left(\frac{\partial g}{\partial Y}\right)\operatorname{Cov}(X,Y).$
Substituting the partial derivatives:
$\operatorname{Var}\left(\frac{X}{Y}\right)\approx\frac{\operatorname{Var}(X)}{\mu_{Y}^{2}}+\frac{\mu_{X}^{2}\operatorname{Var}(Y)}{\mu_{Y}^{4}}-2\frac{\mu_{X}\operatorname{Cov}(X,Y)}{\mu_{Y}^{3}}.$
Application to Value of Information
In the context of the provided study, I estimate the mean value of information for a subsample $s$ using the delta method.
Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$. Define the mean and variance over subsample $s$ as:
$\mu_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right),$
and the mean price impact per dollar over subsample $s$ as:
$\mu_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}.$
Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by:
$E\widehat{\Omega}_{s}\approx\frac{\mu_{\nu s}}{\mu_{\lambda s}}.$
The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as:
$\text{Var}\left(\widehat{\Omega}_{s}\right)\approx\frac{1}{\mu_{\lambda s}^{2}}\left(\Sigma_{\nu s}+\frac{\mu_{\nu s}^{2}}{\mu_{\lambda s}^{2}}\Sigma_{\lambda s}-2\frac{\mu_{\nu s}}{\mu_{\lambda s}}\Sigma_{\nu\lambda s}\right),$
where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as:
$\Sigma_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(\widehat{\sigma}_{it}^{2}-\mu_{\nu s}\right)^{2},$
$\Sigma_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\left(\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}-\mu_{\lambda s}\right)^{2},$
$\Sigma_{\nu\lambda s}=\text{Cov}\left(\widehat{\sigma}_{it}^{2},\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}\right).$
This provides a first-order approximation for the mean and variance of the value of information over the subsample $s$ using the delta method.