### v2

# Linying Lv

Abstract
I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports. Using embeddings from state-of-the-art large language models, I show that textual information in analyst reports explains 10.19% of contemporaneous stock returns out-of-sample, a value that is economically more significant than quantitative forecasts. I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns. The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power. Expressing these findings in economic terms, I estimate that early acquisition of analysts’ reports can yield significant profits. Analysts’ information value peeks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data. JEL Classification: G11, G14, G24
Keywords: Sell-side Analysts, Value of Information, Large Language Models, Explainable AI
“No one ever made a decision because of a number. They need a story.”
― Daniel Kahneman
1 Introduction
Sell-side analysts play a crucial role in financial markets by processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond simple sentiment dimension and the context has been perceived at least as important as the quantifiable outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024). For example, consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences express negative sentiment and pertain to the “earnings analysis” topic, the latter provides additional contextual information and reasoning that can be interpreted and analyzed by LLMs.
I quantify the value of analyst information in analyst reports through three general steps: analyst output representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts from I/B/E/S, such as stock recommendations, earnings forecasts, and target prices. For structured representation of qualitative textual information, I extract contextualized representations of analyst report texts using state-of-the-art LLMs. Each report is mapped into a structured embedding space using the LLMs’ transformer architecture.
With structured text representation, I evaluate how well analyst information explains stock returns using an R-squared measure, which also serves as the terminology of information content. A significant challenge in projecting stock returns onto analyst outputs is the risk of in-sample overfitting due to the high-dimensionality of text embeddings. To address this issue, I employ Ridge regression, which introduces a regularization term to penalize large coefficients. This regularization helps in handling the high-dimensional embeddings while preserving the model’s ability to capture essential relationships between analyst outputs and stock returns. Additionally, I evaluate the out-of-sample explanatory power of quantitative and qualitative measures for abnormal stock returns to assess model stability.
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using each type of information individually. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the substantial and unique information content within analyst report text.
The informativeness of analyst reports varies considerably depending on whether they are associated with forecast revisions or reiterations. Both qualitative and quantitative information provide greater value to the financial market when linked to forecast revisions. Specifically, when analyst reports involve recommendation revisions, the combined use of numerical and textual data yields an out-of-sample $R^{2}$ of 22.63%. In stark contrast, for reports with earnings forecast reiterations, the out-of-sample $R^{2}$ is not statistically significant, suggesting that such reports contribute minimal incremental information to the market.
A notable amount of research emphasizes the role of analysts in information processing around corporate earnings announcements, e.g., Livnat and Zhang (2012), Keskek et al. (2014), Kim and Song (2015), Lobo et al. (2017), and Barron et al. (2017), An investigation on samples within and beyond the earnings announcement windows reveals that the information content of analyst reports peaks in the first week following earnings announcements. The finding supports Huang et al. (2018)’s perspective that timely reactions to corporate disclosures offer valuable information advantages to clients. Using separately trained models on samples within and beyond earnings announcement windows, the magnitude of the out-of-sample $R^{2}$ nearly doubles in the first week following earnings announcement dates compared to other periods. The significant increase underscores the incremental value that analysts add by interpreting and contextualizing earnings information for investors.
Analysts might simply be “piggybacking” on information disclosed during earnings conference calls, I incorporate controls for both the earnings surprise and the text embeddings of the latest earnings conference call transcripts into models to isolate the unique information from analyst reports. The findings indicate that even after accounting for the information from conference calls, analyst reports contain distinct and valuable insights.
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains an open question. To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results show that over half of the explanatory power of the report, in terms of out-of-sample $R^{2}$, is attributed to Income Statement Analyses in analyst reports, particularly the interpretation of realized income numbers. This evidence again emphasizes the critical role analysts play in interpreting financial data and providing the stories and context needed for making investment decisions.
So far, I have measured information value statistically with R-squared. To translate it into economic terms, I adopt Kadan and Manela (forthcoming)’s framework to quantify the dollar value of analyst information to strategic investors. The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts. This multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to the price impact, representing the cost associated with trading on this information.
The findings reveal significant value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports are \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million from both combined. The annualized aggregate information value is estimated at \$6.89 million, based on an average of 15 days with reports released per year. According to Kadan and Manela (forthcoming), this estimate represents a conservative lower bound. Notably, these values increase for large stocks, bold analysts, and reports released promptly following earnings announcements.
My paper relates to three strands of literature. This paper first relates to recent studies on the nature of information conveyed by analysts through their qualitative outputs (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Prior research often relies on manually encoding textual information and sentiment analyses, which is labor-intensive and may overlook nuanced insights. My work provides evidence for the superiority of state-of-the-art LLMs in generating contextualized textual representations. A direct comparison with the Naive Bayes tone classification used in Huang et al. (2014) reveals two main advantages of LLMs: they more effectively capture overall sentiment in analyst reports and extract richer information beyond tone. Furthermore, transitioning from tone analysis alone to a full-text representation using LLM embeddings further boosts the $R^{2}$ measure from 3% to above 10%. The significant improvement in explanatory power demonstrates the potential of LLMs to revolutionize the analysis of textual information. Topic decomposition analysis offers a new dimension for understanding analyst information value by revealing their role in interpreting financial data.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2024), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in financial contexts. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. This research leverages the additive nature of LLM embeddings and proposes a Shapley value decomposition approach for interpretability, offering a new framework for explainable AI.
Third, this paper deepens the understanding of the economics of tipping - a well-documented phenomenon between analysts and institutional investors (Irvine et al., 2007; Christophe et al., 2010). Tipping involves brokerage firms rewarding high-commission clients with early research access, which allows these firms to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends prior work by quantifying the economic dollar value of such insider information.
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation. Section 3 summarizes data construction and empirical results. Section 4 concludes.
2 Methodology
In this section, I introduce how I employ large language models for topic modeling and text embeddings, the information content estimation, the strategic information value estimation, and the decomposition metric used to evaluate the information value of topics. [^1] The details of machine learning models used to explain contemporaneous stock returns are discussed in the Online Appendix C.
2.1 Large Language Models
I harness the power of advanced generative AI models to conduct comprehensive analyses of analyst reports. The methodology employs a multi-faceted approach: fine-tuned BERT models for precise sentence classification based on predefined criteria, ChatGPT prompting for extracting and synthesizing major topics discussed by analysts, and LLaMA embedding for representing entire reports as structured high-dimensional vectors.
2.1.1 ChatGPT Prompting
To avoid ad-hoc classification, I employ ChatGPT prompting to extract exclusive and meaningful topics. The goal is to identify primary topics covered in analyst reports and assign each sentence to a single topic category.
In the first step, I use a random sample of 100 analyst reports and request ChatGPT to analyze the content and identify high-level categories with the following prompt:
Prompt 1:Please read the provided text file of sell-side analyst reports carefully. What are high-level mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
It ends up with 13 categories defined as follows:
  • Executive Summary: Provides a high-level overview of the report’s key findings and conclusions; includes a brief description of the company, its industry, and the purpose of the report; highlights the most important points from the analysis, such as the company’s financial performance, competitive position, and growth prospects.
  • Company Overview: Offers a comprehensive description of the company, including its history, business model, and key products or services; discusses the company’s organizational structure, management team, and corporate governance; analyzes the company’s mission, vision, and strategic objectives.
  • Industry Analysis: Provides an in-depth analysis of the industry in which the company operates; includes information on market size, growth trends, and key drivers; discusses the regulatory environment, technological advancements, and other external factors affecting the industry; analyzes the industry’s competitive dynamics and the company’s position within the industry.
  • Competitive Landscape: Identifies the company’s main competitors and their market share; compares the company’s products, services, and pricing strategies with those of its competitors; analyzes the strengths and weaknesses of the company and its competitors; discusses potential new entrants and substitutes that could disrupt the competitive landscape.
  • Financial Analysis: Analyzes the company’s revenue, expenses, profitability, assets, liabilities, shareholders’ equity, liquidity, solvency, capital structure, cash flows, and key financial ratios; compares the company’s financial performance with that of its competitors and industry benchmarks; discusses trends in the company’s financial performance over time.
  • Business Segments: Provides a detailed analysis of the company’s various business segments or divisions; discusses the financial performance, growth prospects, and challenges of each segment; analyzes the contribution of each segment to the company’s overall revenue and profitability.
  • Growth Strategies: Discusses the company’s strategies for driving future growth, such as organic growth initiatives, product innovations, and geographic expansions; analyzes the company’s mergers and acquisitions (M&A) strategy and potential targets; examines the company’s investments in research and development (R&D) and marketing.
  • Risk Factors: Identifies and analyzes the key risks facing the company, such as market risks, operational risks, financial risks, and legal/regulatory risks; discusses the potential impact of these risks on the company’s financial performance and growth prospects; examines the company’s risk management strategies and mitigation measures.
  • Management and Governance: Provides an overview of the company’s management team, including their experience, expertise, and track record; analyzes the company’s corporate governance practices, such as board composition, executive compensation, and shareholder rights; discusses the company’s succession planning and key person risks.
  • Environmental, Social, and Governance (ESG) Factors: Analyzes the company’s performance and initiatives related to environmental sustainability, social responsibility, and corporate governance; discusses the potential impact of ESG factors on the company’s reputation, risk profile, and financial performance; examines the company’s compliance with ESG regulations and industry standards.
  • Valuation: Estimates the intrinsic value of the company’s shares using various valuation methodologies, such as discounted cash flow (DCF) analysis, relative valuation multiples, and sum-of-the-parts analysis; compares the company’s valuation with that of its peers and historical benchmarks; discusses the key assumptions and sensitivities underlying the valuation analyses.
  • Investment Thesis: Summarizes the key reasons for investing (or not investing) in the company’s shares; discusses the potential catalysts and risks that could impact the company’s valuation and stock price performance; provides a target price or price range for the company’s shares based on the valuation analyses and investment thesis.
  • Appendices and Disclosures: Includes additional supporting materials, such as financial statements, ratio calculations, and detailed segment data; provides important disclosures, such as the analyst’s rating system, potential conflicts of interest, and disclaimers; discusses the limitations and uncertainties of the analysis and the need for further due diligence by investors.
The Financial Analyses category is further divided into four subcategories: Income Statement Analyses, Balance Sheet Analyses, Cash Flow Analyses, and Financial Ratios. Additionally, a category labeled “None of the Above” is included to accommodate sentences that do not fit into any of the predefined topics. In total, this process generates 17 distinct topics for classification. The distribution of topics is relatively stable across years, as illustrated in Figure 1. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, I employ the GPT-4-Turbo model to categorize each sentence in a random sample of 17,028 sentences from 100 analyst reports into a single, most relevant topic.
Prompt 2: Please read the following sentence from a sell-side analyst report of the company {firm} ({ticker}) carefully. Determine which category of information it belongs to in the following 17 categories: {categories}. Pay attention to the sentence in the context of the report.
Output your response in JSON format.
Here is the sentence from the analyst report: {sentence}.
This training sample serves as the training sample for fine-tuning a BERT model, which is subsequently used to categorize all sentence segments in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least prevalent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. The scarcity of Executive Summary content may result from the difficulty of identifying this category at the sentence level. Notably, ESG-related content, while minimal overall, appears mainly after 2020, reflecting the recent rise in ESG discussions.
2.1.2 Text Representations
A systematic textual analysis requires a structured representation of analyst reports. Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. Within LLMs, tokens are mapped to unique identifiers (IDs) based on a predefined vocabulary. This vocabulary is learned during the model’s training phase and contains all the tokens the model can understand.
To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”. This granular approach allows for nuanced interpretation of complex terms. For instance, ‘next-gen workloads’ becomes five separate tokens: ‘next’, ‘-’, ‘gen’, ‘work’, and ‘loads’.
At the heart of LLMs lies the transformer architecture. The transformer’s encoder-decoder structure, combined with multi-head self-attention and feed-forward networks, enables efficient processing of large datasets and state-of-the-art performance. A core output of the transformer architecture is text embeddings. These embeddings are dense, continuous vectors that capture semantic and syntactic information. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
To obtain structured representations of analyst reports, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens. The tokenizer processes the text data, converting it into tokens that the model can understand. By processing the analyst reports through this model, I obtained rich, contextual embeddings that capture the nuanced information within the reports. These embeddings are suitable for various downstream NLP tasks, such as sentiment analysis and economic modeling.
2.2 Information Content
Following Asquith et al. (2005), I use $R^{2}$ as a measure for information content. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations each year with an expanding training window:
$CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (1)
[…]
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ is the structured representation of information from analyst report $j$ on stock $i$ released at day $t$. I use the $60\%$ of data as the initial training sample. Therefore, the out-of-sampel test starts in 2015 and ends in 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ as
$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum_{(i,t)\in\mathcal{T}}CAR_{[-1,+1],it}^{2}},$ (3)
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the portion of return variance that can be explained by analysts’ information.
To evaluate the statistical significance of the $R_{\mathrm{oos}}^{2}$, I employ a modified version of the Diebold and Mariano (2002) test for comparing predictive accuracy between two models. Given the strong cross-sectional dependence in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional average of prediction errors from each model rather than individual returns.
2.3 Information Value
I measure the value of information using equation (4) following Kadan and Manela (forthcoming) who show that it approximates the value of information in a strategic setting with multiple informed investors, each possessing an imperfect signal:
$\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
where $\widehat{\Omega}_{it}$ is the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price of day t-2. $\widehat{r}_{ijt}$ is the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ is the realized $CAR_{[-1,+1]}$, with abnormal returns calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ denotes the number of analyst reports about stock $i$ on day $t$.
To estimate the price impact $\widehat{\lambda}_{it}$, I perform a regression of 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). Here, $y_{itk}$ is calculated as the change in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals. The signed order flow is classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^2] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression model is
$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda). In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The 3-day interval $T$ is divided into one-minute sections $\tau_{0},\tau_{1},\ldots,\tau_{K}$. Thus, each regression has $1170(=390*3)$ observations.
Essentially, the measure reflects the ratio of return volatility to price impact. It is derived based on the theoretical framework of Back et al. (2000), which is a continuous-time version of Kyle (1985) model. Due to risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of the asset price. The intuition is illustrated with an extended single-auction Kyle model incorporating incomplete information in the Online Appendix D.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance. [^3] Kadan and Manela (forthcoming) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply to the analyst case.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To tackle this issue, I apply the delta method to estimate the mean and variance of information value over sample $s$. The delta method used to derive the first-order estimator is discussed in the Online Appendix E.
2.4 Shapley Value Decomposation
To assess the impact of topics within analyst reports, I employ the Shapley Additive exPlanations (SHAP) methodology. SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. It was later applied for machine learning model interpretation by Lundberg and Lee (2017) and Chen et al. (2022). In my context, the approach allows me to attribute the overall explanatory power of an analyst report to its constituent topics. Leveraging the additive nature of text embeddings, I design a novel decomposition method to systematically assess the relative importance of various topics in analyst reports.
LLMs pool report-level embeddings by averaging embeddings across all tokens and all layers. The textual embedding can be represented as:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ is the total number of words in the report. If the tokens can be classified into different topics, the embedding can be fully decomposed into different topic-specific embeddings:
$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
where $P$ is the number of topics, $y_{p}^{emb}$ is the embedding for topic $p$, $N_{p}$ is the number of tokens belonging to topic $p$, and $i_{p}$ is the index of tokens within topic $p$.
One concern raised by the nature of transformer models is that each token in the input sequence can pay attention to other tokens through a mechanism called “self-attention” or “scaled dot-product attention.” This architecture allows the model to weigh the importance of each token relative to every other token and capture contextual information. For example, the embedding of token $i$ contains information from token $j$, which may pertain to a different topic. To isolate the self-attention across topics, I process the embeddings sentence by sentence to ensure the token embeddings only reflect the context information within the sentence. I then take the token-weighted mean of sentence embeddings as the report representation, which is referred to as sentence-segment embeddings.
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embedding of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding reflects the information content of context across sentences.
Second, I calculate the Shapley value of the sentence-segmented embeddings distributed to each topic. Specifically, the Shapley value can be written as:
$\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
where the sum extends over all subsets $S$ of $P$ not containing topic $p$, including the empty set. Note that $\binom{n}{k}$ is the binomial coefficient. $y_{p}^{\text{emb}}$ is the token-weighted sentence embedding of topic $p$, and $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. For each combination of topic embeddings, I calculate the $R_{\text{oos}}^{2}$ accordingly. [^4] Shapley value requires an input for $X$ when it is missing. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ among topics, revealing their relative importance in explaining market reactions.
3 Data and Empirical Results
This section presents the empirical analysis of analyst report information. I begin by examining the information content and decomposing it across different topics. I then quantify the strategic dollar value of analyst information. The results emphasize analysts’ income statement analyses following earning announcements.
3.1 Data
Sell-side analyst reports are obtained from the Mergent Investext Database. I download 223,091 analyst reports of S&P 100 firms from 2000 to 2023. The initial reports are in PDF format, and I use Adobe Acrobat to extract plain text from each report. The text is chunked into sentences. To get clean content of firm-relevant discussions, I fine-tune a BERT model to remove the boilerplate segments. [^5] Boilerplate segments disclose information about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I randomly sampled one report from each of the top 20 brokerage firms producing the largest number of analyst reports and manually labeled each sentence in those reports. The BERT model fine-tuned for boilerplate classification reached an accuracy score of 97.31% on a testing sample.
The Mergent Investext Database provides information on analyst names and the release date of each report. To connect Investext with the I/B/E/S database, I first match the lead analyst name in reports to Analyst ID (AMASKCD) following the procedure of Li et al. (2024). I then match specific reports to the corresponding EPS forecasts, price targets, and recommendation files in the I/B/E/S database. Since multiple reports could be released between a forecast announcement and its revision, I follow Huang et al. (2014) and define the matching window for each report as 2 days before the announcement date to 2 days after the revision date. [^6] In I/B/E/S, each analyst forecast for a specific firm has two time stamps. The announcement date (ANNDATS) is the date the analyst announces the new forecast and revises the old one. The revision date (REVDATS) is the date the forecast becomes invalid or is revised. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. I also compile firm characteristics from the CRSP and Compustat databases. [^7] See Table A2 and Table A3 for detailed variable definitions and summary statistics. After the matching procedure and excluding analyst reports released on earnings announcement days, my final sample comprises 122,252 reports of S&P 100 stocks from 1,305 analysts across 140 unique brokerage firms.
Table 1 presents the summary statistics of the analyst report data, including the number of reports, the number of unique brokerage firms, and the number of unique analysts across years and the Fama-French 12 industries. It also summarizes the average number of pages and tokens per report, which are important indicators of report lengths. Overall, the time trend in the number of analyst reports aligns with that reported in Bonini et al. (2023), and the average number of pages per report is consistent with that in Huang et al. (2014). [^8] The data reveals a peak in the number of analyst reports in 2013, followed by a subtle decline. This trend can be attributed to the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and the introduction of MiFID II in Europe (Bradshaw et al., 2017).
The stock-level price impact is estimated using intraday data from the NYSE TAQ database, covering an out-of-sample period from January 2, 2015, to December 28, 2023. To ensure comparability over time, all prices are adjusted for inflation using the Consumer Price Index (CPI), with December 2020 as the base period.
Given my goal to estimate the information value of analyst outputs, I need to control for other information available to investors. In Section 3.2.2, I compare textual information from earnings conference calls and subsequent analyst reports. Earnings conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Available at https://seekingalpha.com/earnings/earnings-call-transcripts.. These transcripts provide comprehensive record of the calls, including executive presentations and manager-analyst interactions like Q&A sessions. I sample earnings conference call transcripts of S&P 100 constituent firms from 2005 to 2023, merging them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks. This process yielded a final sample of 7,909 firm-quarter-level conference calls.
The set of numerical variables used as input in machine learning models includes:
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ indicates target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $Boldness$ is an indicator variable that takes the value of 1 if revisions are either above both the consensus and previous forecasts or below both. $SR$ represents stock recommendation, following the I/B/E/S rating system, where a rating of 5 denotes a strong sell and a rating of 1 indicates a strong buy. $ERet$ represents the 12-month return forecast, calculated by scaling the 12-month price target by the stock price 50 days before the report release. $Prior\_CAR$ represents the cumulative abnormal return over time period [-10, -2].
Firm-Level Measures: $Size$ is measured by market equity. The Book-to-Market Ratio ($BtoM$) compares the firm’s book value to its market value. Following Huang et al. (2018), the Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ representing its absolute value. The $Miss$ variable is a binary indicator set to one if $SUE$ is positive. $TradingVolume$ is the stock’s trading volume on the latest earnings announcement days standardized by the number of shares outstanding. The Distance to Default ($DD$) variable is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market fluidity.
Industry-Level Measures: $IndustryRecession$ is an indicator variable equal to one if the firm’s industry (defined by Fama-French 48 classification) experiences a negative monthly return in the bottom quintile.
Macroeconomic Measures: The $TimeTrend$ variable captures the temporal dimension by recording the number of years elapsed since the beginning of the sample period.
3.2 Information Content of Analyst Reports
In this section, I examine the information content of analysts’ quantitative and qualitative outputs by projecting contemporaneous cumulative abnormal returns on analyst information. Furthermore, I attribute the information content of analyst report text across topics using the Shapley value decomposition introduced in Section 2.4.
3.2.1 Information Content of Quantitative and Qualitative Information
Table 2 compares the information content embedded in numerical and textual measures, quantified with $R^{2}_{\mathrm{OOS}}$. The estimation model applied in this table is the Ridge regression specified in Equations 1 and 2. The objective variable is ${CAR}_{[-1,+1]}$ centered at the analyst report release date. For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database. I estimate separate Ridge regression models for distinct input configurations and employ an expanding window approach, constructing training sets from 2000 to the year preceding each target year.
Panel A of the table delineates the performance of four types of input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% with forecast revision input, corroborating extant literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst forecast revisions. The inclusion of additional numerical variables in column (2) yields only a marginal improvement, with the $R^{2}_{\mathrm{OOS}}$ reaching 9.08% using 17 numerical features. I compare the estimation with a zero benchmark and report the Diebold-Mariano test t-statistics in columns (2) and (4). The models with revision and numerical inputs significantly outperform a zero-estimation benchmark.
Columns (5) and (6) elucidate the performance of Ridge models with text embeddings as input, showcasing a substantial $R^{2}_{\mathrm{OOS}}$ of 10.19% and a t-statistic of 8.20, underscoring the significant information content of the report text in explaining contemporaneous stock returns. The $R^{2}_{\mathrm{OOS}}$ exhibits relative stability across 2015-2023, with a notable exception during the 2020 pandemic recession. The DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, suggesting substantial information content in analyst report text compared with forecast revisions.
Sarkar and Vafa (2024) raise the look-ahead bias concern associated with pre-trained LLMs. Language model outputs may incorporate future information that should not be available. Since the LLMs I applied were trained on language data ending before 2023, the $R^{2}_{\mathrm{OOS}}$ in 2023 reflects true out-of-sample performance, confirming the robustness of the main findings during this period.
In columns (7) and (8), I further investigate the model performance when combining both forecast revisions and text embeddings in the Ridge regression. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, which represents a 3.27% improvement over the $R^{2}_{\mathrm{OOS}}$ of the revision-only model and a 2.09% improvement over the $R^{2}_{\mathrm{OOS}}$ of the text-only model. I report the DM t-statistic of the comparison between combined input with revision input and text input in columns (7)-(1) and (7)-(5). The overall t-statistics of 3.95 and 3.77 indicate a significant superiority in the combination of text data and revision measures in explaining contemporaneous $CAR$. The DM tests confirm that report text and forecast revision contain distinct information sets that are valued by investors.
The explanatory power might be derived from quantitative information in the numbers mentioned in report text. I conduct robustness checks using content with numbers removed. To address the concern, I exclude numerical data from the written reports before generating text embeddings and report model performance in Table A4. The $R^{2}_{\mathrm{OOS}}$ of the number-free text embedding input model is 10.95%, significantly higher than that of quantitative revision measures (t-statistic = 2.40). The robustness check suggests that the superior performance of text-based models is not driven by the numbers embedded in text.
LLMs exhibit improved language skills as the parameter size increases. I provide a quantitative comparison of LLMs with varying numbers of parameters, including the BERT model, OpenAI text-embedding-3-small model, and LLaMA-3 model. The LLaMA models have a substantially larger number of parameters compared to the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 model. The input in Panel B is report text embeddings, comparable to column (3) in Panel A. A positive t-statistic indicates that the model outperforms the LLaMA-2 benchmark. Overall, the BERT and OpenAI models underperform the LLaMA-2 model. The LLaMA-3 model gives an $R^{2}_{\mathrm{OOS}}$ of 9.66%, which is slightly better but statistically insignificant compared to the LLaMA-2 model. The results in Table 2 suggest that the out-of-sample information content of analyst report text is fairly robust across different LLMs. Pairwise comparisons between LLMs also shed light on how the scale of the model influences performance capabilities.
To test the robustness of the information content in these contemporaneous measures, I also use $CAR$ measures in 2-day and 5-day windows, as applied in Huang et al. (2014) and Asquith et al. (2005), respectively. Additionally, I examine the abnormal stock return on the days that reports are released. In Table A5, I find a similar magnitude in $R^{2}_{\mathrm{OOS}}$ for explaining $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% for the out-of-sample period from 2015 to 2023.
In Table A6, I present the information content of analyst report text across the Fama-French 12 industries. The $R^{2}_{\mathrm{OOS}}$ exceeds 10% in five industries: Shops, Other, Manufacturing, Chemicals, and Durables. Notably, the $R^{2}_{\mathrm{OOS}}$ and DM t-statistics are significantly lower in the Non-Durables, Telecommunications, Energy, and Utilities sectors. The industries with higher $R^{2}_{\mathrm{OOS}}$ typically involve complex products and services that benefit from detailed analyst insights, whereas those with lower $R^{2}_{\mathrm{OOS}}$ may operate in more stable or regulated environments where analyst reports add less incremental value. In Figure A3, I plot analyst information value by Fama-French 12 industries, revealing a pattern consistent with the information content results.
To assess the economic magnitude, I perform OLS regressions specified as follows:
$CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (10)
This model incorporates the out-of-sample estimation of $CAR_{[-1,+1]}$ using report text embeddings or forecast revisions as input. For instance, observations in 2015 are fitted using a Ridge regression model trained on samples from 2000 to 2014. The fitted value of $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension. For comparison, I estimate $\widehat{CAR}_{rev}$ using the three forecast revision measures, following the same procedure.
Table 3 summarizes the summary statistics of the sample used and the regression results. The out-of-sample period starts in 2015. I constrain the samples to include only those with valid previous and current stock recommendations, earnings forecasts, and target prices. The final sample for estimating Equation (10) contains 28,837 observations.
Column (1) of Panel B demonstrates that quantitative measures of recommendation revisions, earnings forecast revisions, and target price revisions all contribute to explaining variations in contemporaneous abnormal stock returns, corroborating the findings of Brav and Lehavy (2003) and Huang et al. (2014). Column (2) reveals that $\widehat{CAR}_{txt}$ alone yields an $R^{2}$ of 10.5%, closely aligning with the $R^{2}_{\mathrm{OOS}}$ value in Table 2. The inclusion of both analyst revision measures and $\widehat{CAR}_{txt}$ as regressors elevates $R^{2}$ from 9.1% in column (1) to 14.9% in column (3), underscoring the incremental value of text information.
The positive and significant coefficient on $\widehat{CAR}_{txt}$ indicates that the information in the report text is not subsumed by forecast revisions. Economically, a one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness increases the 3-day abnormal return by 10 basis points, 40 basis points, 90 basis points, and 120 basis points, respectively.
In columns (4) and (5) of Panel B, I report the coefficient estimates for $\widehat{CAR}_{rev}$, which combines the information from all revision metrics. The $\widehat{CAR}_{rev}$ model alone achieves an $R^{2}$ of 8.9%, while combined input improves it to 14.7%. Economically, a one standard deviation increase in $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{txt}$ corresponds to increases of 100 and 120 basis points in realized $CAR$, respectively. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
I further control for $SUE$ and $Prior\_CAR$ in columns (6)-(7) to capture recent news and earnings surprises. The coefficient estimates on $\widehat{CAR}_{rev}$ and $\widehat{CAR}_{rev}$ remain unchanged. The stability of the coefficients suggests that market-valued analyst information is distinct from recent earnings news and is not simply a reaction to past returns.
3.2.2 Information Content around Earnings Announcements
Some studies have emphasized the relationship between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), in terms of how these two information sources reinforce each other in sequence. I examine the information content of sell-side analyst reports around earnings announcements, the most significant corporate events.
Following Chen et al. (2010), I group analyst reports into 13 weekly bins surrounding earnings announcement dates and aggregate their information content by week. Panel A in Table 4 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for each week. Consistent with Huang et al. (2018), a large portion of analyst reports are released immediately after earnings announcements. Notably, the information content is significant only in the first week, with an $R^{2}_{\mathrm{OOS}}$ of approximately 10% and a t-stat of 10.17. This performance drops sharply in the second week, turns negative in the third and fourth weeks.
The distinct performance observed across weekly bins can be attributed to Ridge models effectively capturing dominant patterns in analyst reports surrounding earnings announcements. To account for potential temporal variation, I train Ridge models using separate samples of analyst reports released surrounding and beyond earnings announcements, enabling the models to extract distinct information content based on the timing of the reports. Panel B in Table 4 compares the model performance during earnings announcement periods and non-earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. The $R^{2}_{\mathrm{OOS}}$ of models trained during earnings announcement periods is consistently higher than those trained beyond these periods. The $R^{2}_{\mathrm{OOS}}$ and t-statistics decline as the windows lengthen, declining from a peak of 11.84% to 7.29% over the course of the earnings announcement week. This trend suggests that the market values information from analysts who react swiftly to earnings announcements.
One potential concern is that analysts may simply restate information from earnings calls, and that market reactions are driven by the earnings announcements themselves rather than analyst reports. If analysts merely restate this information, the measured information content of their reports might actually reflect the text from earnings conference calls. To address this issue, I use earnings conference call transcripts to differentiate the information content between the two text sources. I analyze reports released within one day after earnings announcements, when analysts are most likely to simply restate earnings information.
In column (1) of Panel C, I present the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using embeddings from earnings conference call transcript embeddings as input. [^10] Given the extensive length of earnings conference call transcripts (median length of 13,029 tokens), I employed the LLaMA-3-8B model for generating text embeddings in Panel C. This model accommodates up to 8,192 tokens, compared to the 4,096 token limit of the LLaMA-2-13B model used elsewhere in our study. I truncate the text at the length limits to maintain maximum information. The $R^{2}_{\mathrm{OOS}}$ is 4.20% for the S&P 100 firms, with a DM t-stat of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods. The input comprised the mean of all report text embeddings on the day following the earnings announcement. The Ridge regression yielded an $R^{2}_{\mathrm{OOS}}$ of 9.72%, as shown in column (2), notably surpassing the performance of conference call transcripts alone.
Incorporating both analyst report embeddings and earnings conference call embeddings in Ridge regressions resulted in a significant improvement, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% in column (3). The comparison t-statistic between the transcript-only model and the combined model was 5.12, demonstrating substantial incremental value provided by analyst reports. These findings collectively refute the idea that analysts merely restate information from earnings announcements. Instead, they emphasize the unique and timely value analysts bring to the market.
3.2.3 What Content in Analyst Reports is Valued by Market Investors?
After showing that information in report text is significant, a key question arises: what specific content in analyst reports is the most valuable to the market? To systematically evaluate topic importance, I employ the Shapley value approach to attribute the text-based $R^{2}_{\mathrm{OOS}}$ to the 17 topics described in Section 2.1.1. I use sentence-segmented embeddings as report representation. The sentence-segmented approach mitigates the concern of contextual information from one topic influencing the embeddings of tokens in another. The $R^{2}_{\mathrm{OOS}}$ of text input decreases by 2.25% when I alternate from full-context embeddings to sentence-segmented embeddings. This reduction roughly quantifies the magnitude of contextual information content across sentences.
Intuitively, Shapley values are derived by calculating the reduction in $R^{2}_{\mathrm{OOS}}$ when excluding a specific topic’s embedding across all possible topic combinations. These reductions are then averaged over the possible different permutations into a single importance measure. By design, the sum of Shapley values across all 17 topics precisely equals the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings, ensuring an efficient attribution of explanatory power.
The Shapley value decomposition, as illustrated in Figure 2, reveals a striking hierarchy of topic importance in the written reports. The most valuable topics are Income Statement Analyses and Financial Ratio Analyses, each contributing to 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$. Following these, Investment Thesis and Valuation constitute the second tier of importance, with Shapley values of 1.76% and 1.51%, respectively. The remaining topics exhibit minimal or even negative Shapley values, suggesting negligible explanatory power for contemporaneous market reactions to analyst reports.
The robustness of these findings is further reinforced by additional analyses. As demonstrated in Figure A4, scaling Shapley values by sentence count and token length does not alter the preeminence of Income Statement Analyses. Figure A5 reveals consistent topic rankings across industries.
Figure A6 in the Online Appendix shows how the text-based $R^{2}_{\mathrm{OOS}}$ changes over time. A significant drop in $R^{2}_{\mathrm{OOS}}$ is observed between 2020 and 2021, likely due to the disruptive effects of the COVID-19 pandemic on financial markets and the broader economy. Despite the volatility, the scaled Shapley value for the Income Statement category, represented by the red line, remains relatively stable. Analyst discussions of income statements consistently provide valuable information, even during market disruptions.
The relative importance of topics could stem from analysts’ expertise in certain areas (supply effect) or from investors’ varying responses to different topics (demand effect). Figure A7 compares Shapley values across analysts of different experience levels and brokerage firm sizes. The consistency in topic importance across all four subsamples (experienced vs. inexperienced analysts, large vs. small brokerages) suggests that topic rankings are not driven by differences in analysts’ topical expertise.
Building upon the initial findings, I proceed to study the interpretation role and discovery role of analysts in analyzing income statements. Drawing inspiration from Huang et al. (2018) and Blankespoor et al. (2020), I distinguish between the analyst’s role in collecting and generating otherwise unavailable information (discovery role) and their ability to integrate and process available information into more interpretable signals (interpretation role).
To facilitate the analysis, I categorize Income Statement sentences based on two criteria: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
Prompt 3: Please read the following sentence from a sell-side analyst report for the company {firm} ({ticker}) and classify it based on two criteria:
Information Type:
Information Acquisition (0): Sentences that directly report quantitative financial data, such as earnings, revenue, expenses, or other metrics. Example: ‘Reported EPS of \$1.40 beat consensus of \$0.94 and our estimate of \$1.17.’
Information Interpretation (1): Sentences that provide analysis, context, or interpretation of the financial data, such as discussing trends, comparing performance to forecasts, assessing market impacts, or considering strategic implications. Example: ‘Biogen reported 1Q12 MS franchise sales below our estimate and the Street, as Avonex sales were negatively impacted by unfavorable distribution channel dynamics.’
Time Reference:
Realized Income (0): Sentences referring to concrete, historical results from completed periods. Example: ‘The company’s revenue for the fiscal year 2022 was \$500 million.’
Expected Income (1): Sentences containing subjective predictions, or expectations for future periods. Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
Output your classification as two comma-separated numbers: the first for Information Type (0 for Acquisition, 1 for Integration) and the second for Time Reference (0 for Actual, 1 for Forecast).
Output format: InfoType, TimeRef
Sentence to classify: {report_sent}.
Figure 3 provides a detailed breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, with Income Statement subcategories. Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw data. The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on analysis of actual results than on forward-looking projections. These patterns demonstrate that markets particularly value analysts’ interpretation of realized income, more than their reporting of public information or future expectations.
This insight is reinforced by the findings in section 3.2.2, which show that analyst reports issued promptly after earnings announcements are the most informative. Taken together, these results suggest that the market places the highest value on analysts who can quickly digest financial information from earnings announcements and provide insightful interpretations.
3.2.4 Comparison of Revision Reports and Reiteration Reports
Researchers have documented a robust relationship between market prices and analysts’ forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021). Analysts’ forecast revisions provide a significant and timely source of information to the financial market and may constitute a more valuable signal than reports that merely reiterate previous forecasts.
Therefore, I compare the information content of analyst report text associated with forecast revisions and forecast reiterations. Following the criteria of Huang et al. (2014), I consider a report as a revision report if its release date is within two days of the I/B/E/S forecast announcement date. For each forecast target in recommendations, target prices, and earnings forecasts, I categorize the report independently. Consequently, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 5 summarizes the out-of-sample performance of models with numerical and textual inputs using both reiteration and revision samples. Recommendation revisions are rare events compared to target price revisions and earnings forecast revisions. Earnings forecast revisions are the most frequent, comprising 56.57% of the reports. In contrast, recommendation revision reports and target price revision reports constitute only 2.29% and 31.10%, respectively. When stock recommendations are revised, the explanatory power of various information sources becomes particularly pronounced. The quantitative forecast revision measures alone achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99%. The textual content of analyst reports accompanying recommendation revisions explains a larger portion of CAR variation, with an $R^{2}_{\mathrm{OOS}}$ of 16.80%. The combination of both quantitative and qualitative information further enhances the model’s explanatory power, resulting in a $R^{2}_{\mathrm{OOS}}$ of 22.63%. This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions are considered incrementally informative due to the discreteness and potential biases in analysts’ recommendations (Bradshaw, 2002; Brav and Lehavy, 2003). While reports with target price reiterations yield low information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58% for three-day CAR), revision reports demonstrate substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
As Kothari et al. (2016) note, analysts’ earnings forecast revisions convey both public and private information about cash flow news in financial markets and are the most extensively studied type of analyst output. For models with text from earnings forecast reiteration reports as input, the $R^{2}_{\mathrm{OOS}}$ is insignificant. When focusing on analyst earnings forecast revisions, the $R^{2}_{\mathrm{OOS}}$ for both numerical and textual inputs becomes significant. The combination yields an $R^{2}_{\mathrm{OOS}}$ of 16.37%, with a t-statistic of 11.42, underscoring the greater value of revisions over reiterations.
Collectively, Table 5 shows that report content is more informative when it accompanies forecast revisions rather than reiterations. Additionally, analyst reports associated with recommendation revisions demonstrate the highest value.
3.2.5 Comparison with Analyst Report Sentiment
Huang et al. (2014) emphasize the information content of analyst report text by extracting tone using a Naive Bayes machine learning approach. I compare report-level tone measures derived from both Naive Bayes and BERT classification. Following Huang et al. (2014), I manually label 10,000 sentences as training data for both models. Each sentence is categorized as positive, neutral, or negative using these pre-trained models, and I average the tone at the report level to create $Tone_{N}B$ and $Tone_{B}ERT$ measures. I also construct tone measures separately for Income Statement topics and non-Income Statement topics.
Table 6 presents a comparison of performance between text tone measures and text embeddings. Panel A summarizes the OLS regression of three-day $CAR$ on the tone measures. The coefficient estimate for $Tone\_NB$ in column (1) is 0.019, comparable to the 0.0208 reported by Huang et al. (2014). A one standard deviation increase in the favorableness of the text tone results in an additional $CAR$ of 49 basis points, comparable with the 41 basis points increase in two-day abnormal return reported in Huang et al. (2014). Column (2) summarizes the results of regressing $CAR$ on tone measures extracted within Income Statement topics and non-Income Statement topics. Both coefficients of tone are significantly positive. Regression results with firm-fixed and year-fixed effects are consistent, as reported in column (3).
Columns (4)-(6) summarizes regression results with BERT-based tone measures as explanatory variables. The coefficient estimates’ magnitude increases dramatically, along with significance levels. Economically, a one standard deviation increase in BERT-based text tone favorableness results in an additional three-day $CAR$ of 86 basis points. When both tone measures are included in an OLS regression in column (7), coefficient estimates of Naive Bayes-based tone measure become insignificant, suggesting that BERT model captures sentiment more accurately and effectively.
Panel B of Table 6 compares the $R^{2}_{\mathrm{OOS}}$ of tone measures and embeddings. The $R^{2}_{\mathrm{OOS}}$ of tone inputs ranges from 0.05% to 3.68% in the first four rows. The $R^{2}_{\mathrm{OOS}}$ improves to 9.49% and 12.27% after incorporating forecast revisions and text embeddings. The bottom two rows of Panel B report comparisons of $R^{2}_{\mathrm{OOS}}$ and DM tests between models with “Tone + Rev + text” versus “Rev + text.” The difference between the $R^{2}_{\mathrm{OOS}}$ of models with and without tone measures is not statistically distinguishable from zero.
The empirical results in Table 6 highlight two key advantages of LLMs over conventional NLP methods. First, tone measures extracted using LLMs better explain contemporaneous stock returns than those from the Naive Bayes algorithm. Second, neither BERT nor Naive Bayes tone measures add incremental value beyond text embeddings, suggesting that LLMs capture richer contextual information than simple tone metrics.
3.2.6 Information Content with Alternative Machine Learning Models
In this Section, I examine the information content using three alternative machine learning algorithms, applying the same processes as in Ridge regressions. Each model uses full-context report embeddings as inputs, with the target variable being the three-day $CAR$.
Table 7 presents results from PLS regressions. PLS effectively extracts common components that correlate with the target variable. If text embeddings contain strong common components that predict stock returns, PLS regression should capture this information. The PLS regressions achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to those in Ridge models. DM tests confirm that the $R^{2}_{\mathrm{OOS}}$ is significant at the 1% level across all years.
Next, I replicate the analyses with the XGBoost algorithm. Gradient boosting combines weak estimations to form a strong estimation by constructing sequential trees. Each tree is responsible for training a small sample to fit the residuals of the previous trees and updating the model by learning new information. Therefore, if the high-dimensional feature spaces of text embeddings contain segmented information with discrete importance, XGBoost can potentially identify the most informative aspects of the analyst report embeddings for explaining contemporaneous stock returns. Compared to Ridge and PLS models, XGBoost is capable of capturing complex nonlinear relationships between analyst report embeddings and the target variable. As summarized in Table 7, the XGBoost algorithm underperforms Ridge regressions with an overall $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). The relatively poor performance could be attributed to the inherent complexity of language, where information is distributed across the embedding space rather than concentrated in specific dimensions.
I then proceed with Neural Network models, which form the core of LLMs and the transformer architecture. In a Neural Network, nodes are interconnected, mimicking the neuron structure of the human brain. Since embeddings are encoded as nodes within the Neural Network layers, I expect this architecture to capture long-range relationships between words and the complex contextual information embedded in the text. I construct the Neural Network estimations as an ensemble of five models for each NN configuration. As shown in Table 7, the $R^{2}_{\mathrm{OOS}}$ of Neural Network models ranges from 10.42% to 12.12%, achieving the highest performance among all tested ML algorithms. Performance tends to improve with network depth, aligning with the virtue of complexity principle proposed by Kelly et al. (2022).
The experiments with machine learning models yield noteworthy findings. First, the information content of report text, as indicated by $R^{2}_{\mathrm{OOS}}$, achieves around 10% in certain models, which is significant considering the low signal-to-noise ratio in stock returns. Second, the superior performance of deep learning architectures emphasizes the importance of contextual information and word interdependencies in textual analyses.
3.3 Information Value of Analyst Reports
How to translate statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms? In this section, I quantify the information value of analyst reports using the measure introduced in Section 2.3. This measure reflects the profit derived from a reduction in price uncertainty due to an informative signal, while accounting for price impact. In other words, both greater information content in analyst reports and lower price impact will lead to higher estimated information value.
I first summarize the value of information from analyst reports in Table 8, covering the period from 2015 to 2023. Both the mean and standard deviation are derived using the delta method. For S&P 100 constituent stocks, the average three-day window information value from sell-side analysts is \$0.47 million, with a standard error of \$0.05 million. The mean information value of text alone is \$0.38 million, and the mean value for revisions is \$0.34 million. Considering that a stock typically has reports released on an average of 15 days each year, the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm is \$6.89 million. The narrow standard errors and confidence intervals across different information values indicate consistent estimates. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, indicating that \$1 million dollar order pushes the stock price up by 0.034%. These statistics illustrate that information from analysts is economically meaningful, with both textual information and revisions contributing significantly.
Figure 4 describes the evolution of information value from analyst reports over time. The solid line represents the strategic value of combined numerical and textual information, while the dashed line shows the value of textual information alone. Both lines exhibit significant fluctuations over the period. Despite these fluctuations, there is a general upward trend in information value. For most of the period, the combined value is higher than the textual value alone, suggesting that numerical information often complements the textual content in analyst reports.
The shaded area represents the 95% confidence interval. After 2020, there appears to be greater variability in the information value, as indicated by the wider confidence intervals and more pronounced peaks and troughs. This period corresponds with the global COVID-19 pandemic, which likely introduced additional uncertainty and volatility in financial markets, impacting the value of information from analyst reports.
These patterns indicate that the information value of analyst reports is dynamic, and influenced by market conditions. The relationship between combined and text-only information value fluctuates over time, highlighting the importance of considering both numerical and textual components when analyzing analyst reports.
In the Online Appendix, Table A7 provides the summary statistics of alternative estimates for information value. Using high-frequency measures of realized return volatility and different algorithms for signing trades as buys or sells, the estimates for information value from analyst reports range from \$0.42 million to \$0.64 million. Figure A8 illustrates the time trends of these three alternative estimates, revealing consistent patterns in the evolution of information value over time.
3.3.1 The Value of Analyst Information in the Cross-section of Stocks
The information value of sell-side analysts likely vary in the cross-section of stocks. Specifically, the costs of information processing and dissemination may differ by firm size (Zeghal, 1984). For example, analyst opinions on smaller stocks may trigger larger market reactions, as these stocks generally have less publicly available information (Stickel, 1995). I plot the scatter of information value and market capitalization by stock in Figure A9. There is a general upward trend, indicating that stocks with higher market equity tend to have higher information values. However, the relationship is not strictly linear, as there are notable exceptions. I formally test the relationship by regressing the value of information on stock size, firm fixed effects, and year fixed effects. I also control for the book-to-market ratio in the regressions.
Table 9 shows the results. Stocks with larger market equity have significantly higher information value from analysts. Economically, results in column (4) indicate a 1% increase in market equity is associated with a 0.864% increase in analyst information value. Columns (5)-(8) reveal that the information value of report text is more sensitive to size compared to the combined numerical and textual information value. I further investigate whether the positive size elasticity of information value is driven by higher explained return volatility or lower price impact of large stocks. The decomposition can be expressed logarithmically as:
$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
Panel B of Table 9 shows that the higher information value for large stocks stems primarily from their greater liquidity. Analyst reports of large stocks do not explain higher contemporaneous return volatility, yet their value increases with stock size due to enhanced liquidity.
3.3.2 The Value of Analyst Information in the Cross-section of Analysts
Analysts who make bold forecasts incorporate their private information more completely and provide more relevant information to investors than those who herd in their forecasts (Clement and Tse, 2005). I examine how information value varies with analysts’ herding behavior.
To investigate the cross-section of analyst characteristics, I regress the stock-analyst-day information value on an indicator of forecast boldness, controlling for firm-fixed effects, analyst-fixed effects, and year-fixed effects. I also control for broker size, as Jacob et al. (1999) suggests that larger brokerage firms provide better research. The analyst-specific information value measure is defined as
$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
Table 10 presents the results. In columns (1)-(4) of Panel A, the coefficient estimates on $Bold$ are significantly positive, suggesting that analysts’ information is more strategically valuable when they provide bold forecasts. On average, information from bold reports is 29% more valuable than that from herding reports. This magnitude persists when examining the information value of report text alone. The insignificant coefficients for $Brokersize$ suggest that the value of analyst information is not related to brokerage firm size.
A subsequent research question is whether the relationship between analyst boldness and information value is driven by higher report informativeness or lower price impact. In Panel B, I regress the two decompositions of information value on the boldness indicator. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts provide bold forecasts when they possess relevant private information. In contrast, the coefficient estimates in columns (5)-(8), which examine the relationship between boldness and price impact, are statistically insignificant.
3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. Accordingly, I evaluate the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 illustrates the information value of analyst reports between earnings announcements on a weekly basis, from the first week after the earnings announcement up to the $13^{th}$ week. The information value of analyst reports is highest immediately following earnings announcements, particularly in the first week, peaking at a sample mean of \$0.84 million. This value sharply declines by the second week and stabilizes thereafter, with minor fluctuations. These patterns suggest that the market values analyst reports most highly immediately after earnings announcements, with diminishing importance in subsequent weeks.
To formally validate the observed pattern, I conduct a regression analysis of log information value on an indicator variable $Week$ which takes the value of one if the report is released within one week following earnings announcements and its interaction with earnings announcement day trading volume ($TradingVolume$). This model is motivated by prior literature suggesting that analysts’ interpretative role becomes more prominent with the complexity of financial accounting information (Chen et al., 2010). The trading volume serves as a proxy for investor disagreement, following Banerjee (2011) and Beckmann et al. (2024). The interaction term ($Week\times TradingVolume$) captures how post-earnings timing and investor disagreement jointly affect analyst information value. Stock and year-fixed effects are included to control for time-invariant stock characteristics and market-wide trends.
Table 11 presents regression results showing the influence of earnings announcements and trading volume on information value. Panel A indicates that the information value is significantly higher for reports released within one week of earnings announcements ($Week$), with positive and significant coefficients across all models. The interaction between $Week$ and $TradingVolume$ also shows a significant positive effect, suggesting that the information value increases with higher trading volumes around earnings announcements. The rationale is that investors may rely more on analyst opinions when there is greater disagreement in the interpretation of earnings information.
Panel B reveals significant increases in both log return variance and log price impact following earnings announcements, as evidenced by positive coefficients for the $Week$ variable. Notably, the increase in the information value of analyst reports following earnings announcements is primarily attributed to higher explained return volatility, rather than changes in price impact. These findings indicate that analyst reports provide more valuable insights in the first week after earnings announcements, especially during periods of high investor uncertainty. Investors value analysts’ ability to promptly interpret complex earnings information and make sense of the latest earnings numbers.
4 Conclusion
In this paper, I examine both quantitative and qualitative content of analyst reports using $R^{2}_{\mathrm{OOS}}$. The results show that report text contains more market-valued information content than quantifiable forecast revisions economically, especially right following earnings announcements. A Shapley value decomposition method reveals that analysts’ interpretation of realized income statements is the most valuable content within written reports. The text embedding representation extracts more comprehensive information from analyst reports beyond text sentiment.
I further translate the information content into dollar terms to provide a sense of economic magnitude. Using a measure of explained return volatility relative to price impact, I find that the annualized expected profit from early information acquisition of the analyst reports for an average S&P 100 constituent firm is larger than \$6.89 million. While these results are economically meaningful, readers should exercise caution in generalizing beyond the S&P 100 sample. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The significant potential profit from early access to analyst reports raises concerns about selective distribution. Such practices would violate FINRA Rule 2241, which aims to ensure the objectivity of research and prevent the selective distribution of research reports to preferred clients ahead of others. It is crucial to emphasize that this study presents a hypothetical scenario without assessing or endorsing actual market practices. Rather, by quantifying the potential profit of privileged access to analyst information, this research emphasizes the importance of regulations in ensuring fair disclosure.
A. Additional Figures
(a) (1) Executive Summary
(b) (2) Company Overview
(c) (3) Industry Analysis
(d) (4) Competitive Landscape
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (5) Income Statement Analysis
(b) (6) Balance Sheet Analysis
(c) (7) Cash Flow Analysis
(d) (8) Financial Ratios
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (9) Business Segments
(b) (10) Growth Strategies
(c) (11) Risk Factors
(d) (12) Management and Governance
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (13) ESG Factors
(b) (14) Valuation
(c) (15) Investment Thesis
(d) (16) Appendices and Disclosures
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
Figure A2: Large Language Model Architecture
This figure shows an example of how large language models process a sentence from an analyst report on Nvidia. The model breaks this sentence into tokens, and then generates two types of embeddings: token embeddings and position embeddings. These embeddings then pass through the core of the model - a stack of transformer layers. The output of each transformer layer is a series of vectors representing the tokens in their full context.
Figure A3: Analyst Information Value by Industry
This figure shows the estimated information value of analyst reports for Fama-French 12 industries from 2015Q1 to 2023Q4, quantified in millions of dollars. The solid green line marks the sample mean value of information, with green bars showing the 95% confidence intervals. The horizontal dashed red line marks the sample mean of stocks from all industries. Information value estimates are derived using the delta method and adjusted to 2020 levels using the Consumer Price Index (CPI).
Figure A4: Shapley Values Scaled by Length
This figure presents the SHAP values of topics scaled by the total number of sentences (Panel A) and the total number of tokens (Panel B) for topics with more than 100,000 sentences. The scaling process involves two steps: First, normalize the Shapley value of each topic by dividing it by the number of sentences in that topic. Then, adjust the normalized values to ensure the total sum of SHAP values remains consistent. This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
Figure A5: Shapley Values for Topic Importance across Industries
This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value. For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings. The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
Figure A6: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
This figure shows the $R^{2}_{\mathrm{OOS}}$ of sentence-segmented report embeddings from 2015 to 2023. The color gradient indicates the magnitude of $R^{2}_{\mathrm{OOS}}$. The red line represents the SHAP value of the Income Statements Analyses topic across the years. The shaded area marks the pandemic recession in 2020.
Figure A7: Shapley Values for Topic Importance by Analyst Characteristics
This figure illustrates how 17 different topics contribute to the predictive power of analyst reports across four subsamples, measured using SHAP values. The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), and large (Panel C) versus small brokerage firms (Panel D), with subsamples determined by yearly median. For each subsample, the bars show the SHAP values of individual topics, whose sum equals the out-of-sample R-squared ($R^{2}_{\mathrm{OOS}}$) from the sentence-segmented report embeddings. The red line indicates each topic’s relative importance, calculated as the topic’s SHAP value divided by the total $R^{2}_{\mathrm{OOS}}$.
Figure A8: Alternative Analyst Information Value Estimations Over Time
This figure presents the alternative estimations of analyst information value from 2015Q1 to 2024Q4, reported in millions of dollars and adjusted to 2020 dollars. Quarterly mean approximations are calculated using the delta method. TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Figure A9: Analyst Information Value by Stock
This figure shows the estimated information value of analyst reports for individual stocks from 2015Q1 to 2023Q4, quantified in millions of dollars. Estimates are derived using the delta method. Market equity, averaged daily at the time of each report release, is reported in billions of dollars. All values are adjusted to 2020 levels using the Consumer Price Index (CPI).
B. Additional Tables
Table A1: Topic-Specific Sentence Examples from Analyst Reports
Topics Examples
Executive Summary
Our key takeaways from CEO Jeff Immelt’s presentation at EPG were: Outlook for substantial EPS growth over 2010/12 driven by abatement of credit losses (\$8-9bn in 2010E tapering to \$4bn run-rate) and CRE impair. Looking ahead, guidance was tightened, essentially framing the Street, and commentary suggests a strong outlook for the balance of 2007. Key topics: 1) Update on ABTÕs portfolio of COVID-19 tests; 2) Libre trends, Libre 2 launch and expectations for Libre 3; 3) Elective surgery trends exiting Q1, expectations for 2021 and an update on recent and upcoming new product approvals; and 4) Global trends and impact on EPD and Nutrition.
Company Overview
Oracle Corporation, founded in 1977 and headquartered in Redwood Shores, California, is one of the largest and most prominent companies in the software space – and a technology bellwether. As it has grown, Microsoft has expanded into enterprise software with Windows Server, SQL Server, Dynamics CRM, SharePoint, Azure and Lync; hardware with the Xbox gaming/media platform and the Surface tablet; and online services through MSN and Bing. Altria Group, Inc., is the world’s largest producer and marketer of consumer products, and had revenues of \$80 billion in 2002.
Industry Analysis
According to our global Immunology market model, US Psoriasis (PsO) represented a \$7.7B market in 2016 and is expected to grow at a low-teens CAGR to \$11.8B in 2019E and  \$13B by 2021E driven by more highly effective therapies. Despite record prices, oil demand continues to grow, while supply growth lags and spare production and refining capacity is almost nonexistent. Add to that that some new advertising expectations from PricewaterhouseCoopers of a decline in ad spending of 12% worldwide and 15% in the U.S. for 2009 and continue to decline in 2010.
Competitive Landscape
According to Mercury Research, NVIDIA is now the 3rd largest chipset supplier (consisting of desktop and mobile chipsets, and integrated and non-integrated chipsets), shipping 5.4 million units in calendar Q3 for an 8.2% market share, versus Intel’s shipments of 51 million units (62% share), VIA’s shipments of 9.6 million units (14.4% share), SiS’s shipments of 5.3 million units (8% share), and ATI’s shipments of 4.4 million units (6.6% share). Further, competition in the CDK-4/6 space is rising with Verzenio (abemaciclib) & Kisqali launches placing downward pressure on Ibrance trajectory. We believe that Accenture has recognized that web services could compete directly with client/server as the new systems architecture.
Financial Statement Analysis: Income Statement Analysis
We also note that Alico cannot dividend any of its 2010 earnings to AIG, implying that the income recognized this quarter and throughout 2010 will benefit MetLife results. At the respective midpoints, sales of \$8.5 billion would be down 22% annually and 8% sequentially; and non-GAAP EPS would be down 39%. The growth in operating income from Q1 2009 is largely due to recoveries in partnership income.
Financial Statement Analysis: Balance Sheet Analysis
Long-term borrowings of \$6.59 billion at September 30, 2022 were modest compared to shareholders’ equity of \$37.3 billion. The company finished 2Q16 with \$21.4 billion in cash and short-term investments, up from \$15.6 Growth & Valuation Analysis GROWTH ANALYSIS RISK ANALYSIS billion at the end of 4Q15. Inventory turns improved to 4.5x in 4Q from 4.4x in the same period last year.
Financial Statement Analysis: Cash Flow Analysis
UPS generated \$2.3 billion in operating cash for the quarter. Assuming that the company completes a large portion of its current \$1 billion stock buyback plan in 2008, we estimate that cash per share will be about \$6.20 by the end of the year. We view eBays FCF generation as relatively defensible even in the case of a revenue shortfall.
Financial Statement Analysis: Financial Ratios
The company achieves average scores on our three main measures of financial strength: leverage based on debt-to-cap, profitability and interest coverage. P/S is at 0.6x and EV/EBITDA is 7.7x. The index members currently trade at an average of 16.3-times trailing earnings, which is below the five-year average of 19-times.
Business Segments
The company is organized into three businesses, software, representing the majority of the company’s total revenues, hardware systems and services. Results in 2008 also benefited from the absence of the significant level of mark-to-market losses in the company’s Gas Marketing segment in 2007. The company operates five distinct segments: Americas (71% of FY15 profit); Europe, Middle East, and Africa (4% of FY15 profit); China and Asia Pacific (12% of FY15 profit); Channel Development (13% of FY15 profit).
Growth Strategies
Visa has been especially active on the acquisition front over the last several months. Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle. Specific areas where more investment may be needed: Over the past decade, ROK management often claim that Process automation represents the growth opportunity for the company, but its sales in this market have barely grown in recent years, despite its much smaller sales base compared with established incumbents.
Risk Factors
Risks to our BUY thesis have to do with global competition, changing user behavior, global macro uncertainty, and anything else that can affect FB’s relationship with members, its advertisers or its publishing partners. Risks to achieving our price target include: 1) Apple crushing PayPal; 2) increasing competition in the payments space; 3) heavy investment spending on marketing, point of sale, or technology; and 4) legislative action. In addition to the expenses incurred by patent challenges, product liability and other legal suits could occur and lead to additional liabilities and revenue loss, which could substantially change our financial assumptions.
Management and Governance
Top management changes can be unsettling, and the resulting uncertainty has caused 3M shares to decline. Chairman, President, and CEO Charles Ergen beneficially owns about 53.6% of DISH’s equity securities and has 90.5% voting power. Bill Johnson, the present CEO of Progress Energy, will become president and chief executive officer of the new company.
Environmental, Social, and Governance (ESG) Factors
The assessment of ESG (Environmental, Social & Governance) risks by Baptista Research includes a wide range of considerations that pertain to the long-term sustainability of a company. Sustainalytics assesses the degree to which a company’s enterprise business is affected by ESG issues. Failure to adequately address social risks like labor disputes and community relations could jeopardize the company’s social license to operate in certain regions.
Valuation
Our DCF derives an intrinsic value of ¿\$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC. We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment. Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
Investment Thesis
The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road. As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high. Clearly, our Ford Investment Thesis, which was based in large part on our belief that Ford would be able to offset headwinds (slowing cyclical tailwinds in North America, weakness in South America, weak growth in Europe, slowing growth in China, and regulatory cost headwinds), has been thrown into question.
Appendices and Disclosures
Although the information contained in the subject report has been obtained from sources, we believe to be reliable, its accuracy and completeness cannot be guaranteed. For a complete discussion of the risk factors that could affect the market price of a company’s shares, refer to the most recent Form 10-Q or 10-K that a company has filed with the Securities and Exchange Commission. The Benchmark Company, LLC makes every effort to use reliable, comprehensive information, but we make no representation that it is accurate or complete.
Table A2: Numerical Variables Description
This table shows the definitions of numerical measures.
Numerical Measures Definition and/or sources
Panel A: Firm Level Measures
Size The market value equity of the firm ($CSHOQ*PRCCQ$) at the end of the quarter prior to report release.
BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of quarter prior the report release.
Prior_CAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement. Consensus EPS is the median value of 1-year EPS forecast within 90-days window of all analysts following the firm. The unexpected earnings is scaled by price per share at the fiscal quarter end.
AbsSUE Absolue value of SUE, representing the distance between realized EPS and EPS expectation.
Miss Dummy variable that equals one if the actual EPS is less than the last consensus forecast, and 0 otherwise.
Trading volume Trading volume at earnings announcement day (or the first trading day post earnings announcement), calculated as $VOL/SHROUT$.
Distance to default The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
Fluidity A measure of firms’ product market competition introduced by Hoberg et al. (2014). The data is compiled from the Hoberg and Phillips database.
Panel B: Industry Level Measures
Industry recession An indicator variable that equals one if the FF-48 industry return is negative and in the bottom quintile of FF-48 industry returns and zero otherwise.
Panel C: Macroeconomic Measures
Time trend (ttr) The number of years elapsed from the beginning of the sample.
Panel D: Report Level Measures
$REC\_REV$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
$EF\_REV$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
$TP\_REV$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
Boldness An indicator variable that takes the value of 1 if the EPS forecast revision is above both the analyst’s own prior forecast and the consensus forecast, or else below both, and zero otherwise.
SR Stock recommendation from I/B/E/S rating, with 1 being the most bullish (Strong Buy) and 5 being the most bearish (Sell), based on the ratings provided by the Institutional Brokers’ Estimate System (I/B/E/S).
ERet 12-month return forecast by scaling the 12-month price target by the stock price 1-day before release.
Table A3: Sumamry Statistics for Numerical Measures
This table reports summary statistics of numerical measures.
Mean Std p25 P50 P75 N
Miss 0.35 0.48 0.00 0.00 1.00 102776
Trading volume 18.19 21.98 7.24 11.46 20.29 99908
Prior_CAR 0.00 0.05 -0.02 0.00 0.02 122251
Size 2.39 0.11 2.33 2.39 2.46 117534
BtoM 0.46 0.43 0.18 0.32 0.61 117534
Distance to default 0.00 0.01 0.00 0.00 0.00 86693
Fluidity 7.17 3.66 4.40 6.57 9.44 114942
Time trend 13.02 5.71 9.00 13.00 17.00 122252
Industry recession 0.14 0.35 0.00 0.00 0.00 122252
BrokerSize 71.92 56.71 26.00 53.00 113.00 119233
Firm Experience 7.68 6.72 3.00 6.00 11.00 111192
Number of firms 20.08 17.08 14.00 18.00 23.00 119233
REC_REV 0.00 0.16 0.00 0.00 0.00 120673
EF_REV 0.00 0.00 0.00 0.00 0.00 90625
TP_REV 0.01 0.07 0.00 0.00 0.00 84108
ERet 0.19 0.23 0.07 0.18 0.31 84108
Boldness 0.75 0.43 1.00 1.00 1.00 81182
SR 2.77 0.84 2.00 3.00 3.00 122252
Table A4: Information Content of Analyst Reports with Numbers Removed
This table presents the $R^{2}_{\mathrm{OOS}}$ using analyst reports with numbers removed. The input sets include revision numerical measures, text embeddings from number-free report text, and a combination of both. The estimation targets are the three-day $CAR$ centered around the analyst report release date. The machine learning model employed is Ridge regression, with training samples expanding yearly. The $t$-statistic for the $R^{2}_{\mathrm{OOS}}$ is calculated following Gu et al. (2020). In columns (2), (4), and (6), the benchmark estimation is set to zero. Columns (3)-(1) and (5)-(3) present pairwise tests comparing the performance of models using revision-only input versus text-only input, and text-only input versus combined text and revision input, respectively.
Year Revision only t-stat Text only t-stat Revision plus text t-stat t-stat t-stat
(1) (2) (3) (4) (5) (6) (3)- (1) (5)- (3)
2015 10.30% 3.59 13.19% 5.87 11.10% 3.24 3.97 -1.45
2016 15.47% 11.80 13.73% 5.00 17.91% 9.02 -0.81 3.13
2017 9.16% 5.28 11.66% 5.98 12.33% 6.12 10.58 1.56
2018 9.79% 2.98 11.46% 9.12 13.62% 7.97 1.22 4.09
2019 9.76% 15.69 12.61% 14.73 13.90% 15.17 4.26 2.79
2020 5.20% 3.43 4.42% 4.38 6.98% 5.81 -0.70 8.37
2021 5.31% 4.80 8.77% 5.06 10.66% 29.67 1.54 1.56
2022 10.24% 6.16 16.71% 10.47 17.90% 8.94 9.93 2.38
2023 5.09% 3.86 9.27% 3.60 10.39% 5.09 2.60 1.81
Overall 8.93% 8.51 10.95% 7.87 12.51% 8.83 2.40 3.00
Table A5: Cumulative Abnormal Returns with Alternative Windows
This table reports the $R^{2}_{\mathrm{OOS}}$ of using analyst report text embeddings to estimate $CAR$ with alternative windows. The $R^{2}_{\mathrm{OOS}}$ is calculated annually using a training sample from 2000 to the preceding year. $CAR$ measures are centered on the release day of analyst reports ($T_{0}$). Panel A shows results for the full sample, while Panels B and C focus on periods within and beyond 1-day earnings announcement windows, respectively. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
AR[0] CAR[0,+1] CAR[-2,+2]
year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
Panel A: Full sample period
2015 8.03% 8.53 8.24% 8.36 10.87% 9.40
2016 7.88% 8.27 6.09% 4.68 6.60% 3.11
2017 8.75% 14.35 8.56% 9.29 9.00% 8.78
2018 9.09% 14.95 9.39% 11.26 9.75% 10.54
2019 9.89% 99.70 7.68% 9.45 8.66% 18.93
2020 2.33% 2.33 1.76% 3.15 4.13% 2.14
2021 5.75% 7.32 5.80% 6.62 4.05% 3.57
2022 10.54% 12.32 11.23% 18.20 10.91% 13.65
2023 5.38% 15.44 3.39% 7.00 0.53% 1.48
Overall 7.55% 8.07 6.92% 6.77 7.57% 6.79
Panel B: Earnings announcement period
Overall 9.78% 5.58 7.87% 5.10 13.14% 8.02
Panel C: Non-earnings announcement period
Overall 5.77% 5.78 5.62% 5.64 4.96% 4.91
Table A6: Information Content of Analyst Reports across Industries
This table reports the $R^{2}_{\mathrm{OOS}}$ of Fama-French 12 industries, using the definition from Kenneth French’s website. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
year Shops Other Manuf Chems
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 13.27% 6.03 657 17.69% 5.03 585 10.84% 3.19 808 15.05% 1.74 147
2016 10.43% 2.46 556 3.34% 0.48 593 10.37% 5.38 811 23.91% 3.25 111
2017 16.65% 3.49 506 20.45% 2.67 544 25.03% 13.64 722 11.04% 0.59 82
2018 10.86% 4.32 406 21.65% 3.40 405 3.32% 1.55 553 25.80% 1.87 77
2019 21.04% 9.24 394 7.49% 2.37 610 16.25% 12.07 568 5.00% 3.22 106
2020 4.39% 2.55 499 -0.30% -0.53 499 4.06% 1.41 542 2.65% 1.60 115
2021 10.27% 2.13 312 9.70% 3.66 362 8.66% 4.32 367 1.46% 0.30 102
2022 22.75% 3.83 320 15.98% 2.66 437 14.26% 11.46 420 7.00% 2.10 110
2023 17.78% 1.76 353 3.49% 2.00 339 7.15% 2.71 360 19.31% 6.54 76
Overall 13.88% 6.03 4003 11.27% 3.69 4374 11.39% 5.57 5151 11.35% 3.55 926
year Durbl BusEq Hlth NoDur
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 16.83% 5.77 216 15.66% 6.41 1378 5.75% 2.72 1242 0.68% 0.31 241
2016 4.95% 0.79 209 10.24% 6.09 1303 11.07% 2.90 1185 17.14% 5.18 242
2017 -2.17% -0.30 282 10.36% 3.12 1154 3.37% 2.39 1060 -2.14% -0.68 278
2018 18.29% 5.57 208 0.32% 0.31 1002 12.47% 6.58 918 1.29% 0.13 324
2019 15.31% 2.26 222 10.17% 6.13 1036 10.02% 7.68 884 18.51% 2.30 277
2020 5.73% 0.79 168 8.11% 4.14 1100 2.86% 1.84 950 -0.37% -0.28 265
2021 10.01% 5.31 150 11.43% 1.49 731 7.66% 4.58 592 -6.38% -0.60 195
2022 7.99% 2.58 162 12.08% 4.77 822 12.05% 2.71 850 5.24% 1.11 187
2023 12.05% 2.37 140 10.45% 11.80 962 3.52% 1.32 844 -5.69% -0.71 158
Overall 11.02% 3.24 1757 9.68% 6.77 9488 7.53% 5.71 8525 7.50% 1.38 2167
year Money Telcm Enrgy Utils
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 18.13% 4.41 1153 1.68% 0.00 403 -4.21% -1.45 303 18.47% 3.69 401
2016 7.64% 3.53 1108 -10.18% -1.22 329 15.80% 2.83 252 -1.84% -0.23 310
2017 1.77% 0.46 1071 8.60% 5.16 307 -18.18% -8.75 307 -39.46% -6.99 315
2018 13.55% 8.48 879 10.65% 2.46 219 8.38% 1.29 275 -23.76% -5.53 215
2019 2.10% 2.25 977 1.53% 0.52 351 -10.28% -9.07 340 -2.36% -0.27 193
2020 1.42% 0.76 952 -8.84% -2.83 360 5.72% 2.62 434 -7.29% -2.36 165
2021 2.62% 0.47 614 6.15% 3.27 188 -15.04% -6.83 251 -4.01% -1.16 103
2022 4.34% 2.13 667 12.84% 3.50 143 16.14% 4.44 286 8.31% 2.71 132
2023 -0.63% -0.18 569 8.99% 6.89 182 -27.33% -5.03 290 -6.43% -1.12 120
Overall 5.18% 3.31 7990 3.32% 1.39 2482 1.90% 0.31 2738 -1.97% -0.88 1954
Table A7: Summary Statistics for Alternative Analyst Information Value Estimations
The analyst information value is measured as the explained return variance divided by the price impact, with the results reported in millions of dollars. All dollar values are adjusted for inflation to reflect 2020 values using the Consumer Price Index (CPI). The sample comprises analyst reports targeting common constituent firms of the S&P 100 index from 2015Q1 to 2023Q4. Both mean and standard deviation are computed using the delta method.TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. Intraday denotes that $\sigma_{v}^{2}$ does not include overnight return volatility. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Mean SE 95%CI 99%CI N
Information value (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17669
Information value of text (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17672
Information value of revisions (TAQ volitility), \$M 0.42 0.04 [0.34, 0.50] [0.31, 0.53] 17672
Information value (Intraday), \$M 0.42 0.04 [0.34, 0.51] [0.32, 0.53] 17672
Information value of text (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17669
Information value of revisions (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17672
Information value (EMO), \$M 0.34 0.04 [0.27, 0.41] [0.25, 0.43] 17672
Information value of text (EMO), \$M 0.35 0.04 [0.27, 0.42] [0.25, 0.44] 17672
Information value of revisions (EMO), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17669
Information value (CLNV), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17672
Information value of text (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.41] 17672
Information value of revisions (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.42] 17672
C. Machine Learning Models
Ridge Regression
Ridge regression addresses multicollinearity by adding a regularization term to the least squares objective function. The ridge regression estimator is given by:
[…]
where $\alpha$ is the regularization parameter that controls the trade-off between fitting the data and shrinking the coefficients.
To find the optimal value of $\alpha$, cross-validation is used over a grid of values ranging from $10^{-10}$ to $10^{10}$. The cross-validation process ensures that the chosen model generalizes well to unseen data, preventing overfitting while capturing the predictive power of the text embeddings.
Partial Least Square Regression
To mitigate the risk of overfitting inherent in high-dimensional text embeddings, I employ Partial Least Squares (PLS) for dimensionality reduction.
The optimization problem can be expressed as follows:
[…]
where $\Omega$ is a $K\times P$ transformation matrix that reduces the $K$ predictors in $X_{i}$ to $P$ lower-dimensional components.
The extraction of the $j$-th PLS component is guided by the following objective function:
[…]
In essence, this approach sequentially extracts components that maximize the covariance with the outcome variable, while ensuring orthogonality to previously extracted components.
Extreme Gradient Boosting
Tree-based approaches are commonly applied in stock return forecasting literature (see, e.g., Gu et al., 2020; Cao et al., 2024; Bonini et al., 2023). XGBoost is an advanced implementation of tree-based machine learning models that builds an ensemble of decision trees. In XGBoost, each tree is added sequentially to correct the errors of previous trees. The main idea is to combine the outputs of multiple weak learners (decision trees) to create a strong learner. XGBoost incorporates regularization techniques, such as L1 (Lasso) and L2 (Ridge), to prevent overfitting and enhance model generalization.
In comparison to random forests, which build a multitude of independent trees and aggregate their predictions, XGBoost constructs trees sequentially, with each tree designed to correct the errors of the preceding ones. While random forests rely on bagging, a method that combines the predictions of various trees to reduce variance, XGBoost uses boosting, an approach that aims to reduce both bias and variance by focusing on difficult-to-predict instances in subsequent iterations. This boosting approach allows XGBoost to effectively capture and leverage the nuanced information embedded in textual data, leading to a more accurate estimation of stock returns.
Neural Networks
To extend beyond the linear modeling approach, I explore the use of Neural Networks to estimate $CAR$ using text embeddings derived from analyst reports. A Neural Network can capture complex non-linear relationships between the text embeddings and the $CAR$, potentially improving the model’s fitting capabilities. Consider a three-layer Neural Network as the prediction model. The prediction problem can be formulated as follows:
$f(X_{i};\theta)=W_{3}\sigma\left(W_{2}\sigma\left(W_{1}X_{i}+b_{1}\right)+b_{2}\right)+b_{3},$ (16)
where $\sigma(\cdot)$ is the ReLU activation function, $W_{i}$ and $b_{i}$ represent the weights and biases for layer $i$, respectively. The Neural Network architecture consists of an input layer representing the text embeddings, followed by multiple hidden layers. The specific architecture employed includes 32 neurons in the first hidden layer, followed by 16, 8, 4, and 2 neurons in subsequent layers. This structure is flexible and can be adjusted by adding or removing layers as necessary to optimize performance.
The training process involves optimizing the weights and biases to minimize the loss function, which, in this case, is the mean-square loss. To regularize the model and prevent overfitting, early stopping is implemented, halting training once the validation loss ceases to decrease. Additionally, following Gu et al. (2020), the models are retrained five times, and the final estimation is obtained by averaging the outputs of these five models, forming an ensemble estimation.
D. Theory
I provide an intuition for the measure of strategic value by discussing a simple extension of Kyle (1985) model. In the Kyle model, there is one risky asset with a payoff $\tilde{v}\sim N\left(p_{0},\Sigma_{0}\right)$. Three types of traders exist: a strategic trader with insider information, a market maker who sets prices in a perfectly competitive market, and an uninformed trader who trades $\tilde{u}\sim N\left(0,\sigma_{u}^{2}\right)$, where $\tilde{u}$ is independent of $\tilde{v}$. Illiquidity is measured by Kyle’s lambda ($\lambda$). Kyle’s lambda depends on private information and liquidity trading. I extend the model by considering a case where $\phi$ percentage of variance in $\tilde{v}$ is explained by the informed trader’s information.
A Single Auction Model
There are two periods, $t_{0}$ and $t_{1}$. The asset is traded with asymmetric information at period $t_{0}$, and the value $\tilde{v}$ is realized at period $t_{1}$. I assume without loss of generality that
$\tilde{v}=P_{0}+\tilde{s}+\tilde{\epsilon},$
where $\tilde{s}$ is a mean-zero signal observed by the informed trader at $t_{0}$, and $\tilde{\epsilon}$ is the combination of residual information and noise. I assume that the signal $\tilde{s}$ and residual $\tilde{\epsilon}$ are uncorrelated, that is, $\sigma_{v}^{2}=\sigma_{s}^{2}+\sigma_{\epsilon}^{2}$.
Let
$\phi=\frac{\operatorname{var}(\tilde{s})}{\operatorname{var}\left(\tilde{v}\right)}=\frac{\sigma_{s}^{2}}{\sigma_{v}^{2}}.$
This measure $\phi$ is the ”R-square” of projecting $\tilde{v}$ on $\tilde{s}$, or the percentage of explainable variance in $\tilde{v}$ using signal $\tilde{s}$.
After observing $\tilde{s}$, the informed trader submits a market order $\tilde{x}=X(\tilde{s})$, and the uninformed trader trades a zero-mean random variable $\tilde{z}$ that is normally distributed and independent of $\tilde{v}$. The profit of the informed trader is given by: $\tilde{\pi}=(\tilde{v}-\tilde{p})\tilde{v}$. The insider has a rational guess of $P(\tilde{x}+\tilde{u})$ and understands that his order $\tilde{x}$ will move the price against him.
The market maker observes the order flow $\tilde{y}\stackrel{{\scriptstyle\text{def}}}{{=}}\tilde{x}+\tilde{z}$ and determines the equilibrium price $\tilde{p}=P(\tilde{x}+\tilde{u})$ to break even. The assumptions for the market maker are risk neutrality and perfect competition, which drives the profits for market makers to zero.
Equilibrium
An equilibrium is a set of $X$ and $P$ satisfying
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s],$
$\tilde{p}(X,P)=E[\tilde{v}\mid\tilde{x}+\tilde{u}].$
The first condition is profit maximization, stating that given the market maker’s pricing rule, the insider chooses a strategy $X$ maximizing her conditional expected profit, taking into account the pricing rule. The second condition is market efficiency. Given the insiders’ trading strategy, the market maker sets the price to be the expected value of the security.
Conjecture:
$P(\tilde{y})=\mu+\lambda\tilde{y},$
$X(\tilde{s})=\alpha+\beta\tilde{s}.$
The profit of the insider can be written as:
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s]$ $=E[(\tilde{v}-\mu-\lambda(\tilde{u}+x))x\mid\tilde{s}=s]$
$=(P_{0}+s+-\mu-\lambda x)x.$
Traders take into account that her order flow will move the price against her, which serves to restrain her position size.
Solving for optimal profit, I get:
$x^{*}=\frac{P_{0}+s-\mu}{2\lambda}=\alpha+\beta v.$
Hence, I can express $\alpha$ and $\beta$ :
$\begin{gathered}\beta=\frac{1}{2\lambda},\\ \alpha=\frac{P_{0}-\mu}{2\lambda}=(P_{0}-\mu)\beta.\end{gathered}$
When the market maker puts a higher weight on the order flow in setting the price, the trader puts a lower weight on her information.
I now look at the price-setting rule:
$\mu+\lambda y=E\{\tilde{v}\mid\alpha+\beta\tilde{s}+\tilde{u}=y\}.$
Essentially, the market maker observes a normally distributed signal about
$\mu+\lambda y$ $=E[\tilde{v}\mid y]$
$=\bar{v}+\frac{\operatorname{cov}(\tilde{v},\alpha+\beta\tilde{s}+\tilde{u})}{\operatorname{var}(\alpha+\beta\tilde{s}+\tilde{u})}(\beta\tilde{s}-\beta\bar{s}+\tilde{u})$
$=p_{0}+\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}}(y-\alpha).$
Hence, I can express $\lambda$ and $\mu$ :
$\lambda=\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}},$
$\mu=p_{0}-\alpha\lambda.$
There is a unique linear equilibrium given by
$\mu=P_{0},$
$\lambda=\frac{\sigma_{s}}{2\sigma_{u}},$
$\alpha=0,$
$\beta=\frac{\sigma_{u}}{\sigma_{s}}.$
Discussion
Kyle’s Lambda
The parameter $\lambda$ is universally known as Kyle’s lambda. Formally, it is the impact on the equilibrium price of a unit order. Its reciprocal ($1/\lambda$) measures the liquidity (or depth) of the market. If $1/\lambda=\frac{2\sigma_{u}}{\sigma_{s}}$ is larger, then the market is more liquid, either because there is less private information in $\sigma_{s}$ or there is more liquid trade in the sense of $\sigma_{u}$.
Value of Private Information
Notice that the equilibrium strategy of the informed trader is $\tilde{x}=\beta s$. The unconditional expected gain of the informed trader is
$\mathrm{E}[\tilde{x}[\tilde{v}-p(\tilde{x}+\tilde{z})]]$ $=\beta\mathrm{E}[s(\tilde{v}-\mu-\lambda\beta s-\lambda\tilde{u})]$
$=\beta(1-\lambda\beta)\sigma_{s}^{2}$
$=\frac{\sigma_{s}\sigma_{u}}{2}$
$=\frac{\phi\sigma_{v}^{2}}{4\lambda}.$
The expected gain for the informed trader is maximized when she has more private information or when there is more liquidity trading. On the other hand, liquidity traders incur losses equivalent to the informed trader’s gains, but they accept these losses due to other motives for trading. The equilibrium price ensures that market makers do not profit or lose in expectation.
E. Delta Method Approximation
I use the delta method to approximate the expected value and variance of a ratio of two random variables. This method relies on a first-order Taylor expansion. Specifically, consider two random variables $X$ and $Y$ with means $\mu_{X}$ and $\mu_{Y}$, and variances $\sigma_{X}^{2}$ and $\sigma_{Y}^{2}$, respectively. We are interested in the ratio $Z=\frac{X}{Y}$ and want to approximate the mean $E[Z]$ and the variance $\text{Var}(Z)$.
Delta Method Approximation
The key idea is to approximate the function $g(X,Y)=\frac{X}{Y}$ using a Taylor series expansion around the means $\mu_{X}$ and $\mu_{Y}$. For the function $g(X,Y)$, I use a first-order Taylor expansion around $(\mu_{X},\mu_{Y})$:
$g(X,Y)\approx g(\mu_{X},\mu_{Y})+\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}(X-\mu_{X})+\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}(Y-\mu_{Y}).$
The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are:
$\frac{\partial g}{\partial X}=\frac{1}{Y},\quad\frac{\partial g}{\partial Y}=-\frac{X}{Y^{2}}.$
Evaluating these at $(\mu_{X},\mu_{Y})$, I get:
$\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}=\frac{1}{\mu_{Y}},\quad\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}=-\frac{\mu_{X}}{\mu_{Y}^{2}}.$
Substituting the partial derivatives into the Taylor expansion, I obtain:
$\frac{X}{Y}\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}(X-\mu_{X})-\frac{\mu_{X}}{\mu_{Y}^{2}}(Y-\mu_{Y}).$
Taking the expectation on both sides:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}E[X-\mu_{X}]-\frac{\mu_{X}}{\mu_{Y}^{2}}E[Y-\mu_{Y}].$
Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation simplifies to:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}.$
Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by:
$\operatorname{Var}(Z)\approx\left(\frac{\partial g}{\partial X}\right)^{2}\operatorname{Var}(X)+\left(\frac{\partial g}{\partial Y}\right)^{2}\operatorname{Var}(Y)+2\left(\frac{\partial g}{\partial X}\right)\left(\frac{\partial g}{\partial Y}\right)\operatorname{Cov}(X,Y).$
Substituting the partial derivatives:
$\operatorname{Var}\left(\frac{X}{Y}\right)\approx\frac{\operatorname{Var}(X)}{\mu_{Y}^{2}}+\frac{\mu_{X}^{2}\operatorname{Var}(Y)}{\mu_{Y}^{4}}-2\frac{\mu_{X}\operatorname{Cov}(X,Y)}{\mu_{Y}^{3}}.$
Application to Value of Information
In the context of the provided study, I estimate the mean value of information for a subsample $s$ using the delta method.
Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$. Define the mean and variance over subsample $s$ as:
$\mu_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right),$
and the mean price impact per dollar over subsample $s$ as:
$\mu_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}.$
Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by:
$E\widehat{\Omega}_{s}\approx\frac{\mu_{\nu s}}{\mu_{\lambda s}}.$
The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as:
$\text{Var}\left(\widehat{\Omega}_{s}\right)\approx\frac{1}{\mu_{\lambda s}^{2}}\left(\Sigma_{\nu s}+\frac{\mu_{\nu s}^{2}}{\mu_{\lambda s}^{2}}\Sigma_{\lambda s}-2\frac{\mu_{\nu s}}{\mu_{\lambda s}}\Sigma_{\nu\lambda s}\right),$
where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as:
$\Sigma_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(\widehat{\sigma}_{it}^{2}-\mu_{\nu s}\right)^{2},$
$\Sigma_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\left(\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}-\mu_{\lambda s}\right)^{2},$
$\Sigma_{\nu\lambda s}=\text{Cov}\left(\widehat{\sigma}_{it}^{2},\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}\right).$
This provides a first-order approximation for the mean and variance of the value of information over the subsample $s$ using the delta method.