### v3

# Linying Lv

Abstract
I examine the value of information from sell-side analysts by analyzing a large corpus of their written reports. Using embeddings from state-of-the-art large language models, I show that textual information in analyst reports explains 10.19% of contemporaneous stock returns out-of-sample, a value that is economically more significant than quantitative forecasts. I then perform a Shapley value decomposition to assess how much each topic within the reports contributes to explaining stock returns. The results show that analysts’ income statement analyses account for more than half of the reports’ explanatory power. Expressing these findings in economic terms, I estimate that early acquisition of analysts’ reports can yield significant profits. Analysts’ information value peeks in the first week following earnings announcements, highlighting their vital role in interpreting new financial data. JEL Classification: G11, G14, G24
Keywords: Sell-side Analysts, Value of Information, Large Language Models, Explainable AI
“No one ever made a decision because of a number. They need a story.”
― Daniel Kahneman
1 Introduction
Sell-side analysts play a crucial role in financial markets by producing and processing information for investors. Their research reports, which include both quantitative forecasts and qualitative analyses, are widely used by investors in decision-making (Barber et al., 2001; Kong et al., 2021). This raises a fundamental question: Do the analysts generate value for their clients, or are they merely peddling expensive noise?
Quantifying the value of analyst information, especially that contained in written reports, has been a persistent challenge in finance literature due to the unstructured nature of textual data. Prior studies have shed light on the incremental information value of report text through sentiment analyses (Asquith et al., 2005; Huang et al., 2014; Huang et al., 2018). However, analyst reports are meant to provide contextual information beyond the simple sentiment dimension. The context has been perceived at least as important as the quantifiable outputs. According to the annual survey of Institutional Investor magazine, investors consistently rank “Written Reports” as more valuable than earnings forecasts and stock recommendations. Yet, empirical evidence supporting this survey feedback is limited, and it remains unclear what specific discussion holds the most value for investors.
The advent of large language models (LLMs) has significantly enhanced our ability to quantify textual meaning beyond sentiment and topics, allowing us to capture both contextual information and the reasoning logic within the text (Chen et al., 2022; Li et al., 2024). Consider the following two statements: “Reported EPS of \$0.94 miss consensus of \$1.40 and our estimate of \$1.17.” and “Reported 1Q12 MS franchise sales are below our estimate, as sales were negatively impacted by unfavorable distribution channel dynamics.” While both sentences analyze earnings in a negative tone, the latter provides additional contextual information and reasoning. The question now becomes: do investors disregard obvious facts and figures in favor of details in lines?
I investigate the question through three general steps: analyst output representation, econometric modeling, and value decomposition. Analyst outputs include both quantitative information and qualitative information. The quantitative data consists of numerical forecasts like stock recommendations, earnings forecasts, and target prices. For qualitative textual information, I extract contextualized representations by mapping each report into a structured LLM embedding space.
With structured representation, I evaluate how well analyst information explains stock returns as evidence of information content. A significant challenge in projecting stock returns onto text embeddings is the risk of in-sample overfitting due to high dimensionality. To address this issue, I employ Ridge regressions. The simple model penalizes large coefficients while still preserving the model’s ability to capture essential signals that move the market. Additionally, I use out-of-sample R-squared to evaluate model performances.
I find that the textual information generates an out-of-sample $R^{2}$ of 10.19% for three-day cumulative abnormal returns ($CAR_{[-1,+1]}$), surpassing the 9.01% explained by numerical information. When both textual and numerical measures are combined, the explainable variation in $CAR$ increases to 12.28%, which is significantly higher than using either information type alone. These findings remain robust across various LLM text representations, machine learning algorithms, and $CAR$ windows, highlighting the distinct and economically meaningful information content embedded in analyst report narratives.
The informativeness of analyst reports varies substantially with forecast changes. Analyst reports accompanied by forecast revisions provide significantly more valuable information to financial markets compared to reiterations. For reports containing recommendation revisions, the combination of numerical and textual information generates an out-of-sample $R^{2}$ of 22.63%. In contrast, reports merely reiterating previous earnings forecasts yield statistically insignificant out-of-sample $R^{2}$ is not statistically significant, indicating they provide minimal new information to market participants.
Previous research emphasizes analysts’ role in information processing around corporate earnings announcements( e.g., Livnat and Zhang, 2012, Keskek et al., 2014, Kim and Song, 2015, Lobo et al., 2017, and Barron et al., 2017). My analysis on samples during and outside earnings announcement periods reveals that analyst reports’ information content peaks in the first week following earnings releases. The finding aligns with Huang et al. (2018), who argue that prompt responses to corporate disclosures provide valuable insights to clients. Using distinct models trained on samples during and outside earnings announcement windows, the out-of-sample $R^{2}$ approximately doubles in the first week following earnings announcement dates compared to other periods. The substantial increase highlights the incremental value analysts add by interpreting and contextualizing earnings information for investors.
To address the concern that my results merely capture analysts’ “piggybacking” on earnings conference call information or post-earning-announcement drift (PEAD), I control for both earnings surprises and the latest earnings conference call transcripts using text embeddings. The results show that analyst reports provide distinct and valuable insights even after accounting for these alternative explanations.
As raised in the review paper of Bradshaw et al. (2017), identifying the qualitative attributes of analyst reports that investors find most salient remains largely unsolved. I develop a systematic approach to quantify the importance of different report content. By exploiting the additive property of text embeddings and implementing a Shapley value decomposition, I attribute the explanatory power to 17 major topics discussed in analyst reports.
To systematically quantify the importance of different content in analyst reports, I leverage the additive feature of text embeddings and design a Shapley value decomposition approach to fully attribute the explanatory power to 17 major topics discussed in analyst reports. The results reveal that Income Statement Analyses, particularly interpretations of realized earnings, account for over half of the reports’ explanatory power in terms of out-of-sample $R^{2}$. These results reinforce the critical role of analysts in interpreting financial data and constructing the narratives and context essential for investment decisions.
To translate statistical measures of information content into economic terms, I adopt Kadan and Manela (2024)’s framework to quantify the dollar value of analyst information for strategic investors. The multi-client framework provides a realistic foundation for analyzing the complex dynamics of analyst reports dissemination. The information value can be interpreted as the total expected profits for strategic investors either because: (1) some investors have early access to analyst views, or (2) they conduct independent research and derive insights comparable to those of analysts. Essentially, the measure is defined as the ratio of the explainable return variance—which quantifies the reduction in uncertainty attributable to analyst insights—to price impact, representing the cost of trading on this information.
The analysis reveals substantial economic value in analyst information. For an average S&P 100 stock over a three-day window, strategic investors’ expected profits from early access to analyst reports amount to \$0.34 million from numerical information, \$0.38 million from textual information, and \$0.47 million when combined. Based on an average of 15 report days annually, the aggregate annualized information value reaches \$6.89 million—a conservative lower bound according to Kadan and Manela (2024). This value increases for large-cap stocks, bold analysts, and reports released promptly following earnings announcements.
My paper relates to three strands of literature. It first connects to recent studies of market reactions to security analyst reports (Huang et al., 2018; Chen et al., 2023; Chi et al., 2024). Earlier investigations primarily rely on manual text coding and sentiment analysis, which are labor-intensive and may overlook nuanced insights. My work demonstrates the superiority of state-of-the-art LLMs in generating contextualized textual representations. Compared to Huang et al. (2014)’s Naive Bayes tone classification, LLMs offer two main advantages: better capture of overall report sentiment and extraction of richer information beyond tone. Moving from simple tone analysis to comprehensive LLM embeddings boosts the $R^{2}$ from 3% to over 10%. The substantial improvement in explanatory power highlights LLMs’ potential to transform textual analysis. Additionally, the topic decomposition approach provides novel insights into analysts’ value by revealing their role in interpreting financial data.
Second, this work contributes to the growing field of machine learning and artificial intelligence in finance. Recent studies like Chen et al. (2022), Li et al. (2024), Jha et al. (2024), and Beckmann et al. (2024) have showcased the potential of implementing LLMs in a financial paradigm. However, the interpretability of LLMs remains an issue, and solutions like prompting can introduce subjectivity and manipulation concerns. By exploiting the additive property of LLM embeddings and developing a Shapley value decomposition approach, my research provides a new framework for explainable AI.
Third, this paper advances the understanding of analyst tipping - the practice of brokerage firms providing early research access to high-commission clients (Irvine et al., 2007; Christophe et al., 2010). Tipping allows brokerages to offset research costs by earning commissions from clients who benefit from the privileged information. Prior studies, such as those by Green (2006) and Kadan et al. (2018), have explored the short-term profitability generated from early access to analyst recommendation revisions, with typical returns ranging from 1.02% to 1.71%. This paper extends by quantifying the economic dollar value of such privileged information.
The rest of the paper is organized as follows. Section 2 introduces methodology, measures of analyst information value, and interpretation framework. Section 3 describes the data and empirical results. Section 4 concludes.
2 Methodology
This section describes the key methodologies: the use of large language models for topic modeling and text embeddings, estimation of information content and strategic information value, and decomposition metrics for evaluating topic-specific information value. [^1] The details of machine learning models used to explain contemporaneous stock returns are provided in the Online Appendix C.
2.1 Large Language Models
The analysis of analyst reports leverages three advanced generative AI models. A fine-tuned BERT model classifies sentences according to predefined criteria, ChatGPT extracts and synthesizes major topics through prompting, and LLaMA generates high-dimensional vector embeddings to represent report content.
2.1.1 ChatGPT Prompting
To avoid ad-hoc classification, I employ ChatGPT prompting to extract exclusive and meaningful topics. The goal is to identify primary topics covered in analyst reports and assign each sentence to a single topic category.
In the first step, I use a random sample of 100 analyst reports and request ChatGPT to analyze the content and identify high-level categories with the following prompt:
Prompt 1:Please read the provided text file of sell-side analyst reports carefully. What are high-level mutually exclusive topics covered in these reports? Make sure that each sentence from the text file can be assigned to one of the topics. Here is the report content: {text}.
It ends up with 16 categories defined in Table LABEL:tab:_topic. A category labeled “None of the Above” is included to accommodate sentences that do not fit into predefined topics. To provide a visual representation of the content within each topic, Figure A1 in the Online Appendix presents word clouds for each category. Furthermore, Table LABEL:tab:_topic_example offers illustrative sentences extracted directly from analyst reports, providing concrete examples of the discussions typical to each category.
In the second step, a systematic approach is required to assign topic classifications to a large corpus of 6,975,114 sentences. To begin with, I employ the GPT-4-Turbo model to categorize each sentence in a random sample of 17,028 sentences from 100 analyst reports into a single, most relevant topic.
Prompt 2: Please read the following sentence from a sell-side analyst report of the company {firm} ({ticker}) carefully. Determine which category of information it belongs to in the following 17 categories: {categories}. Pay attention to the sentence in the context of the report.
Output your response in JSON format.
Here is the sentence from the analyst report: {sentence}.
This labeled sample is used to fine-tune a BERT model, which then classifies all sentences in the corpus. The fine-tuned model demonstrates high reliability, achieving an accuracy score of 89% on a testing sample. The distribution of topics is relatively stable across years, as illustrated in Figure 1. Income Statement Analyses and Financial Ratios are the most frequently discussed topics in analyst reports, accounting for 17.23% and 15.65% of all sentences, respectively. These are followed by Risk Factors (8.32%), Valuation (8.10%), and Investment Thesis (7.87%), which collectively form the next tier of frequently discussed topics. Appendices and Disclosures (0.50%), Executive Summary (0.48%), and Environmental, Social, and Governance (ESG) issues (0.23%) emerge as the least frequent topics.
Several factors contribute to these observed frequencies. The low occurrence of Appendices and Disclosures can be attributed to pre-boilerplate cleaning procedures implemented in the data preparation phase. Executive Summary sections appear infrequently because they are difficult to identify at the sentence level. While ESG discussions are minimal overall, they increase notably after 2020, reflecting growing market interest in sustainability issues.
2.1.2 Text Representations
I need a structured representation of writtrn reports. Unlike traditional NLP methods, the LLM framework begins with a contextual representation of tokenized text, a process that involves several key steps. Prior to tokenization, the text undergoes normalization, which may include converting to lowercase, removing punctuation, and handling special characters. Depending on the strategy, tokenization involves breaking down text into smaller units, which can be words, subwords, or even characters. These tokens are then mapped to unique identifiers based on a predefined vocabulary learned during model training.
To illustrate this process, consider the following sentence from an analyst report on NVIDIA: “We expect to see more GPU/SmartNIC integration as next-gen workloads grow and CPUs become a bigger bottleneck in the data center.” The LLaMA tokenizer breaks this into 34 sub-word tokens, including “‘We’, ‘expect’, ‘to’, ‘see’, ‘more’, ‘GPU’, ‘/’, ‘Sm’, ‘art’, ‘N’, ‘IC’, ‘integration’, ‘as’, ‘next’, ‘-’, ‘gen’, ‘work’, ‘loads’, ‘grow’, ‘and’, ‘CPU’, ‘s’, ‘become’, ‘a’, ‘bigger’, ‘bott’, ‘l’, ‘ene’, ‘ck’, ‘in’, ‘the’, ‘data’, ‘center’, ‘.’”. This granular approach allows for nuanced interpretation of complex terms. For instance, ‘next-gen workloads’ becomes five separate tokens: ‘next’, ‘-’, ‘gen’, ‘work’, and ‘loads’.
At the heart of LLMs lies the transformer architecture. A core output of the transformer is text embeddings - dense, continuous vectors that capture semantic and syntactic information. Figure A2 in the Online Appendix illustrates how LLMs generate text embeddings for a sentence from an analyst report on Nvidia. These high-dimensional representations allow models to capture complex patterns in text, surpassing traditional word-based techniques.
For this analysis, I used the pre-trained LLaMA-2-13B model and its tokenizer from MetaAI. The input length limit of LLaMA-2 models is 4,096 tokens, which is sufficient for most analyst reports, given their median length of 1,393 tokens and mean length of 2,055 tokens. The resulting contextual embeddings capture nuanced report information and are suitable for downstream tasks like sentiment analysis and economic modeling.
2.2 Information Content
Following Asquith et al. (2005), I consider $R^{2}$ as a proxy for information content. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations annually using an expanding window:
$CAR_{[-1,+1],it}=\beta_{0}+\beta^{\prime}y_{ijt}^{\mathrm{AI}}+\epsilon_{ijt},$ (1)
[…]
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ for stock $i$ at day $t$. The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ using:
$R_{\mathrm{oos}}^{2}=1-\frac{\sum_{(i,t)\in\mathcal{T}}\left(CAR_{[-1,+1],it}-\widehat{CAR}_{[-1,+1],it}\right)^{2}}{\sum_{(i,t)\in\mathcal{T}}CAR_{[-1,+1],it}^{2}},$ (3)
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the proportion of return variance attributable to analyst information.
To assess the statistical significance of the $R_{\mathrm{oos}}^{2}$, I implement a modified version of the Diebold and Mariano (2002) test for predictive accuracy comparison. Given the strong cross-sectional dependence inherent in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional averages of prediction errors.
2.3 Information Value
Following Kadan and Manela (2024), I evaluate the value of information using a measure that approximates information value in a strategic setting where multiple informed investors possess imperfect signals:
$\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
where $\widehat{\Omega}_{it}$ represents the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price on day t-2. $\widehat{r}_{ijt}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$. Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ represents the number of analyst reports for stock $i$ on day $t$.
To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). The order flow is derived from changes in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals, with trades classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^2] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression specification is:
$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda). In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
The measure reflects the ratio of return volatility to price impact and is grounded in the theoretical framework of Back et al. (2000), which is a continuous-time extension of Kyle (1985) model. Under risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of asset price. The intuition is illustrated with an extended single-auction Kyle model with incomplete information in Online Appendix D.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance. [^3] Kadan and Manela (2024) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply in the analyst context.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To address this challenge, I apply the delta method to estimate the mean and variance of information value over sample $s$. The derivation of the first-order estimator using the delta method is detailed in Online Appendix E.
2.4 Shapley Value Decomposition
To quantify the relative contribution of different topics, I implement the Shapley Additive exPlanations (SHAP) methodology. SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. It was later applied for machine learning model interpretation (Lundberg and Lee, 2017 and Chen et al., 2022). The approach allows for attributing an analyst report’s overall explanatory power to its constituent topics. Leveraging the additive nature of text embeddings, I develop a novel decomposition method to systematically evaluate the relative importance of various topics in analyst reports.
The report-level embeddings are generated as the average of embeddings across all tokens and layers. The report embeddings can be formalized as:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ denotes the total token count of the report. Given the tokens can be classified into distinct topics, the embeddings can be fully decomposed into topic-specific components:
$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
where $P$ denotes the total number of topics, $y_{p}^{emb}$ represents the embedding for topic $p$, $N_{p}$ is the number of tokens classified as topic $p$, and $i_{p}$ indicates the index of tokens within topic $p$.
A consideration in the transformer architectures is the self-attention mechanism, which allows the model to compute relative importance weights between tokens. This architecture enables each token in the input sequence to “talk” with each other. Specifically, the embedding of token $i$ incorporates contextual information from token $j$, even when these tokens belong to distinct topics. To mitigate this cross-topic attention effect, I process embeddings sentence by sentence to ensure the token embeddings only reflect intra-sentence contextual information. I then take the token-weighted mean of sentence embeddings as the report representation, which I designate as sentence-segment embeddings:
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embeddings of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding quantifies the information content of context across sentences.
Subsequently, I compute the Shapley value across topics using sentence-segmented embeddings. The Shapley value is defined as:
$\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
where the sum extends over all subsets $S\subseteq P\backslash\{p\}$, including the empty set. $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. The $R_{\text{oos}}^{2}$ is calculated for each unique combination of topic embeddings.  [^4] Shapley value requires a defined input for missing $X$ values. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ across topics, revealing their relative contributions to explaining market reactions.
3 Data and Empirical Results
In this section, I present the empirical analysis of analyst report information, beginning with an examination of information content and its decomposition across topics. I then quantify the strategic dollar value of analyst information. The results emphasize analysts’ income statement analyses following earning announcements.
3.1 Data
I refer to the Mergent Investext database for 223,091 sell-side analyst reports for S&P 100 firms spanning the years 2000 to 2023. The reports, initially in PDF format, are converted to plain text using Adobe Acrobat. The extracted text is then segmented into sentences. To ensure a clean dataset focused on firm-relevant discussions, I fine-tune a BERT model to filter out boilerplate content.  [^5] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I construct a training dataset by randomly sampling one report from each of the top 20 brokerage firms by report volume and manually label each sentence. The BERT model fine-tuned for boilerplate classification achieves a 97.31% accuracy rate on test data.
The Mergent Investext Database provides analyst identification and report release dates. To connect Investext with the I/B/E/S database, I first match lead analyst names to their corresponding Analyst IDs (AMASKCD) following Li et al. (2024). I then match individual reports to corresponding EPS forecasts, price targets, and recommendation files in I/B/E/S. Given that multiple reports may be issued between a forecast announcement and its subsequent revision, I adopt Huang et al. (2014)’s approach and define the matching window as spanning from two days before the announcement date to two days after the revision date. [^6] I/B/E/S records two timestamps for each analyst forecast: the announcement date (ANNDATS) when the new forecast is issued and the old one is revised, and the revision date (REVDATS) when the forecast becomes invalid or undergoes revision. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. Additional firm characteristics are sourced from CRSP and Compustat. [^7] Detailed variable definitions and summary statistics are provided in Table A2 and Table A3, respectively. After the matching procedure and excluding reports released on earnings announcement days, the final sample consists of 122,252 reports covering S&P 100 stocks, produced by 1,305 analysts across 140 distinct brokerage firms.
Table 2 provides summary statistics of analyst reports, including the number of reports, unique brokerage firms, and analysts across years and the Fama-French 12 industries. It also reports average page and token counts per report as measures of report length. The temporal distribution of analyst reports aligns with patterns documented in Bonini et al. (2023), while the average page count per report corresponds with findings in Huang et al. (2014). [^8] The data reveals a peak in the number of analyst reports in 2013, followed by a gradual decline. This trend is likely influenced by the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and MiFID II in Europe (Bradshaw et al., 2017).
Stock-level price impacts are estimated using intraday data from the NYSE TAQ database for an out-of-sample period spanning January 2, 2015, to December 28, 2023. To ensure temporal comparability, all prices are adjusted for inflation using the Consumer Price Index (CPI), normalized to December 2020 as the base period.
To isolate the information value of analyst outputs, I control for alternative information sources available to investors. In Section 3.2.2, I conduct a comparative analysis between earnings conference call content and subsequent analyst reports. Conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts.. These transcripts provide detailed documentation of executive presentations and manager-analyst interactions during Q&A sessions. I sample transcripts of S&P 100 constituent firms from 2005 to 2023, matching them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks, yielding a final sample of 7,909 firm-quarter-level conference calls.
The set of numerical variables used as input in machine learning models includes:
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 days pre-release. $Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both. $SR$ denotes stock recommendations following the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell). $ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 days pre-release. $Prior\_CAR$ measures cumulative abnormal return over [-10, -2].
Firm-Level Measures: $Size$ is quantified by market equity. Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value. Following Huang et al. (2018), Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value. $Miss$ is a binary indicator for positive $SUE$. $TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding. Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market fluidity.
Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the firm’s industry (defined by Fama-French 48 classification) experiences negative monthly returns in the bottom quintile.
Macroeconomic Measures: $TimeTrend$ is a temporal variable measuring the number of years since the start of the sample period.
3.2 Information Content of Analyst Reports
This section investigates market reactions to analysts’ quantitative and qualitative outputs. I employ the Shapley value decomposition methodology introduced in Section 2.4 to attribute the information content across distinct topics.
3.2.1 Quantitative and Qualitative Information
Table 3 compares the information content embedded in numerical and textual measures. The analysis employs Ridge regression as specified in Equations 1 and 2, with ${CAR}_{[-1,+1]}$ centered at the report release date serving as the dependent variable. For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database. The estimation utilizes distinct input configurations with an expanding window approach, where training sets span from 2000 to the year preceding each target year.
Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for forecast revision input, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst revisions. The incorporation of additional 14 numerical features in column (2) yields a modest improvement to 9.08% $R^{2}_{\mathrm{OOS}}$. Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the Diebold-Mariano test t-statistics reported in columns (2) and (4).
Columns (5) and (6) reveal that Ridge models with text embeddings achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text. The $R^{2}_{\mathrm{OOS}}$ remains stable from 2015 to 2023, except during the 2020 pandemic recession. A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating substantial information content of report text relative to forecast revisions.
Addressing the look-ahead bias concern raised by Sarkar and Vafa (2024) regarding pre-trained LLMs, I examine the $R^{2}_{\mathrm{OOS}}$ in 2023. Since the employed LLMs were trained on pre-2023 data, this period represents true out-of-sample performance, validating the robustness of the main findings.
Columns (7) and (8) examine the model performance when integrating both forecast revisions and text embeddings. This combined approach achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively. The DM t-statistics of 3.95 and 3.77 for comparisons between combined input versus revision input and text input demonstrate that report text and forecast revisions contain distinct, complementary information sets valued by investors.
The explanatory power might stem from quantitative information embedded in the numbers mentioned within the text. To address this concern, I conduct robustness checks using content with numbers removed. Table A4 presents the performance of models using number-free text embeddings. The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% significantly exceeds that of quantitative revision measures (t-statistic = 2.40), suggesting that the superior performance of text-based models is not driven by embedded numerical content.
I examine the scaling law relationship by comparing LLMs with varying parameters, including BERT, OpenAI text-embedding-3-small, and LLaMA-3 models. The LLaMA models feature significantly more parameters than the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 benchmark. The input in Panel B is report text embeddings, comparable to column (3) in Panel A. A positive t-statistic indicates superior performance relative to the LLaMA-2 model. While smaller models like BERT and OpenAI underperform LLaMA-2, the larger LLaMA-3 model shows only marginally better performance with an $R^{2}_{\mathrm{OOS}}$ of 9.66%, though the improvement is statistically insignificant. These results demonstrate the robustness of analyst report text information content across different LLMs and illuminate the relationship between model scale and performance capabilities.
To validate the robustness of contemporaneous measures, I examine alternative $CAR$ windows, including 2-day and 5-day intervals following Huang et al. (2014) and Asquith et al. (2005), respectively. Table A5 shows consistent $R^{2}_{\mathrm{OOS}}$ magnitudes for $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% during the 2015-2023 out-of-sample period.
Table A6 presents the cross-industry analysis of analyst report text information content across Fama-French 12 classifications. Five industries exhibit $R^{2}_{\mathrm{OOS}}$ exceeding 10%: Shops, Other, Manufacturing, Chemicals, and Durables. Conversely, Non-Durables, Telecommunications, Energy, and Utilities sectors show significantly lower $R^{2}_{\mathrm{OOS}}$ and DM t-statistics. This pattern, visualized in Figure A3, suggests that industries with complex products and services derive greater value from detailed analyst insights, while those operating in stable or regulated environments benefit less from incremental analyst information.
To evaluate the economic significance, I estimate OLS regressions specified as follows:
$CAR_{[-1,+1],it}=\alpha+\beta_{1}\widehat{CAR}_{txt,it}+\beta_{2}\widehat{CAR}_{rev,it}+\epsilon_{it}.$ (10)
This model incorporates out-of-sample $CAR_{[-1,+1]}$ estimations using either report text embeddings or forecast revisions as input. The estimates for 2015, for example, are derived from Ridge regression models trained on 2000-2014 data. The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension, while $\widehat{CAR}_{rev}$ is estimated using three forecast revision measures under identical procedures.
Table 4 presents sample summary statistics and regression results. The analysis sample is restricted to observations with valid current and previous stock recommendations, earnings forecasts, and target prices. Panel B’s Column (1) confirms that recommendation revisions, earnings forecast revisions, and target price revisions all contribute significantly to explaining contemporaneous abnormal stock return variations, consistent with Brav and Lehavy (2003) and Huang et al. (2014). Column (2) shows that $\widehat{CAR}{txt}$ alone achieves an $R^{2}$ of 10.5%. When both analyst revision measures and $\widehat{CAR}{txt}$ are included as regressors in column (3), the $R^{2}$ increases to 14.9%. In economic terms, one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness corresponds to 3-day abnormal return increases of 10, 40, 90, and 120 basis points, respectively. Columns (4) and (5) examine $\widehat{CAR}{rev}$, which aggregates all revision metrics. The $\widehat{CAR}{rev}$-only model achieves an $R^{2}$ of 8.9%, while combined inputs reach 14.7%. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
Columns (6)-(7) introduce $SUE$ and $Prior_{C}AR$ controls to capture earnings surprises and recent news. The stability of $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ coefficients suggests that market-valued analyst information is distinct from recent earnings news and does not merely reflect reactions to past returns.
3.2.2 Earnings Announcements
Building on research examining the interplay between analyst reports and corporate disclosures (Frankel et al., 2006; Chen et al., 2010; Huang et al., 2018), I analyze the information content of sell-side analyst reports surrounding earnings announcements.
Following Chen et al. (2010), I group analyst reports into 13 weekly bins around earnings announcement dates. Panel A in Table 5 presents the $R^{2}_{\mathrm{OOS}}$ and DM t-stats for weekly report samples. Consistent with Huang et al. (2018), analyst reports cluster predominantly in the immediate post-earnings announcement period. The information content only exhibits strong significance only in the first week, achieving an $R^{2}_{\mathrm{OOS}}$ of approximately 10% with a t-statistic of 10.17.
The variation in performance across weekly bins may reflect Ridge models capturing dominant patterns in analyst reports issued around earnings announcements. To address potential temporal variation, I train separate Ridge models on analyst reports released during earnings announcement periods and those issued beyond these periods, allowing the models to extract distinct information content based on report timing. Panel B in Table 5 compares the model performance within and beyond earnings announcement periods using 1-day, 2-day, 3-day, and 7-day windows. Models trained on earnings announcement periods consistently outperform those trained on non-announcement periods. The $R^{2}_{\mathrm{OOS}}$ and t-statistics exhibit a declining pattern as windows lengthen, decreasing from 11.84% to 7.29% over the earnings announcement week, suggesting markets particularly value swift analyst responses to earnings news.
To address concerns that analysts merely restate earnings call information and that market reactions primarily reflect earnings announcements rather than analyst insights, I conduct a comparative analysis of information content between these two sources. I focus on reports released one day after earnings announcements, when analysts would be most likely to simply repackage earnings information.
Column (1) of Panel C presents the $R^{2}_{\mathrm{OOS}}$ for $CAR_{[-1,+1]}$ using earnings conference call transcript embeddings as input. [^10] Due to the length of earnings conference call transcripts (median 13,029 tokens), I employ the LLaMA-3-8B model for text embeddings in Panel C, which accommodates 8,192 tokens versus the 4,096-token limit of the LLaMA-2-13B model used elsewhere. Texts are truncated at these limits to maximize information retention. The $R^{2}_{\mathrm{OOS}}$ reaches 4.20% with a DM t-statistic of 5.16. Since the conference call transcript data began in 2005, I retrained the Ridge regression model with analyst report text input to align the sample periods. Using report text embeddings from the day following earnings announcements, this analysis achieves an $R^{2}_{\mathrm{OOS}}$ of 9.72% (column 2), substantially outperforming conference call transcripts alone.
The combination of analyst report and earnings conference call embeddings yields further improvements, with the $R^{2}_{\mathrm{OOS}}$ increasing to 11.96% (column 3). A comparison t-statistic of 5.12 between transcript-only and combined models demonstrates analysts’ substantial incremental contribution. These findings collectively refute the notion that analysts merely restate earnings information, instead highlighting their distinct and timely value to the market.
3.2.3 What Content in Analyst Reports is Valued?
Having established the significance of report text information, I examine which specific content components most influence market response. I employ the Shapley value methodology to decompose the text-based $R^{2}_{\mathrm{OOS}}$ across the 17 topics detailed in Table LABEL:tab:_topic. The analysis uses sentence-segmented embeddings as report representations to mitigate cross-topic contextual effects. When switching from full-context to sentence-segmented embeddings, the $R^{2}_{\mathrm{OOS}}$ decreases by 2.25%, quantifying the magnitude of inter-sentence contextual information.
Figure 2 illustrates the Shapley value decomposition, revealing a clear hierarchy in topic importance. Income Statement Analyses and Financial Ratio Analyses emerge as the most valuable components, contributing 67% and 45% of text-based $R^{2}_{\mathrm{OOS}}$, respectively. Investment Thesis and Valuation form a second tier of importance, with Shapley values of 1.76% and 1.51%. The remaining topics show minimal or negative Shapley values, indicating limited market reactions.
The findings’ robustness is supported by several additional analyses. Figure A4 shows that normalizing Shapley values by sentence count and token length maintains Income Statement Analyses’ dominance. Figure A5 demonstrates consistency in topic rankings across different industries.
The evolution of text-based $R^{2}_{\mathrm{OOS}}$ is depicted in Figure A6 of the Online Appendix. While a notable decline in $R^{2}_{\mathrm{OOS}}$ occurs between 2020 and 2021, likely reflecting COVID-19 pandemic disruptions in financial markets, the scaled Shapley value for Income Statement analysis (shown by the red line) exhibits remarkable stability. This persistence suggests that analyst commentary on income statements maintains its informational value even during market turbulence.
To disentangle whether topic importance stems from analysts’ expertise in certain areas (supply effect) or investors’ response (demand effect), Figure A7 compares Shapley values across analyst experience levels and brokerage firm sizes. The consistent topic hierarchy across all four subsamples (experienced/inexperienced analysts and large/small brokerages) indicates that topic rankings are not driven by variations in analyst expertise.
I proceed to examine analysts’ dual roles in income statement analysis: their discovery role in collecting and generating otherwise unavailable information and their interpretation role in integrating and processing available information into more interpretable signals (Huang et al., 2018; Blankespoor et al., 2020). To facilitate the analysis, I categorize Income Statement sentences based on: (1) Information Acquisition versus Information Interpretation, and (2) Realized Income versus Expected Income. This categorization is achieved through a prompt-based classification system, as detailed in Prompt 3.
Prompt 3: Please read the following sentence from a sell-side analyst report for the company {firm} ({ticker}) and classify it based on two criteria:
Information Type:
Information Acquisition (0): Sentences that directly report quantitative financial data, such as earnings, revenue, expenses, or other metrics. Example: ‘Reported EPS of \$1.40 beat consensus of \$0.94 and our estimate of \$1.17.’
Information Interpretation (1): Sentences that provide analysis, context, or interpretation of the financial data, such as discussing trends, comparing performance to forecasts, assessing market impacts, or considering strategic implications. Example: ‘Biogen reported 1Q12 MS franchise sales below our estimate and the Street, as Avonex sales were negatively impacted by unfavorable distribution channel dynamics.’
Time Reference:
Realized Income (0): Sentences referring to concrete, historical results from completed periods. Example: ‘The company’s revenue for the fiscal year 2022 was \$500 million.’
Expected Income (1): Sentences containing subjective predictions, or expectations for future periods. Example: ‘The company expects to achieve revenue of \$550 million in fiscal year 2023 based on current market conditions.’
Output your classification as two comma-separated numbers: the first for Information Type (0 for Acquisition, 1 for Integration) and the second for Time Reference (0 for Actual, 1 for Forecast).
Output format: InfoType, TimeRef
Sentence to classify: {report_sent}.
Figure 3 provides a granular breakdown of $R^{2}_{\mathrm{OOS}}$ across 18 topics, including Income Statement subcategories. Income interpretation contributes three times as much to $R^{2}_{\mathrm{OOS}}$ as income acquisition, suggesting that markets value context and explanation over raw figures. The contribution from realized income discussions is triple that of income expectations, indicating markets place greater value on actual results than forward-looking projections. These patterns emphasizes that markets particularly value analysts’ ability to interpret realized income, beyond mere information reporting or future expectations.
These findings align with the results from section 3.2.2, where analyst reports released immediately after earnings announcements demonstrate the highest information content. Taken together, these results suggest that the market places the highest value on analysts who can rapidly digest earnings announcement information and provide insightful interpretations.
3.2.4 Comparison of Revision Reports and Reiteration Reports
Given the established relationship between market prices and analyst forecast revisions (e.g., Gleason and Lee, 2003; Asquith et al., 2005; Bradshaw et al., 2021), I examine the differential information content between revision and reiteration reports. Following the criteria of Huang et al. (2014), I classify reports as revisions when released within two days of the I/B/E/S forecast announcement date. This classification is applied independently across three forecast targets—recommendations, target prices, and earnings. Therefore, a single report may be classified as both an earnings forecast revision report and a recommendation reiteration report.
Table 6 presents out-of-sample model performance across reiteration and revision samples. The frequency of revisions varies considerably: earnings forecast revisions are most common (56.57% of reports), followed by target price revisions (31.10%), while recommendation revisions are relatively rare (2.29%). Recommendation revisions exhibit particularly strong explanatory power. Quantitative forecast revision measures achieve an $R^{2}_{\mathrm{OOS}}$ of 14.99% when accompanying recommendation revisions, while report text yields 16.80%. The combination enhances explanatory power to 22.63%. This finding complements the literature that attributes large stock price changes to analysts’ recommendation revisions (Loh and Stulz, 2011; Bradley et al., 2014).
Target price revisions, noted for their incremental information value due to the limitations of discrete recommendations (Bradshaw, 2002; Brav and Lehavy, 2003), show marked differences between revisions and reiterations. While reiteration reports demonstrate minimal information content (combined $R^{2}_{\mathrm{OOS}}$ of 3.58%), revision reports exhibit substantial value ($R^{2}_{\mathrm{OOS}}$ = 20.88%, t-stat = 11.18).
Earnings forecast revisions show a similar pattern. Reiteration reports yield insignificant $R^{2}_{\mathrm{OOS}}$, while revision reports demonstrate significant explanatory power. The combination of numerical and textual inputs for revision reports achieves an $R^{2}_{\mathrm{OOS}}$ of 16.37% (t-statistic = 11.42), highlighting revisions’ superior value over reiterations.
Overall, Table 6 shows that report content carries greater information content when accompanying forecast revisions rather than reiterations, with recommendation revision reports exhibiting the highest value.
3.2.5 Comparison with Analyst Report Sentiment
The sentiment analysis serves two objectives: to evaluate how LLMs perform compared to traditional NLP methods, and to assess the relative information content of embeddings versus tone measures. I compare tone measures derived from both Huang et al. (2014)’s Naive Bayes approach and BERT classification. I manually label 10,000 sentences as training data for both models. After training the models to categorize sentences as positive, neutral, or negative, I average these classifications at the report level to create $Tone\_NB$ and $Tone\_BERT$ measures.
Table 7 contrasts the performance of text tone measures against text embeddings. Panel A presents OLS regression results of three-day $CAR$ on tone measures. Column (1) shows a coefficient estimate of 0.022 for $Tone\_NB$, aligning with Huang et al. (2014)’s reported 0.0208. A one standard deviation increase in text tone favorableness corresponds to a 54 basis point increase in $CAR$, comparable to their reported 41 basis point increase in two-day abnormal returns. Column (2) reveals significantly positive coefficients for tone measures for both Income Statement and non-Income Statement topics. The results remain consistent under firm-fixed and year-fixed effects in column (3). BERT-based tone measures (columns 4-6) demonstrate substantially larger coefficient estimates with higher significance levels. A one standard deviation increase in BERT-based tone favorableness yields a 95 basis point increase in three-day $CAR$. When both measures are included (column 7), Naive Bayes-based tone coefficients become insignificant, indicating BERT’s superior sentiment capture.
Panel B of Table 7 compares the $R^{2}_{\mathrm{OOS}}$ across models. Tone inputs alone achieve $R^{2}_{\mathrm{OOS}}$ ranging from 0.05% to 3.68% in the first four rows. Including forecast revisions and text embeddings improves $R^{2}_{\mathrm{OOS}}$ to 9.49% and 12.27%, respectively. The panel’s final rows show that adding tone measures to text embedding models yields no statistically significant improvement in $R^{2}_{\mathrm{OOS}}$.
The results highlight two key advantages of LLMs over conventional NLP methods: stronger explanatory power for stock returns through better sentiment capture, and the ability to extract richer textual information beyond simple sentiment measures.
3.2.6 Alternative Machine Learning Models
I examine the information content using three alternative machine learning algorithms, following the same procedure of Ridge regressions. Results are presented in Table 8.
PLS regressions extract common components that correlate with the target variable. If text embeddings contain strong common components that correlates with stock returns, PLS should outperform. Using this method, I achieve an $R^{2}_{\mathrm{OOS}}$ of 8.26%, with year-to-year variations similar to Ridge models. DM tests confirm significance at the 1% level across all years.
The XGBoost algorithm, which builds sequential trees to combine weak estimations into stronger ones, allows for testing whether text embeddings contain discrete informational components of varying importance. Despite its ability to capture nonlinear relationships, XGBoost underperforms Ridge regressions with an $R^{2}_{\mathrm{OOS}}$ of 5.97% (DM t-stat = 6.76). This relatively weaker performance suggests that linguistic information is distributed across the embedding space rather than concentrated in specific dimensions.
Neural Network models, which underpin LLMs and transformer architectures, show superior performance. Given their interconnected node structure mirroring human neural networks, these models are well-suited to capture word relationships and contextual information. Using ensembles of five models for each configuration, Neural Networks achieve the highest $R^{2}_{\mathrm{OOS}}$ among tested algorithms, ranging from 10.42% to 12.12%. Performance improves with network depth, consistent with Kelly et al. (2022)’s virtue of complexity principle.
Two key findings emerge from these experiments. First, the significant $R^{2}_{\mathrm{OOS}}$ across models demonstrates meaningful information content in report text, particularly given the low signal-to-noise ratio in stock returns. Second, although Neural Networks achieve the highest performance, most of the textual information can be captured by simple linear models like Ridge regressions.
3.3 Information Value of Analyst Reports
To translate the statistical measure $R^{2}_{\mathrm{OOS}}$ into economic terms, this section quantifies the information value of analyst reports using the methodology established in Section 2.3. This measure captures the economic profit generated from reduced price uncertainty, while incorporating price impact considerations. The measure’s construction implies that both higher information content and lower price impact contribute positively to the estimated information value.
Table 9 presents the statistical evidence estimated with the delta method. For S&P 100 constituent stocks, sell-side analysts generate an average three-day window information value of \$0.47 million (SE = \$0.05 million). This value can be decomposed into its constituent components: textual content contributes \$0.38 million, while numerical revisions account for \$0.34 million. Extrapolating from the typical frequency of analyst coverage—approximately 15 report days per year—the annualized lower bound investors would expect to profit from early access to analyst reports on an S&P 100 firm amounts to \$6.89 million. The precision of these estimates is supported by consistently narrow standard errors and confidence intervals across different information value measures. The average price impact ($\widehat{\lambda}_{it}/P_{it_{-}}$) is 0.34, suggesting that a \$1 million order pushes the stock price up by 0.034%. The statistics demonstrate that analyst information value is economically meaningful, with both textual content and numerical forecasts contributing significantly.
Figure 4 illustrates the dynamics of analyst report information value. The analysis tracks both the strategic value of combined numerical and textual information (solid line) and the isolated value of textual information (dashed line). While both measures exhibit substantial temporal variation, they demonstrate an overall upward trend. The consistently higher combined value relative to textual value suggests complementarity between numerical and textual components.
The 95% confidence interval reveals increased variability in information value post-2020, characterized by wider confidence bounds and more pronounced peaks and troughs. The heightened volatility coincides with the COVID-19 pandemic period, which introduced substantial market uncertainty. The patterns demonstrate that the dynamic nature of analyst information value and its sensitivity to market conditions.
The Online Appendix provides additional robustness checks. Table A7 presents alternative information value estimates using high-frequency realized return volatility measures and varied trade-signing algorithms, yielding estimates ranging from \$0.42 million to \$0.64 million. Figure A8 demonstrates that these alternative specifications produce consistent temporal patterns in information value evolution.
3.3.1 The Value of Analyst Information in the Cross-section of Stocks
Analyst information value may vary in the cross-section of stocks. Prior literature suggests that information processing and dissemination costs vary with firm size (Zeghal, 1984), and analyst opinions may trigger larger market reactions for smaller stocks due to their relatively limited public information environment (Stickel, 1995). Figure A9 illustrates this relationship through a scatter plot of information value against stock market capitalization. While the plot reveals a positive association between market equity and information value, the relationship exhibits notable nonlinearities. To formally examine this relationship, I estimate regressions of information value on stock size, incorporating firm and year fixed effects and the book-to-market ratio control.
Table 10 shows that stocks with larger market equity have significantly higher information value from analysts. The economic magnitude is substantial: estimates from column (4) indicate that a 1% increase in market equity corresponds to a 0.864% increase in analyst information value. Columns (5)-(8) reveal that textual information value exhibits greater size sensitivity compared to the combined numerical and textual information value. To better understand the mechanisms driving this positive size elasticity, I decompose the information value into two components: explained return volatility and price impact. The decomposition can be expressed logarithmically as:
$\log\widehat{\Omega}_{it}=\underbrace{\log\left[r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right]}_{\text{log explained return variance}}-\underbrace{\log\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}}_{\text{log price impact}}.$ (11)
Panel B of Table 10 shows that higher analyst information value in larger stocks stems primarily from their greater liquidity. While analyst reports do not better explain return volatility in large stocks, the reduced price impact of liquid stocks enhances their information value.
3.3.2 The Value of Analyst Information in the Cross-section of Analysts
Prior literature suggests that analysts who make bold forecasts incorporate private information more comprehensively than those who herd (Clement and Tse, 2005). This section examines the relationship between information value and analysts’ herding behavior.
To analyze cross-sectional variation in analyst characteristics, I estimate regressions of stock-analyst-day information value on forecast boldness indicators, incorporating firm, analyst, and year fixed effects. Following Jacob et al. (1999), who suggests larger brokerage firms provide superior research, I include broker size as a control variable. The analyst-specific information value measure is defined as:
$\widehat{\Omega}_{ijt}=\frac{r_{it}^{2}-\left(r_{it}-\widehat{r}_{ijt}\right)^{2}}{\lambda_{it}}\cdot p_{it_{-}}.$ (12)
Table 11 presents the regression results. Panel A, columns (1)-(4), reveals significantly positive coefficient estimates on $Bold$, indicating more strategically valuable information when analysts provide bold forecasts. Quantitatively, information from bold reports demonstrates 29% higher value compared to herding reports—a magnitude that persists when examining textual content alone. The insignificant coefficients for $Brokersize$ suggest that broker size is related to analyst information value.
To understand the underlying mechanisms, Panel B decomposes the relationship between analyst boldness and information value into informativeness and price impact components. As shown in columns (1)-(4), information from analysts who issue bold forecasts explains higher return volatility, supporting the argument of Clement and Tse (2005) that analysts issue bold forecasts when possessing relevant private information. Columns (5)-(8) reveal no statistically significant relationship between boldness and price impact, suggesting that the enhanced value of bold forecasts stems primarily from their information content rather than market microstructure effects.
3.3.3 The Value of Analyst Information around Earnings Announcements
Previous evidence underlines the information content of analyst reports in the first week following earnings announcements. Accordingly, this section evaluates the information value of analyst reports issued promptly following earnings announcements and beyond that.
Figure 5 examines different samples of analyst reports categorized by their timing relative to earnings announcements, from reports released in the first week after earnings announcements to those issued in the $13^{th}$ week. Reports issued in the first post-earnings week exhibit the highest information value, with a sample mean of \$0.84 million. Reports from subsequent weeks show a sharp decline in value. This pattern again suggests that markets place premium value on timely post-earnings analyst reports.
To formally test the observed pattern, I estimate regressions of log information value on an indicator variable $Week$, which identifies reports released within one week after earnings announcements. Following Banerjee (2011) and Beckmann et al. (2024), I include earnings announcement day trading volume ($TradingVolume$) as a proxy for investor disagreement. The model incorporates an interaction term ($Week\times TradingVolume$) to capture the joint effect of post-earnings timing and investor disagreement on analyst information value. The specification builds on Chen et al. (2010)’s finding that analysts’ interpretative role becomes more critical with increasing financial information complexity. The model includes stock and year fixed effects.
Table 12 presents regression results. Panel A documents significantly higher information value for reports issued within the first post-earnings week, as evidenced by positive and significant coefficients on $Week$ across all models. The positive and significant interactions between $Week$ and $TradingVolume$ indicate that information value increases with trading volume on earnings announcement days. The results are consistent with investors placing greater reliance on analyst interpretations during periods of heightened disagreement.
Panel B decomposes the earnings announcement effects into return variance and price impact, showing significant increases in both components. The coefficient magnitudes indicate that the timing effect is primarily driven by higher explained return volatility, rather than changes in price impact. These findings lead to conclude that investors value the most of analysts’ interpretation role during periods of heightened post-earnings uncertainty.
4 Conclusion
This paper examines both the quantitative and qualitative content of analyst reports. I find that report text contains more market-valued information content than quantifiable forecast revisions economically, especially shortly following earnings announcements. A Shapley value decomposition identifies analysts’ interpretation of income statements as the most valuable component of written reports.
I further translate the information content into economic terms. Using a measure of explained return volatility relative to price impact, I find that early acquisition of the analyst reports for an average S&P 100 constituent firm yields an annualized profit larger than \$6.89 million. Cross-sectional analysis reveals that the information value from analysts is higher for large stocks, bold forecasts, and reports released in the week following earnings announcements.
These findings have important implications for regulators, particularly concerning the fair distribution of information in financial markets. The substantial potential profits from early access to analyst reports raise concerns about selective distribution practices—a direct violation of FINRA Rule 2241, which mandates research objectivity and prohibits preferential distribution to certain clients. This study presents a hypothetical scenario for quantifying information value rather than an assessment of market practices. By documenting the magnitude of potential profits from privileged access, my research emphasizes the importance of regulations for fair disclosure.
A. Additional Figures
(a) (1) Executive Summary
(b) (2) Company Overview
(c) (3) Industry Analysis
(d) (4) Competitive Landscape
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (5) Income Statement Analysis
(b) (6) Balance Sheet Analysis
(c) (7) Cash Flow Analysis
(d) (8) Financial Ratios
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (9) Business Segments
(b) (10) Growth Strategies
(c) (11) Risk Factors
(d) (12) Management and Governance
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
(a) (13) ESG Factors
(b) (14) Valuation
(c) (15) Investment Thesis
(d) (16) Appendices and Disclosures
Figure A1: Word Clouds of Topics
This figure presents word clouds for 16 key topics frequently discussed in analyst reports. Each word cloud visually represents the most common terms associated with the topic, with word size indicating term frequency. The topics include Executive Summary, Company Overview, Industry Analysis, Competitive Landscape, Income Statement Analysis, Balance Sheet Analysis, Cash Flow Analysis, Financial Ratios, Business Segments, Growth Strategies, Risk Factors, Management and Governance, ESG Factors, Valuation, Investment Thesis, and Appendices and Disclosures. The “None of the Above” category is excluded from the visualization.
Figure A2: Large Language Model Architecture
This figure shows an example of how large language models process a sentence from an analyst report on Nvidia. The model breaks this sentence into tokens, and then generates two types of embeddings: token embeddings and position embeddings. These embeddings then pass through the core of the model - a stack of transformer layers. The output of each transformer layer is a series of vectors representing the tokens in their full context.
Figure A3: Analyst Information Value by Industry
This figure shows the estimated information value of analyst reports for Fama-French 12 industries from 2015Q1 to 2023Q4, quantified in millions of dollars. The solid green line marks the sample mean value of information, with green bars showing the 95% confidence intervals. The horizontal dashed red line marks the sample mean of stocks from all industries. Information value estimates are derived using the delta method and adjusted to 2020 levels using the Consumer Price Index (CPI).
Figure A4: Shapley Values Scaled by Length
This figure presents the SHAP values of topics scaled by the total number of sentences (Panel A) and the total number of tokens (Panel B) for topics with more than 100,000 sentences. The scaling process involves two steps: First, normalize the Shapley value of each topic by dividing it by the number of sentences in that topic. Then, adjust the normalized values to ensure the total sum of SHAP values remains consistent. This ensures that the scaled values reflect the relative importance of topics while preserving the magnitude of total $R^{2}_{\mathrm{OOS}}$.
Figure A5: Shapley Values for Topic Importance across Industries
This figure shows the relative importance of 17 topics in analyst reports across the Fama-French 12 industries, quantified using the SHAP value. For each industry, the bars represent the SHAP values of different topics, with their sum equaling the $R^{2}_{\mathrm{OOS}}$ of the sentence-segmented report embeddings. The red line shows each topic’s SHAP value as a proportion of the total $R^{2}_{\mathrm{OOS}}$.
Figure A6: $R^{2}_{\mathrm{OOS}}$ of Sentence-Segmented Embeddings by Year
This figure shows the $R^{2}_{\mathrm{OOS}}$ of sentence-segmented report embeddings from 2015 to 2023. The color gradient indicates the magnitude of $R^{2}_{\mathrm{OOS}}$. The red line represents the SHAP value of the Income Statements Analyses topic across the years. The shaded area marks the pandemic recession in 2020.
Figure A7: Shapley Values for Topic Importance by Analyst Characteristics
This figure illustrates how 17 different topics contribute to the predictive power of analyst reports across four subsamples, measured using SHAP values. The subsamples compare analysts with above-median (Panel A) versus below-median work experience (Panel B), and large (Panel C) versus small brokerage firms (Panel D), with subsamples determined by yearly median. For each subsample, the bars show the SHAP values of individual topics, whose sum equals the out-of-sample R-squared ($R^{2}_{\mathrm{OOS}}$) from the sentence-segmented report embeddings. The red line indicates each topic’s relative importance, calculated as the topic’s SHAP value divided by the total $R^{2}_{\mathrm{OOS}}$.
Figure A8: Alternative Analyst Information Value Estimations Over Time
This figure presents the alternative estimations of analyst information value from 2015Q1 to 2024Q4, reported in millions of dollars and adjusted to 2020 dollars. Quarterly mean approximations are calculated using the delta method. TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Figure A9: Analyst Information Value by Stock
This figure shows the estimated information value of analyst reports for individual stocks from 2015Q1 to 2023Q4, quantified in millions of dollars. Estimates are derived using the delta method. Market equity, averaged daily at the time of each report release, is reported in billions of dollars. All values are adjusted to 2020 levels using the Consumer Price Index (CPI).
B. Additional Tables
Table A1: Topic-Specific Sentence Examples from Analyst Reports
Topics          Examples
Executive Summary
Our key takeaways from CEO Jeff Immelt’s presentation at EPG were: Outlook for substantial EPS growth over 2010/12 driven by abatement of credit losses (\$8-9bn in 2010E tapering to \$4bn run-rate) and CRE impair. Looking ahead, guidance was tightened, essentially framing the Street, and commentary suggests a strong outlook for the balance of 2007. Key topics: 1) Update on ABTÕs portfolio of COVID-19 tests; 2) Libre trends, Libre 2 launch and expectations for Libre 3; 3) Elective surgery trends exiting Q1, expectations for 2021 and an update on recent and upcoming new product approvals; and 4) Global trends and impact on EPD and Nutrition.
Company Overview
Oracle Corporation, founded in 1977 and headquartered in Redwood Shores, California, is one of the largest and most prominent companies in the software space – and a technology bellwether. As it has grown, Microsoft has expanded into enterprise software with Windows Server, SQL Server, Dynamics CRM, SharePoint, Azure and Lync; hardware with the Xbox gaming/media platform and the Surface tablet; and online services through MSN and Bing. Altria Group, Inc., is the world’s largest producer and marketer of consumer products, and had revenues of \$80 billion in 2002.
Industry Analysis
According to our global Immunology market model, US Psoriasis (PsO) represented a \$7.7B market in 2016 and is expected to grow at a low-teens CAGR to \$11.8B in 2019E and  \$13B by 2021E driven by more highly effective therapies. Despite record prices, oil demand continues to grow, while supply growth lags and spare production and refining capacity is almost nonexistent. Add to that that some new advertising expectations from PricewaterhouseCoopers of a decline in ad spending of 12% worldwide and 15% in the U.S. for 2009 and continue to decline in 2010.
Competitive Landscape
According to Mercury Research, NVIDIA is now the 3rd largest chipset supplier (consisting of desktop and mobile chipsets, and integrated and non-integrated chipsets), shipping 5.4 million units in calendar Q3 for an 8.2% market share, versus Intel’s shipments of 51 million units (62% share), VIA’s shipments of 9.6 million units (14.4% share), SiS’s shipments of 5.3 million units (8% share), and ATI’s shipments of 4.4 million units (6.6% share). Further, competition in the CDK-4/6 space is rising with Verzenio (abemaciclib) & Kisqali launches placing downward pressure on Ibrance trajectory. We believe that Accenture has recognized that web services could compete directly with client/server as the new systems architecture.
Financial Statement Analysis: Income Statement Analysis
We also note that Alico cannot dividend any of its 2010 earnings to AIG, implying that the income recognized this quarter and throughout 2010 will benefit MetLife results. At the respective midpoints, sales of \$8.5 billion would be down 22% annually and 8% sequentially; and non-GAAP EPS would be down 39%. The growth in operating income from Q1 2009 is largely due to recoveries in partnership income.
Financial Statement Analysis: Balance Sheet Analysis
Long-term borrowings of \$6.59 billion at September 30, 2022 were modest compared to shareholders’ equity of \$37.3 billion. The company finished 2Q16 with \$21.4 billion in cash and short-term investments, up from \$15.6 Growth & Valuation Analysis GROWTH ANALYSIS RISK ANALYSIS billion at the end of 4Q15. Inventory turns improved to 4.5x in 4Q from 4.4x in the same period last year.
Financial Statement Analysis: Cash Flow Analysis
UPS generated \$2.3 billion in operating cash for the quarter. Assuming that the company completes a large portion of its current \$1 billion stock buyback plan in 2008, we estimate that cash per share will be about \$6.20 by the end of the year. We view eBays FCF generation as relatively defensible even in the case of a revenue shortfall.
Financial Statement Analysis: Financial Ratios
The company achieves average scores on our three main measures of financial strength: leverage based on debt-to-cap, profitability and interest coverage. P/S is at 0.6x and EV/EBITDA is 7.7x. The index members currently trade at an average of 16.3-times trailing earnings, which is below the five-year average of 19-times.
Business Segments
The company is organized into three businesses, software, representing the majority of the company’s total revenues, hardware systems and services. Results in 2008 also benefited from the absence of the significant level of mark-to-market losses in the company’s Gas Marketing segment in 2007. The company operates five distinct segments: Americas (71% of FY15 profit); Europe, Middle East, and Africa (4% of FY15 profit); China and Asia Pacific (12% of FY15 profit); Channel Development (13% of FY15 profit).
Growth Strategies Visa has been especially active on the acquisition front over the last several months. Combined with the Horizon assets and an emerging pipeline, there is enough in AMGN’s portfolio to offset potential headwinds and allow the company to grow through its patent cycle. Specific areas where more investment may be needed: Over the past decade, ROK management often claim that Process automation represents the growth opportunity for the company, but its sales in this market have barely grown in recent years, despite its much smaller sales base compared with established incumbents.
Risk Factors
Risks to our BUY thesis have to do with global competition, changing user behavior, global macro uncertainty, and anything else that can affect FB’s relationship with members, its advertisers or its publishing partners. Risks to achieving our price target include: 1) Apple crushing PayPal; 2) increasing competition in the payments space; 3) heavy investment spending on marketing, point of sale, or technology; and 4) legislative action. In addition to the expenses incurred by patent challenges, product liability and other legal suits could occur and lead to additional liabilities and revenue loss, which could substantially change our financial assumptions.
Management and Governance
Top management changes can be unsettling, and the resulting uncertainty has caused 3M shares to decline. Chairman, President, and CEO Charles Ergen beneficially owns about 53.6% of DISH’s equity securities and has 90.5% voting power. Bill Johnson, the present CEO of Progress Energy, will become president and chief executive officer of the new company.
Environmental, Social, and Governance (ESG) Factors
The assessment of ESG (Environmental, Social & Governance) risks by Baptista Research includes a wide range of considerations that pertain to the long-term sustainability of a company. Sustainalytics assesses the degree to which a company’s enterprise business is affected by ESG issues. Failure to adequately address social risks like labor disputes and community relations could jeopardize the company’s social license to operate in certain regions.
Valuation
Our DCF derives an intrinsic value of \$100 for ABBV by discounting cash flows through 2024E, and assuming a -5% terminal growth & 7.6% WACC. We value MET based on a Sum-Of-The-Parts (SOTP) analysis based on our 2021 EPS estimate and using peer comps across each business segment. Our target price is based on a five-year discounted cash flow (DCF) valuation that employs a 5% discount rate and 20x terminal-year FCF multiple.
Investment Thesis The results for the back half of the year will still be complex and confusing given the purchase accounting impacts and the full quarter of HSBC, but we do believe that there is a pay-off at the end of the road. As good as it gets: With its record multi-year backlog, Boeing’s revenue profile over the rest of decade is generally considered secure, and expectations for execution and cash already appear high. Clearly, our Ford Investment Thesis, which was based in large part on our belief that Ford would be able to offset headwinds (slowing cyclical tailwinds in North America, weakness in South America, weak growth in Europe, slowing growth in China, and regulatory cost headwinds), has been thrown into question.
Appendices and Disclosures
Although the information contained in the subject report has been obtained from sources, we believe to be reliable, its accuracy and completeness cannot be guaranteed. For a complete discussion of the risk factors that could affect the market price of a company’s shares, refer to the most recent Form 10-Q or 10-K that a company has filed with the Securities and Exchange Commission. The Benchmark Company, LLC makes every effort to use reliable, comprehensive information, but we make no representation that it is accurate or complete.
Table A2: Numerical Variables Description
This table shows the definitions of numerical measures.
Numerical Measures Definition and/or sources
Panel A: Firm Level Measures
Size The market value equity of the firm ($CSHOQ*PRCCQ$) at the end of the quarter prior to report release.
BtoM The book value of equity ($SEQ+TXDB+ITCB-PREF$) scaled by the market value of equity ($CSHOQ*PRCCQ$) t the end of quarter prior the report release.
Prior_CAR The cumulated 10-day abnormal return ending 2 days before release. The abnormal return is calculated as the raw return minus the buy-and-hold market value-weighted return.
SUE Earnings surprise, calculated as the actual EPS minus the last consensus EPS forecast before the earnings announcement. Consensus EPS is the median value of 1-year EPS forecast within 90-days window of all analysts following the firm. The unexpected earnings is scaled by price per share at the fiscal quarter end.
AbsSUE Absolue value of SUE, representing the distance between realized EPS and EPS expectation.
Miss Dummy variable that equals one if the actual EPS is less than the last consensus forecast, and 0 otherwise.
Trading volume Trading volume at earnings announcement day (or the first trading day post earnings announcement), calculated as $VOL/SHROUT$.
Distance to default The distance to default calculated following Merton (1974). The proxy is compiled from the National University of Singapore’s Credit Research Initiative (NUS CRI).
Fluidity A measure of firms’ product market competition introduced by Hoberg et al. (2014). The data is compiled from the Hoberg and Phillips database.
Panel B: Industry Level Measures
Industry recession An indicator variable that equals one if the FF-48 industry return is negative and in the bottom quintile of FF-48 industry returns and zero otherwise.
Panel C: Macroeconomic Measures
Time trend (ttr) The number of years elapsed from the beginning of the sample.
Panel D: Report Level Measures
$REC\_REV$ Recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm.
$EF\_REV$ Earnings forecast revision, calculated as the current report’s EPS forecast minus the last EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
$TP\_REV$ Target price revision, calculated as the current report’s target price minus the last target price in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release.
Boldness An indicator variable that takes the value of 1 if the EPS forecast revision is above both the analyst’s own prior forecast and the consensus forecast, or else below both, and zero otherwise.
SR Stock recommendation from I/B/E/S rating, with 1 being the most bullish (Strong Buy) and 5 being the most bearish (Sell), based on the ratings provided by the Institutional Brokers’ Estimate System (I/B/E/S).
ERet 12-month return forecast by scaling the 12-month price target by the stock price 1-day before release.
Table A3: Sumamry Statistics for Numerical Measures
This table reports summary statistics of numerical measures.
Mean Std p25 P50 P75 N
Miss 0.35 0.48 0.00 0.00 1.00 102776
Trading volume 18.19 21.98 7.24 11.46 20.29 99908
Prior_CAR 0.00 0.05 -0.02 0.00 0.02 122251
Size 2.39 0.11 2.33 2.39 2.46 117534
BtoM 0.46 0.43 0.18 0.32 0.61 117534
Distance to default 0.00 0.01 0.00 0.00 0.00 86693
Fluidity 7.17 3.66 4.40 6.57 9.44 114942
Time trend 13.02 5.71 9.00 13.00 17.00 122252
Industry recession 0.14 0.35 0.00 0.00 0.00 122252
BrokerSize 71.92 56.71 26.00 53.00 113.00 119233
Firm Experience 7.68 6.72 3.00 6.00 11.00 111192
Number of firms 20.08 17.08 14.00 18.00 23.00 119233
REC_REV 0.00 0.16 0.00 0.00 0.00 120673
EF_REV 0.00 0.00 0.00 0.00 0.00 90625
TP_REV 0.01 0.07 0.00 0.00 0.00 84108
ERet 0.19 0.23 0.07 0.18 0.31 84108
Boldness 0.75 0.43 1.00 1.00 1.00 81182
SR 2.77 0.84 2.00 3.00 3.00 122252
Table A4: Information Content of Analyst Reports with Numbers Removed
This table presents the $R^{2}_{\mathrm{OOS}}$ using analyst reports with numbers removed. The input sets include revision numerical measures, text embeddings from number-free report text, and a combination of both. The estimation targets are the three-day $CAR$ centered around the analyst report release date. The machine learning model employed is Ridge regression, with training samples expanding yearly. The $t$-statistic for the $R^{2}_{\mathrm{OOS}}$ is calculated following Gu et al. (2020). In columns (2), (4), and (6), the benchmark estimation is set to zero. Columns (3)-(1) and (5)-(3) present pairwise tests comparing the performance of models using revision-only input versus text-only input, and text-only input versus combined text and revision input, respectively.
Year Revision only t-stat Text only t-stat Revision plus text t-stat t-stat t-stat
(1) (2) (3) (4) (5) (6) (3)- (1) (5)- (3)
2015 10.30% 3.59 13.19% 5.87 11.10% 3.24 3.97 -1.45
2016 15.47% 11.80 13.73% 5.00 17.91% 9.02 -0.81 3.13
2017 9.16% 5.28 11.66% 5.98 12.33% 6.12 10.58 1.56
2018 9.79% 2.98 11.46% 9.12 13.62% 7.97 1.22 4.09
2019 9.76% 15.69 12.61% 14.73 13.90% 15.17 4.26 2.79
2020 5.20% 3.43 4.42% 4.38 6.98% 5.81 -0.70 8.37
2021 5.31% 4.80 8.77% 5.06 10.66% 29.67 1.54 1.56
2022 10.24% 6.16 16.71% 10.47 17.90% 8.94 9.93 2.38
2023 5.09% 3.86 9.27% 3.60 10.39% 5.09 2.60 1.81
Overall 8.93% 8.51 10.95% 7.87 12.51% 8.83 2.40 3.00
Table A5: Cumulative Abnormal Returns with Alternative Windows
This table reports the $R^{2}_{\mathrm{OOS}}$ of using analyst report text embeddings to estimate $CAR$ with alternative windows. The $R^{2}_{\mathrm{OOS}}$ is calculated annually using a training sample from 2000 to the preceding year. $CAR$ measures are centered on the release day of analyst reports ($T_{0}$). Panel A shows results for the full sample, while Panels B and C focus on periods within and beyond 1-day earnings announcement windows, respectively. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
AR[0] CAR[0,+1] CAR[-2,+2]
year $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat $R^{2}_{\mathrm{OOS}}$ t-stat
Panel A: Full sample period
2015 8.03% 8.53 8.24% 8.36 10.87% 9.40
2016 7.88% 8.27 6.09% 4.68 6.60% 3.11
2017 8.75% 14.35 8.56% 9.29 9.00% 8.78
2018 9.09% 14.95 9.39% 11.26 9.75% 10.54
2019 9.89% 99.70 7.68% 9.45 8.66% 18.93
2020 2.33% 2.33 1.76% 3.15 4.13% 2.14
2021 5.75% 7.32 5.80% 6.62 4.05% 3.57
2022 10.54% 12.32 11.23% 18.20 10.91% 13.65
2023 5.38% 15.44 3.39% 7.00 0.53% 1.48
Overall 7.55% 8.07 6.92% 6.77 7.57% 6.79
Panel B: Earnings announcement period
Overall 9.78% 5.58 7.87% 5.10 13.14% 8.02
Panel C: Non-earnings announcement period
Overall 5.77% 5.78 5.62% 5.64 4.96% 4.91
Table A6: Information Content of Analyst Reports across Industries
This table reports the $R^{2}_{\mathrm{OOS}}$ of Fama-French 12 industries, using the definition from Kenneth French’s website. The Overall row reports the $R^{2}_{\mathrm{OOS}}$ and $t$-statistics for 2015-2023. The t-statistics for $R^{2}_{\mathrm{OOS}}$ are calculated using zero benchmarks estimation following Gu et al. (2020).
year Shops Other Manuf Chems
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 13.27% 6.03 657 17.69% 5.03 585 10.84% 3.19 808 15.05% 1.74 147
2016 10.43% 2.46 556 3.34% 0.48 593 10.37% 5.38 811 23.91% 3.25 111
2017 16.65% 3.49 506 20.45% 2.67 544 25.03% 13.64 722 11.04% 0.59 82
2018 10.86% 4.32 406 21.65% 3.40 405 3.32% 1.55 553 25.80% 1.87 77
2019 21.04% 9.24 394 7.49% 2.37 610 16.25% 12.07 568 5.00% 3.22 106
2020 4.39% 2.55 499 -0.30% -0.53 499 4.06% 1.41 542 2.65% 1.60 115
2021 10.27% 2.13 312 9.70% 3.66 362 8.66% 4.32 367 1.46% 0.30 102
2022 22.75% 3.83 320 15.98% 2.66 437 14.26% 11.46 420 7.00% 2.10 110
2023 17.78% 1.76 353 3.49% 2.00 339 7.15% 2.71 360 19.31% 6.54 76
Overall 13.88% 6.03 4003 11.27% 3.69 4374 11.39% 5.57 5151 11.35% 3.55 926
year Durbl BusEq Hlth NoDur
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 16.83% 5.77 216 15.66% 6.41 1378 5.75% 2.72 1242 0.68% 0.31 241
2016 4.95% 0.79 209 10.24% 6.09 1303 11.07% 2.90 1185 17.14% 5.18 242
2017 -2.17% -0.30 282 10.36% 3.12 1154 3.37% 2.39 1060 -2.14% -0.68 278
2018 18.29% 5.57 208 0.32% 0.31 1002 12.47% 6.58 918 1.29% 0.13 324
2019 15.31% 2.26 222 10.17% 6.13 1036 10.02% 7.68 884 18.51% 2.30 277
2020 5.73% 0.79 168 8.11% 4.14 1100 2.86% 1.84 950 -0.37% -0.28 265
2021 10.01% 5.31 150 11.43% 1.49 731 7.66% 4.58 592 -6.38% -0.60 195
2022 7.99% 2.58 162 12.08% 4.77 822 12.05% 2.71 850 5.24% 1.11 187
2023 12.05% 2.37 140 10.45% 11.80 962 3.52% 1.32 844 -5.69% -0.71 158
Overall 11.02% 3.24 1757 9.68% 6.77 9488 7.53% 5.71 8525 7.50% 1.38 2167
year Money Telcm Enrgy Utils
$R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N $R^{2}_{\mathrm{OOS}}$ t-stat N
2015 18.13% 4.41 1153 1.68% 0.00 403 -4.21% -1.45 303 18.47% 3.69 401
2016 7.64% 3.53 1108 -10.18% -1.22 329 15.80% 2.83 252 -1.84% -0.23 310
2017 1.77% 0.46 1071 8.60% 5.16 307 -18.18% -8.75 307 -39.46% -6.99 315
2018 13.55% 8.48 879 10.65% 2.46 219 8.38% 1.29 275 -23.76% -5.53 215
2019 2.10% 2.25 977 1.53% 0.52 351 -10.28% -9.07 340 -2.36% -0.27 193
2020 1.42% 0.76 952 -8.84% -2.83 360 5.72% 2.62 434 -7.29% -2.36 165
2021 2.62% 0.47 614 6.15% 3.27 188 -15.04% -6.83 251 -4.01% -1.16 103
2022 4.34% 2.13 667 12.84% 3.50 143 16.14% 4.44 286 8.31% 2.71 132
2023 -0.63% -0.18 569 8.99% 6.89 182 -27.33% -5.03 290 -6.43% -1.12 120
Overall 5.18% 3.31 7990 3.32% 1.39 2482 1.90% 0.31 2738 -1.97% -0.88 1954
Table A7: Summary Statistics for Alternative Analyst Information Value Estimations
The analyst information value is measured as the explained return variance divided by the price impact, with the results reported in millions of dollars. All dollar values are adjusted for inflation to reflect 2020 values using the Consumer Price Index (CPI). The sample comprises analyst reports targeting common constituent firms of the S&P 100 index from 2015Q1 to 2023Q4. Both mean and standard deviation are computed using the delta method.TAQ volatility indicates the explained return volatility is calculated as $\frac{r^{2}-(r-\widehat{r})^{2}}{r^{2}}\cdot\sigma_{v}^{2}$, where $\sigma_{v}^{2}$ is the realized volatility from one-minute log returns. Intraday denotes that $\sigma_{v}^{2}$ does not include overnight return volatility. EMO and CLNV refer to trade-signing algorithms developed by Ellis et al. (2000) and Lee and Ready (1991), respectively.
Mean SE 95%CI 99%CI N
Information value (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17669
Information value of text (TAQ volitility), \$M 0.47 0.05 [0.38, 0.56] [0.35, 0.58] 17672
Information value of revisions (TAQ volitility), \$M 0.42 0.04 [0.34, 0.50] [0.31, 0.53] 17672
Information value (Intraday), \$M 0.42 0.04 [0.34, 0.51] [0.32, 0.53] 17672
Information value of text (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17669
Information value of revisions (Intraday), \$M 0.38 0.04 [0.30, 0.46] [0.28, 0.48] 17672
Information value (EMO), \$M 0.34 0.04 [0.27, 0.41] [0.25, 0.43] 17672
Information value of text (EMO), \$M 0.35 0.04 [0.27, 0.42] [0.25, 0.44] 17672
Information value of revisions (EMO), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17669
Information value (CLNV), \$M 0.34 0.04 [0.26, 0.43] [0.23, 0.46] 17672
Information value of text (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.41] 17672
Information value of revisions (CLNV), \$M 0.31 0.04 [0.23, 0.39] [0.21, 0.42] 17672
C. Machine Learning Models
Ridge Regression
Ridge regression addresses multicollinearity by adding a regularization term to the least squares objective function. The ridge regression estimator is given by:
[…]
where $\alpha$ is the regularization parameter that controls the trade-off between fitting the data and shrinking the coefficients.
To find the optimal value of $\alpha$, cross-validation is used over a grid of values ranging from $10^{-10}$ to $10^{10}$. The cross-validation process ensures that the chosen model generalizes well to unseen data, preventing overfitting while capturing the predictive power of the text embeddings.
Partial Least Square Regression
To mitigate the risk of overfitting inherent in high-dimensional text embeddings, I employ Partial Least Squares (PLS) for dimensionality reduction.
The optimization problem can be expressed as follows:
[…]
where $\Omega$ is a $K\times P$ transformation matrix that reduces the $K$ predictors in $X_{i}$ to $P$ lower-dimensional components.
The extraction of the $j$-th PLS component is guided by the following objective function:
[…]
In essence, this approach sequentially extracts components that maximize the covariance with the outcome variable, while ensuring orthogonality to previously extracted components.
Extreme Gradient Boosting
Tree-based approaches are commonly applied in stock return forecasting literature (see, e.g., Gu et al., 2020; Cao et al., 2024; Bonini et al., 2023). XGBoost is an advanced implementation of tree-based machine learning models that builds an ensemble of decision trees. In XGBoost, each tree is added sequentially to correct the errors of previous trees. The main idea is to combine the outputs of multiple weak learners (decision trees) to create a strong learner. XGBoost incorporates regularization techniques, such as L1 (Lasso) and L2 (Ridge), to prevent overfitting and enhance model generalization.
In comparison to random forests, which build a multitude of independent trees and aggregate their predictions, XGBoost constructs trees sequentially, with each tree designed to correct the errors of the preceding ones. While random forests rely on bagging, a method that combines the predictions of various trees to reduce variance, XGBoost uses boosting, an approach that aims to reduce both bias and variance by focusing on difficult-to-predict instances in subsequent iterations. This boosting approach allows XGBoost to effectively capture and leverage the nuanced information embedded in textual data, leading to a more accurate estimation of stock returns.
Neural Networks
To extend beyond the linear modeling approach, I explore the use of Neural Networks to estimate $CAR$ using text embeddings derived from analyst reports. A Neural Network can capture complex non-linear relationships between the text embeddings and the $CAR$, potentially improving the model’s fitting capabilities. Consider a three-layer Neural Network as the prediction model. The prediction problem can be formulated as follows:
$f(X_{i};\theta)=W_{3}\sigma\left(W_{2}\sigma\left(W_{1}X_{i}+b_{1}\right)+b_{2}\right)+b_{3},$ (16)
where $\sigma(\cdot)$ is the ReLU activation function, $W_{i}$ and $b_{i}$ represent the weights and biases for layer $i$, respectively. The Neural Network architecture consists of an input layer representing the text embeddings, followed by multiple hidden layers. The specific architecture employed includes 32 neurons in the first hidden layer, followed by 16, 8, 4, and 2 neurons in subsequent layers. This structure is flexible and can be adjusted by adding or removing layers as necessary to optimize performance.
The training process involves optimizing the weights and biases to minimize the loss function, which, in this case, is the mean-square loss. To regularize the model and prevent overfitting, early stopping is implemented, halting training once the validation loss ceases to decrease. Additionally, following Gu et al. (2020), the models are retrained five times, and the final estimation is obtained by averaging the outputs of these five models, forming an ensemble estimation.
D. Theory
I provide an intuition for the measure of strategic value by discussing a simple extension of Kyle (1985) model. In the Kyle model, there is one risky asset with a payoff $\tilde{v}\sim N\left(p_{0},\Sigma_{0}\right)$. Three types of traders exist: a strategic trader with insider information, a market maker who sets prices in a perfectly competitive market, and an uninformed trader who trades $\tilde{u}\sim N\left(0,\sigma_{u}^{2}\right)$, where $\tilde{u}$ is independent of $\tilde{v}$. Illiquidity is measured by Kyle’s lambda ($\lambda$). Kyle’s lambda depends on private information and liquidity trading. I extend the model by considering a case where $\phi$ percentage of variance in $\tilde{v}$ is explained by the informed trader’s information.
A Single Auction Model
There are two periods, $t_{0}$ and $t_{1}$. The asset is traded with asymmetric information at period $t_{0}$, and the value $\tilde{v}$ is realized at period $t_{1}$. I assume without loss of generality that
$\tilde{v}=P_{0}+\tilde{s}+\tilde{\epsilon},$
where $\tilde{s}$ is a mean-zero signal observed by the informed trader at $t_{0}$, and $\tilde{\epsilon}$ is the combination of residual information and noise. I assume that the signal $\tilde{s}$ and residual $\tilde{\epsilon}$ are uncorrelated, that is, $\sigma_{v}^{2}=\sigma_{s}^{2}+\sigma_{\epsilon}^{2}$.
Let
$\phi=\frac{\operatorname{var}(\tilde{s})}{\operatorname{var}\left(\tilde{v}\right)}=\frac{\sigma_{s}^{2}}{\sigma_{v}^{2}}.$
This measure $\phi$ is the ”R-square” of projecting $\tilde{v}$ on $\tilde{s}$, or the percentage of explainable variance in $\tilde{v}$ using signal $\tilde{s}$.
After observing $\tilde{s}$, the informed trader submits a market order $\tilde{x}=X(\tilde{s})$, and the uninformed trader trades a zero-mean random variable $\tilde{z}$ that is normally distributed and independent of $\tilde{v}$. The profit of the informed trader is given by: $\tilde{\pi}=(\tilde{v}-\tilde{p})\tilde{v}$. The insider has a rational guess of $P(\tilde{x}+\tilde{u})$ and understands that his order $\tilde{x}$ will move the price against him.
The market maker observes the order flow $\tilde{y}\stackrel{{\scriptstyle\text{def}}}{{=}}\tilde{x}+\tilde{z}$ and determines the equilibrium price $\tilde{p}=P(\tilde{x}+\tilde{u})$ to break even. The assumptions for the market maker are risk neutrality and perfect competition, which drives the profits for market makers to zero.
Equilibrium
An equilibrium is a set of $X$ and $P$ satisfying
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s],$
$\tilde{p}(X,P)=E[\tilde{v}\mid\tilde{x}+\tilde{u}].$
The first condition is profit maximization, stating that given the market maker’s pricing rule, the insider chooses a strategy $X$ maximizing her conditional expected profit, taking into account the pricing rule. The second condition is market efficiency. Given the insiders’ trading strategy, the market maker sets the price to be the expected value of the security.
Conjecture:
$P(\tilde{y})=\mu+\lambda\tilde{y},$
$X(\tilde{s})=\alpha+\beta\tilde{s}.$
The profit of the insider can be written as:
$E[\tilde{\pi}(X,P)\mid\tilde{s}=s]$ $=E[(\tilde{v}-\mu-\lambda(\tilde{u}+x))x\mid\tilde{s}=s]$
$=(P_{0}+s+-\mu-\lambda x)x.$
Traders take into account that her order flow will move the price against her, which serves to restrain her position size.
Solving for optimal profit, I get:
$x^{*}=\frac{P_{0}+s-\mu}{2\lambda}=\alpha+\beta v.$
Hence, I can express $\alpha$ and $\beta$ :
$\begin{gathered}\beta=\frac{1}{2\lambda},\\ \alpha=\frac{P_{0}-\mu}{2\lambda}=(P_{0}-\mu)\beta.\end{gathered}$
When the market maker puts a higher weight on the order flow in setting the price, the trader puts a lower weight on her information.
I now look at the price-setting rule:
$\mu+\lambda y=E\{\tilde{v}\mid\alpha+\beta\tilde{s}+\tilde{u}=y\}.$
Essentially, the market maker observes a normally distributed signal about
$\mu+\lambda y$ $=E[\tilde{v}\mid y]$
$=\bar{v}+\frac{\operatorname{cov}(\tilde{v},\alpha+\beta\tilde{s}+\tilde{u})}{\operatorname{var}(\alpha+\beta\tilde{s}+\tilde{u})}(\beta\tilde{s}-\beta\bar{s}+\tilde{u})$
$=p_{0}+\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}}(y-\alpha).$
Hence, I can express $\lambda$ and $\mu$ :
$\lambda=\frac{\beta\sigma_{s}^{2}}{\beta^{2}\sigma_{s}^{2}+\sigma_{u}^{2}},$
$\mu=p_{0}-\alpha\lambda.$
There is a unique linear equilibrium given by
$\mu=P_{0},$
$\lambda=\frac{\sigma_{s}}{2\sigma_{u}},$
$\alpha=0,$
$\beta=\frac{\sigma_{u}}{\sigma_{s}}.$
Discussion
Kyle’s Lambda
The parameter $\lambda$ is universally known as Kyle’s lambda. Formally, it is the impact on the equilibrium price of a unit order. Its reciprocal ($1/\lambda$) measures the liquidity (or depth) of the market. If $1/\lambda=\frac{2\sigma_{u}}{\sigma_{s}}$ is larger, then the market is more liquid, either because there is less private information in $\sigma_{s}$ or there is more liquid trade in the sense of $\sigma_{u}$.
Value of Private Information
Notice that the equilibrium strategy of the informed trader is $\tilde{x}=\beta s$. The unconditional expected gain of the informed trader is
$\mathrm{E}[\tilde{x}[\tilde{v}-p(\tilde{x}+\tilde{z})]]$ $=\beta\mathrm{E}[s(\tilde{v}-\mu-\lambda\beta s-\lambda\tilde{u})]$
$=\beta(1-\lambda\beta)\sigma_{s}^{2}$
$=\frac{\sigma_{s}\sigma_{u}}{2}$
$=\frac{\phi\sigma_{v}^{2}}{4\lambda}.$
The expected gain for the informed trader is maximized when she has more private information or when there is more liquidity trading. On the other hand, liquidity traders incur losses equivalent to the informed trader’s gains, but they accept these losses due to other motives for trading. The equilibrium price ensures that market makers do not profit or lose in expectation.
E. Delta Method Approximation
I use the delta method to approximate the expected value and variance of a ratio of two random variables. This method relies on a first-order Taylor expansion. Specifically, consider two random variables $X$ and $Y$ with means $\mu_{X}$ and $\mu_{Y}$, and variances $\sigma_{X}^{2}$ and $\sigma_{Y}^{2}$, respectively. We are interested in the ratio $Z=\frac{X}{Y}$ and want to approximate the mean $E[Z]$ and the variance $\text{Var}(Z)$.
Delta Method Approximation
The key idea is to approximate the function $g(X,Y)=\frac{X}{Y}$ using a Taylor series expansion around the means $\mu_{X}$ and $\mu_{Y}$. For the function $g(X,Y)$, I use a first-order Taylor expansion around $(\mu_{X},\mu_{Y})$:
$g(X,Y)\approx g(\mu_{X},\mu_{Y})+\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}(X-\mu_{X})+\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}(Y-\mu_{Y}).$
The partial derivatives of $g(X,Y)=\frac{X}{Y}$ are:
$\frac{\partial g}{\partial X}=\frac{1}{Y},\quad\frac{\partial g}{\partial Y}=-\frac{X}{Y^{2}}.$
Evaluating these at $(\mu_{X},\mu_{Y})$, I get:
$\left.\frac{\partial g}{\partial X}\right|_{(\mu_{X},\mu_{Y})}=\frac{1}{\mu_{Y}},\quad\left.\frac{\partial g}{\partial Y}\right|_{(\mu_{X},\mu_{Y})}=-\frac{\mu_{X}}{\mu_{Y}^{2}}.$
Substituting the partial derivatives into the Taylor expansion, I obtain:
$\frac{X}{Y}\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}(X-\mu_{X})-\frac{\mu_{X}}{\mu_{Y}^{2}}(Y-\mu_{Y}).$
Taking the expectation on both sides:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}+\frac{1}{\mu_{Y}}E[X-\mu_{X}]-\frac{\mu_{X}}{\mu_{Y}^{2}}E[Y-\mu_{Y}].$
Since $E[X-\mu_{X}]=0$ and $E[Y-\mu_{Y}]=0$, the approximation simplifies to:
$E\left[\frac{X}{Y}\right]\approx\frac{\mu_{X}}{\mu_{Y}}.$
Using the delta method, the variance of $Z=\frac{X}{Y}$ is approximated by:
$\operatorname{Var}(Z)\approx\left(\frac{\partial g}{\partial X}\right)^{2}\operatorname{Var}(X)+\left(\frac{\partial g}{\partial Y}\right)^{2}\operatorname{Var}(Y)+2\left(\frac{\partial g}{\partial X}\right)\left(\frac{\partial g}{\partial Y}\right)\operatorname{Cov}(X,Y).$
Substituting the partial derivatives:
$\operatorname{Var}\left(\frac{X}{Y}\right)\approx\frac{\operatorname{Var}(X)}{\mu_{Y}^{2}}+\frac{\mu_{X}^{2}\operatorname{Var}(Y)}{\mu_{Y}^{4}}-2\frac{\mu_{X}\operatorname{Cov}(X,Y)}{\mu_{Y}^{3}}.$
Application to Value of Information
In the context of the provided study, I estimate the mean value of information for a subsample $s$ using the delta method.
Let $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}=\frac{r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ be the value of information for stock $i$ on date $t$. Define the mean and variance over subsample $s$ as:
$\mu_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(r_{if}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}\right),$
and the mean price impact per dollar over subsample $s$ as:
$\mu_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}.$
Using the delta method approximation for the mean of the ratio, the mean value of information over subsample $s$ is given by:
$E\widehat{\Omega}_{s}\approx\frac{\mu_{\nu s}}{\mu_{\lambda s}}.$
The variance of the ratio $\widehat{\Omega}_{it}=\frac{\widehat{\sigma}_{it}^{2}}{\widehat{\lambda}_{it}/P_{it_{-}}}$ over the subsample $s$ is estimated as:
$\text{Var}\left(\widehat{\Omega}_{s}\right)\approx\frac{1}{\mu_{\lambda s}^{2}}\left(\Sigma_{\nu s}+\frac{\mu_{\nu s}^{2}}{\mu_{\lambda s}^{2}}\Sigma_{\lambda s}-2\frac{\mu_{\nu s}}{\mu_{\lambda s}}\Sigma_{\nu\lambda s}\right),$
where $\Sigma_{\nu s}$, $\Sigma_{\lambda s}$, and $\Sigma_{\nu\lambda s}$ are defined as:
$\Sigma_{\nu s}=\frac{1}{|s|}\sum_{it\in s}\left(\widehat{\sigma}_{it}^{2}-\mu_{\nu s}\right)^{2},$
$\Sigma_{\lambda s}=\frac{1}{|s|}\sum_{it\in s}\left(\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}-\mu_{\lambda s}\right)^{2},$
$\Sigma_{\nu\lambda s}=\text{Cov}\left(\widehat{\sigma}_{it}^{2},\frac{\widehat{\lambda}_{it}}{P_{it_{-}}}\right).$
This provides a first-order approximation for the mean and variance of the value of information over the subsample $s$ using the delta method.